A Complete Bibliography of the Journal of Number Theory (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 May 2024 Version 1.30

Title word cross-reference

(1, 2) [203]. (2, 2) [947]. (2, 2n, 3) [583]. (4, 2, p) [945]. (χ, b) [770]. (m − 1, 1) [967]. 0 [189]. 1/2 [497]. 1/2 < ℜ(s) < 1 [999]. 16 [62]. 2 [66, 79, 702, 813, 852, 913]. 2^{n} − k [1120]. 3 [189, 751, 756, 923, 1041]. 3k − 4 [206]. 4 [205]. 5 [95, 189]. 7 [95, 189, 753]. 8k + 3 [681]. > [294]. ^{1} [1022]. ^{2} [214]. _{2} [300, 544]. _{4} [467]. _{n} [761]. _{n}F_{n−1} [1079]. A [544, 553, 1091]. {a, 3a} [454]. A_{n} [137]. abc [995]. α [781]. αβ [306, 462, 1108]. AX^{d} + C [204]. b [371]. s̄pt_{ω}(n) [735]. \binom{n}{k} = \binom{m}{l} + d [70]. modp^{k} [927]. C [551]. C_{1} [617]. C_{p^{2}} [713]. D [534, 641]. d(n) [631]. Δ_{(1)}(x) [614]. E_{5} [282, 446]. E_{7} [282]. E_{7,3} [908]. E_{8} [846]. ℓ [533, 956]. F [641, 761]. f(q) [430]. F_{1} [671]. F_{n} ± a(10^{m} − 1)9 = k! [752]. F_{n} ± F_{m} = y^{a} [319]. F_{q}[T] [627, 1012]. \frac{L′}{L} (1/2 + ϵ, χ_{D}) [703]. G [302, 448]. Γ [332, 572, 745]. Γ^{∗}(k) [579]. Γ_{0}^{+}(2) [382]. Γ_{0}^{+}(N) [585]. Γ_{1}^{Δ}(n) [543]. GL(2) [255, 259]. GL(2) × GL(2) [742]. GL(2) × GL(3) [1007]. GL(3) [259]. GL(3)L [1007]. GL(3) × GL(3) [1110]. GL_{2}(D) [1038]. GL_{2}(F) [402].

GL_{3}(R) [587]. GL_{N} [749]. GSp(4) [836, 1111]. GSp_{4}(Z_{p}) [850]. h^{1} ≠ h_{1} [413]. Ẑ [471]. j [170, 693, 898]. K [37, 106, 175, 258, 805, 917, 938, 1045, 1078, 1117]. K^{×}K^{×p^{m}} [809]. L [15, 20, 39, 44, 74, 75, 88, 116, 120, 121, 151, 178, 234, 251, 268, 271, 287, 293, 294, 340, 342, 345, 347, 363, 389, 435, 475, 529, 549, 582, 587, 602, 620, 627, 674, 686, 695, 697, 700, 731, 742, 747, 767, 779, 811, 826, 833, 835, 836, 847, 886, 920, 925, 948, 963, 976–978, 991, 1017, 1023, 1031, 1063, 1070, 1112, 1116, 1118, 1119]. L(\frac{1}{2}, χ) [136]. L(\frac{1}{2}, Sym^{2}f × g [1104]. L(\frac{1}{2}, Sym^{2}f × g) [7]. L(s) + L(2s) + ⋯ + L(Ns) [807]. L^{2} [954]. L^{2}(F_{2}((T))) [866]. L^{2}(Q_{2}) [866]. λ [386, 881, 959, 1024]. ⌊αp + β⌋ [411]. m [132, 571, 856]. M_{22} [494]. N [236]. P^{n} [35]. Q [25, 45, 224, 341, 494, 756, 923]. Q(−5l) [446]. Q(ζ_{2ℓ+1})^{+} [533]. Z/pZ [199]. Z[[x]] [183]. Z_{m} × Z_{n} [241]. Q(ap_{1}p_{2}) [740]. Q( ^{p} \sqrt{n}) [233]. p [472]. pm [472]. GL(2) [54, 61, 178, 810]. GL_{2} [76, 121]. GL_{3} × GL_{1} [121]. GL_{n} [83]. PGL_{n}(R) [13]. SL(3) [224]. SL_{2}(F_{p}) [32]. Co_{3} [11]. mn ≤ x [241]. μ [595]. N [56, 219, 245, 492, 944]. N(D) [641]. p [18, 47, 59, 126, 214, 320, 389, 402, 435, 491, 492, 526, 533, 566, 570, 575, 590, 651, 659, 666, 720, 786, 800, 811, 871, 886, 903, 904, 920, 922, 925, 927, 929, 952, 1042, 1076, 1079, 1081, 1088, 1092]. p + 2k_{1}^{r_{1}} + ⋯ + 2k_{t}^{r_{t}} [1055]. P^{+}(n) [379]. P^{+}(n + 1) [379]. p_{A}(n, k) [953]. PGL_{2}Q [763]. ψ_{n}^{2} [481]. q [301, 739, 889]. R [41, 415, 930]. R^{d} [466]. S [131, 294, 734]. S(t) [537]. S_{1}(t) [537]. σ [18, 1072]. σ(2n + 1) ≥ σ(2n) [219]. SL(3, Z) [911]. SL_{2} [1071]. SL_{2}(o) [829]. SL_{2}(Z) [884]. SL_{3}(Z) [883]. ∑_{j=1}^{k} jF_{j}^{p} = F_{n}^{q} [260]. T [54, 413, 418, 547, 549, 554, 742, 801, 1057]. θ_{3}(q) [242]. × [925]. U(1) [193]. U(n + 1) × U(n) [559]. U_{2n+1} × Res_{E/F} GL_{m}(m > n) [234]. U_{p} [317]. ε [47]. φ [1072]. φ_{n} [481]. x [37]. x^{2} + y^{2} + z^{2} + k [633]. x^{6} + ax^{3} + b [728]. X_{0}(14) [403]. X_{0}(N) [799]. X_{0}(p) [337]. y^{2} = x^{5} + ax [378]. Y^{2} = X^{6} + 1 [79]. Z [160, 169, 1062, 1096]. Z^{n} [591]. Z_{p} [526, 556]. ζ [332]. ζ(1/2 + it) [376, 537, 1030]. ζ(s) [848, 999]. ∥a^{x} − b^{y}∥ = k [710]. |L(1, χ)| [377, 630, 754].

-adic [18, 126, 320, 389, 402, 435, 566, 575, 590, 651, 659, 666, 720, 739, 786, 801, 811, 852, 871, 886, 889, 903, 920, 925, 927, 929, 952, 1088, 1092]. -arithmetic [734]. -ary [132]. -Aspect [54, 742]. -class [47, 702, 1081]. -colored [930]. -conjectures [47]. -continued [41, 944]. -coordinates [37]. -cycles [756]. -diamond [938]. -dimensional [923]. -divisible [169]. -elongated [813]. -extension [471]. -extensions [137, 526, 713]. -factor [572]. -factors [745, 770]. -function [18, 332, 742, 811, 835, 886, 1062]. -functions [15, 20, 39, 74, 75, 88, 116, 121, 151, 160, 234, 251, 268, 287, 293, 294, 302, 340, 345, 347, 363, 389, 435, 475, 487, 529, 587, 602, 620, 627, 674, 695, 697, 700, 731, 767, 779, 826, 833, 836, 920, 925, 963, 976– 978, 991, 1007, 1017, 1023, 1031, 1063, 1070, 1112, 1116, 1118, 1119]. -generalized [37, 1045]. -Genocchi [956]. -groups [415]. -harmonic [553]. -indivisible [917, 1117]. -invariant [170, 693, 898]. -invariants [386, 595, 881, 959, 1024]. -layered [258]. -modular [913]. -modules [547, 549, 1091]. -motives [413, 551]. -motivic [554]. -multipartition [418]. -norm [954]. -operators [317]. -orbits [306, 1108]. -parts [131]. -primary [781]. -products [301]. -rank [1078]. -rational [800]. -rationality [533, 904, 1042]. -rd [1041]. -repunits [371]. -scheme [671]. -series [120, 342, 686]. -spheres [56]. -term [751]. -th [175, 534, 570, 922]. -theory [805]. -torsion [491]. -tuple [571]. -values [44, 178, 271, 549, 582, 747, 948].

1 [294]. 1-498 [557]. 167 [90]. 186 [269]. 189 [888]. 194 [817]. 195 [452]. 196 [113].

200 [758]. 2019 [506]. 202 [372, 1104]. 2020 [211, 230, 247]. 2021 [272, 289, 311, 329, 350, 374, 393, 409, 428, 442, 456, 478]. 2022 [517, 557, 580, 603, 618, 660, 706, 729, 759]. 2023 [789, 819, 843, 861, 874, 890, 900, 918, 940, 950, 961, 974]. 2024 [992, 1003, 1014, 1033, 1050, 1067, 1089, 1105]. 206 [184]. 209 [310]. 216 [373]. 217 [602]. 218 [991]. 219 [818]. 239 [960]. 240 [889]. 246 [959]. 251 [1088].

A1 [9]. A2 [9]. ABC [158, 346]. Abel [93]. Abelian [4, 17, 40, 60, 155, 201, 257, 276, 365, 452, 482, 521, 530, 549, 573, 673, 678, 709, 738, 788, 793, 880, 916, 984, 1076]. abélienne [530]. abéliennes [738]. Absolute [85, 314, 671, 719, 721, 807]. Absolutely [365]. action [111]. Addendum [90]. addition [780]. Additive [30, 55, 73, 105, 236, 392, 709, 845, 879, 893, 954]. additively [978]. adeles [130]. Adelic [130, 648]. adic [18, 126, 320, 389, 402, 435, 566, 575, 590, 651, 659, 666, 720, 739, 786, 801, 811, 852, 871, 886, 889, 903, 920, 925, 927, 929, 952, 1088, 1092]. admitting [577, 898]. affine [167, 663]. affinoids [209, 801]. after [149]. against [867]. Alfonsín [869]. algebra [61, 191, 834]. Algebraic [1, 22, 53, 63, 65, 90, 104, 109, 296, 332, 383, 435, 441, 487, 497, 568, 617, 637, 726, 737, 743, 746, 748, 777, 846, 923, 943, 1008, 1038]. Algebraic-integer [748]. Algebraicity [293, 529, 935, 1079]. algebras [12, 24]. algorithm [705, 784, 860]. Algorithmic [387]. algorithms [944]. Alladi [339]. Almost [107, 138, 152, 202, 338, 356, 514, 594, 623, 965]. Almost-prime [107, 152]. along [123]. alternate [935, 998]. alternating [198, 459, 732, 776]. always [1117]. among [36]. Anabelian [721]. analog [82, 276]. analogs [385, 1075]. analogue [53, 264, 364, 808]. Analogues [339]. analysis [599]. analytic [134, 147, 291, 516, 666, 768]. Anderson [169, 413, 1091]. Andrew [787]. Andrews [228, 495, 504]. angles [744]. Ankeny [1042]. Annihilators [171]. annuli [613]. anti [678]. anti-cyclotomic [678]. Anticyclotomic [496, 595, 948]. Apéry [433]. Apéry-like [433]. Apollonian [208]. Appendix [560, 787]. application [813, 913, 931, 1113]. Applications [51, 102, 124, 430, 535, 638, 651, 655, 734, 877, 1109]. applying [7, 1104]. approach [858, 997]. approximants [368]. approximately [937]. approximation [127, 221, 222, 320, 407, 762, 839, 849, 926, 955, 1100]. Approximations [886]. April [71, 310, 329, 557, 861, 1033]. Arakelov [130]. arbitrary [783]. arboreal [112, 694]. Archimedean [982]. areal [996]. argument [420, 668]. arguments [107, 715]. arise [691]. arising [65, 84, 666, 893, 1076]. Arithmetic [100, 130, 150, 157, 210, 262, 369, 388, 391, 436, 440, 445, 449, 460, 501, 504, 510, 535, 601, 629, 651, 656, 688, 694, 734, 756, 778, 785, 809, 816, 832, 833, 894, 1006, 1109]. Arithmetically [937]. arrangements [335]. Artin [175, 176, 200, 620, 669, 697, 873, 1042]. Artin-Chowla [1042]. ary [132]. Asai [76, 340]. Aspect [39, 54, 742]. associated [89, 585, 638, 648, 683, 743, 808, 828, 944, 972, 1017, 1087]. associators [256]. Asymptotic [93, 123, 168, 277, 502, 625, 628, 634, 677, 838, 945, 1029]. asymptotics [6, 758, 963]. Atkinson [848]. attached [367]. attack [370]. Attractor [1065]. Attractors [733]. August [163, 409, 918, 1105]. automata [1048]. automatic [180]. automorphic [39, 300, 770, 771, 825, 826, 928, 1116, 1119]. automorphism [119, 365, 743]. automorphisms [970]. Average [2, 59, 156, 159, 269, 638, 1110]. Averages [638, 977]. avoiding [422]. Ax [982]. ayant [106].

Backward [870, 1018]. balanced [532, 625, 886]. Banff [651]. Banks [138]. Barnes [763]. Barry [465]. base [58, 864, 997, 998]. based [600]. bases [277, 439, 477, 895, 935]. basilica [694]. basis [544, 566, 692, 740, 1057]. Bass [431, 934]. Bateman [68, 254]. Bateman-like [254]. BBGS [334]. BBP [858]. BBP-type [858]. be [617]. Beatty [57, 243]. behavior [357, 715]. below [228]. Belyĭ’s [528]. Benford [723]. Bernoulli [555]. Berthelot [653]. Bessel [836]. Bessenrodt [953]. beta [777, 1020]. beta-expansion [1020]. between [93, 110, 146, 338, 404, 600, 744, 879, 897, 917, 1002, 1109, 1117]. beyond [1120]. bi [1016]. bi-quadratic [1016]. Bianchi [575, 622, 811, 920]. Bias [69, 229, 824, 867]. biases [917, 1117]. bidegree [203]. Bielliptic [245]. Biennial [506]. Bifurcations [915]. Big [64, 738, 850, 982]. bilinear [322, 794]. Binary [55, 172, 179, 208, 387, 610, 622, 664, 674, 683, 743, 933, 1114]. binomial [36]. biprojective [203]. biquadratic [125, 458, 473, 498]. Birch [67, 872]. block [696]. Board [10, 27, 50, 72, 98, 118, 141, 164, 187, 212, 231, 248, 273, 290, 312, 330, 351, 375, 394, 410, 429, 443, 457, 479, 505, 518, 539, 558, 581, 604, 619, 642, 661, 707, 730, 760, 790, 820, 844, 862, 875, 891, 901, 919, 941, 951, 962, 975, 993, 1004, 1015, 1034, 1051, 1068, 1090, 1106]. Bogomolov [512]. Bombieri [486]. Borcherds [412, 698, 933]. Bordellès [635]. Bose [668]. both [385, 560]. bound [44, 54, 123, 161, 175, 363, 464, 1037, 1116]. bounded [692]. Bounding [693, 859]. Bounds [116, 142, 149, 188, 298, 343, 377, 386, 416, 430, 462, 591, 624, 647, 704, 754, 794, 878, 912, 1002, 1028, 1080]. boxes [381]. brackets [197, 1053]. Branches [12]. Brauer [205, 640, 679, 1000]. Brocot [675]. broken [938]. Bruhat [1075]. BSD [593]. bundle [438]. bundles [474]. Burgess [110, 464].

C [283]. Calabi [108, 654, 1005]. calculating [48, 641]. Cantor [280]. Capitulation [530, 916]. Carlitz [550, 552, 553, 555, 674, 737]. Carmichael [460]. Cartan [122]. Case [416, 559, 704, 834, 840, 857, 957, 1114]. Catalan [416, 436]. Category [551]. Central [24, 44, 75, 271, 475, 529, 571, 634, 731, 805, 859, 1007, 1112]. Certain [4, 21, 42, 100, 104, 114, 147, 382, 452, 488, 504, 609, 626, 728, 767, 914, 957, 968, 973, 990, 1064, 1108, 1116]. certaines [21]. Cesàro [447]. Ceva [356]. changes [614]. Characterisation [983]. characteristic [200, 209, 528, 663, 713, 979, 1091]. characters [151, 318, 976, 1017, 1112]. Chebotarev [243, 787]. Chebyshev [867]. checking [705]. Chen [988]. Chevalley [916]. Chow [979]. Chowla [668, 1042]. Chromatic [292]. Chudnovsky [785]. Cilleruelo [1099]. circle [154, 208, 438, 637, 810, 1046]. circles [489]. circulant [63]. Class [47, 96, 123, 171, 188, 220, 233, 240, 254, 359, 444, 446, 455, 460, 471, 519, 530, 538, 550, 583, 646, 702, 799, 839, 896, 916, 977, 1016, 1056, 1059, 1081, 1092, 1097]. classes [530, 822, 917, 1117]. classical [216, 331, 572, 588, 605]. classicality [782]. Classification [19, 43, 285, 387, 792]. Classifying [732]. closed [427, 485]. closure [233, 871]. CM [121, 157, 276, 665, 738, 793, 863, 877, 946, 965, 1024, 1052]. CM-points [157]. co [392]. co-minimal [392]. cocycle [763]. cocycles [622]. coefficient [307, 972]. Coefficients [36, 66, 189, 224, 242, 283, 301, 313, 419, 423, 430, 477, 483, 509, 552, 561, 609, 628, 701, 736, 868, 909, 911, 914, 928, 954, 957, 969, 977, 1005, 1010, 1039, 1058, 1064, 1107]. Cohen [197, 1053]. Cohomology [65, 83, 259, 449, 467, 509, 568, 651, 883, 884, 1041, 1054]. coincident [294]. Collision [1047]. colored [930, 985]. Combinations [294, 382]. combinatorics [30, 535]. coming [565]. Comments [274]. Common [177, 432, 481, 571]. commutative [491, 516]. commutator [724]. compact [157, 785, 823]. compactification [291, 543]. companion [267]. Comparing [496, 902]. Comparison [560]. complement [263]. complements [168, 390, 893]. complete [43]. Completed [130]. completely [180]. complex [56, 91, 150, 667, 802, 972, 1065]. components [863]. composite [176]. Composition [91, 1021]. compositions [739, 889]. computation [482, 990]. Computations [1086]. compute [899, 1121]. Computing [427, 650]. concavity [953, 1078]. concerning [241]. Conditional [175]. conditions [243, 673, 1032]. conductor [155, 582]. Conference [506, 651]. confirmed [399]. Congruence [450, 520, 656, 673, 813, 883, 884, 917, 930, 1075]. Congruences [18, 189, 266, 317, 361, 435, 455, 472, 495, 722, 735, 956, 985, 1044, 1066]. Congruent [496, 513, 681]. conifold [1005]. Conjecture [23, 35, 68, 73, 96, 138, 176, 192, 226, 229, 239, 244, 308, 348, 352, 353, 355, 362, 373, 379, 399, 435, 484, 512, 513, 534, 559, 635, 644, 680, 690, 700, 724, 753, 798, 804, 853, 869, 872, 916, 927, 934, 948, 988, 995, 1026, 1046, 1054]. conjectures [47, 338, 346, 379, 465, 499, 531, 586, 964, 1042]. conjugates [726]. connected [132, 863]. connection [110, 881, 959]. connections [30]. Consecutive [146, 338, 632, 897, 904]. consequence [512]. considerations [227]. Constant [153, 170, 882, 981]. constants [35, 110, 314, 594, 667, 1037]. constraints [501]. constructed [532]. Constructing [147, 360, 461]. Construction [678, 846, 1053, 1065]. Constructions [169, 612]. containing [426]. continuation [134, 768]. Continued [3, 41, 149, 222, 385, 476, 667, 733, 841, 851, 888, 944, 980]. Continuous [489]. contributions [345]. convergence [368, 488, 807]. Converse [344, 588]. convex [613]. Convolution [30, 249, 515, 810, 1064, 1095, 1107, 1110]. Conway [11]. Cooper [355]. coordinates [37]. Coppersmith [370]. Coprime [611]. coranks [496]. correlation [52, 313, 503, 909, 1036, 1084]. correlations [589, 623, 827]. correspondence [340, 514, 560]. correspondences [78, 310]. Corrigendum [184, 269, 310, 372, 373, 452, 538, 602, 758, 817, 818, 888, 889, 959, 960, 991, 1088, 1104]. coset [939]. cotangent [438]. Counting [157, 159, 203, 466, 607, 677, 822, 923, 1025, 1059]. counts [750, 942]. courtes [21]. cover [224, 1071]. Covering [494]. covers [200]. cranks [815]. critère [714]. criteria [414, 421]. criterion [664, 916]. Critical [121, 529, 576, 634, 686, 978, 1018, 1043, 1070]. crop [489]. crystal [223]. crystalline [85, 666]. crystallographic [38]. Crystals [654]. cube [340, 1025]. cubes [210, 262]. cubic [14, 15, 39, 303, 335, 405, 473, 641, 683, 755, 824, 885, 1000]. cubic-line [335]. cubics [304, 818]. curve [79, 139, 209, 327, 427, 448, 543, 665, 766, 828, 1013]. curves [44, 78, 89, 105, 122, 159, 173, 182, 188, 200, 229, 232, 245, 253, 263, 267, 303, 310, 318, 328, 337, 341, 364, 381, 405, 408, 419, 437, 444, 468, 470, 481, 500, 513, 521, 525, 565, 598, 693, 727, 741, 792, 797, 812, 852, 885, 898, 946, 965, 995, 1024, 1031, 1052, 1103]. Cusp [13, 66, 178, 279, 313, 590, 624, 628, 649, 878, 954]. Cuspidal [259, 467, 597, 770, 799, 859, 909]. cycle [119]. cycles [64, 324, 400, 672, 756]. cyclic [171, 473, 527, 638, 755, 837, 1081]. cyclotomic [28, 453, 471, 480, 550, 678, 719, 881, 959, 1024].

Dai [635]. Davenport [110, 160]. David [540]. December [247, 478, 759, 974]. decidability [780]. decomposable [215]. decomposition [325, 771]. Dedekind [156, 217, 380, 395, 697, 815, 943]. Deducing [492]. definability [994]. defined [265, 728, 845, 1052]. definition [802]. Deformations [652]. degeneracy [263]. degenerate [31, 80]. degeneration [448]. Degree [63, 66, 145, 205, 307, 640, 842, 857, 871, 970, 977, 987]. degree-2 [842]. degrees [984]. del [205, 640]. Delange [1006]. delta [54]. Demailly [35]. denominators [753, 1100]. dense [711, 845, 1108]. Densities [93, 133, 345, 1027]. Density [29, 154, 155, 347, 418, 492, 495, 504, 684, 699, 766, 787, 828, 1060, 1063, 1119]. d’entiers [106]. dependent [632]. Depth [191, 256, 1111]. Depth-graded [191]. depths [560]. d’Erdos [106]. derivative [75]. derivatives [104, 500, 765, 1062, 1070, 1118]. descent [605, 658]. description [223]. desingularized [523]. detecting [731]. Determination [702]. Deuring [274]. developments [643]. Diagonal [368, 401, 626, 876]. diagonalizable [469]. diamond [938]. Dichotomous [942]. difference [129]. differences [199, 503, 573, 584]. different [439]. differential [332, 524, 656, 1053]. differentials [232, 253, 982]. Digit [439, 777, 915]. digital [490]. digits [305, 354]. dihedral [45, 871]. dimension [62, 190, 365, 462, 476, 567, 617, 833, 955, 1108]. dimensional [62, 237, 568, 923, 1084]. dimensions [52, 612, 772]. diophantine [70, 87, 114, 127, 221, 270, 407, 454, 488, 621, 752, 755, 762, 806, 849, 955, 989, 994, 1100, 1101]. Dirichlet [16, 20, 116, 151, 184, 223, 251, 268, 287, 318, 345, 354, 475, 602, 683, 701, 723, 847, 976, 986, 991, 1017, 1023, 1058, 1062, 1063, 1112, 1118]. discrepancy [613, 1084]. Discrete [251, 501, 721, 765, 855, 1118]. Discriminant [641, 645, 659]. Discriminant-stability [659]. discriminants [32]. Disjointness [31, 1046]. display [1117]. displays [654]. Distance [600, 726]. distances [146]. distinct [1093]. distinguished [357, 761]. Distinguishing [1010]. Distribution [15, 109, 139, 159, 160, 195, 224, 238, 250, 306, 341, 383, 440, 502, 703, 725, 744, 914, 1102]. Distributions [358, 887, 1103]. divided [584, 655]. Divisibility [29, 132, 170, 384, 780, 872, 914, 946, 1032, 1082]. divisible [169, 781, 887, 1019]. division [12, 61, 468, 521, 570]. divisor [316, 445, 799, 808, 1107]. divisors [69, 89, 170, 177, 263, 391, 711, 716, 870, 956, 960, 1010]. Domain [254, 357]. domains [613, 791, 1065]. Double [7, 65, 108, 224, 298, 605, 704, 878, 939, 1005, 1104]. doubling [261, 572, 837]. drawn [489]. Drinfeld [143, 235, 291, 357, 441, 472, 541– 546, 548, 564, 645, 646, 648–650, 972, 1022, 1047, 1057]. dual [62, 733]. duality [903]. due [653]. Duke [697]. DVRs [144]. Dwork [111]. Dyer [67, 872]. Dynamic [126]. Dynamical [34, 165].

Editorial [10, 27, 50, 72, 98, 118, 141, 164, 187, 212, 231, 248, 273, 290, 312, 330, 351, 375, 394, 410, 429, 443, 457, 479, 505, 518, 539, 558, 581, 604, 619, 642, 661, 707, 730, 760, 790, 820, 844, 862, 875, 891, 901, 919, 941, 951, 962, 975, 993, 1004, 1015, 1034, 1051, 1068, 1090, 1106]. Effective [142, 263, 286, 490, 528, 630, 658, 775, 783, 912, 926]. Egyptian [796]. Eichler [636]. eigenform [913]. eigenforms [120, 197, 561]. eigenvalue [712]. eigenvalues [266, 921]. eight [327]. Eisenstein [134, 224, 309, 331, 382, 412, 585, 597, 736, 751, 815, 906, 952, 1044, 1086, 1088]. elementary [161, 265, 858, 1054]. elements [60, 299, 327, 640]. Elliott [23, 379]. Elliptic [29, 44, 89, 105, 139, 159, 170, 173, 188, 267, 318, 328, 335, 341, 364, 408, 412, 444, 468, 470, 481, 500, 513, 521, 565, 598, 665, 693, 705, 741, 766, 812, 828, 882, 885, 898, 946, 965, 1024, 1031, 1052, 1103]. elongated [813]. embedding [984]. encoded [809]. Endomorphism [143, 482, 546, 645, 650]. endomorphisms [343, 545, 845]. Engel [305]. ensemble [136]. Entire [25, 263, 388]. entries [803]. enumerating [99]. Epsilon [333]. Epstein [855]. equal [209, 349, 840]. equation [70, 96, 152, 260, 270, 319, 416, 524, 576, 710, 752, 811]. Equations [37, 87, 220, 399, 520, 536, 583, 626, 629, 712, 806, 840, 989, 1053, 1071, 1085, 1101]. Equidistribution [51, 213, 438, 1029]. equilateral [356]. equivalence [422, 832]. equivalent [336, 937]. Equivariant [549]. Eratosthenes [1001]. Erdos [106, 110, 270, 308, 338, 407, 490, 690, 708]. Ergodic [556, 997]. ergodicity [838]. Erratum [113]. error [217, 241, 447, 689, 848, 1084]. Esnault [653]. Estimate [400, 1030]. Estimates [208, 225, 534, 537, 630, 697, 775, 958, 1102, 1119]. Estimation [21]. eta [504, 609, 815, 939, 957]. eta-quotients [504, 609, 939, 957]. étale [67, 278, 449, 509]. Euclidean [622, 841, 1016]. Euler [8, 22, 87, 105, 128, 314, 332, 372, 671, 678, 851, 976]. Evaluations [996]. even [209, 228, 513, 907, 957, 1060]. even-regular [957]. Eventual [1078]. everywhere [898]. exact [579]. examples [656]. Exceptional [14, 116, 125, 283, 562, 836, 908]. exchanges [777]. Existence [402, 667, 866]. Existential [780, 994]. expanders [281]. Expanding [237, 425]. expansion [124, 385, 634, 1020, 1039]. expansions [41, 103, 244, 305, 439, 490, 567, 667, 677, 739, 777, 889, 915, 997, 998]. expected [662]. Explicit [13, 116, 181, 395, 422, 464, 537, 697, 740, 754, 848, 1002, 1030, 1102]. exponent [370, 840]. Exponential [225, 252, 279, 298, 399, 563, 695, 704, 722, 767, 801, 864, 868, 931, 1101]. exponential-type [563]. Exponents [51, 586, 849]. expressions [1082]. Ext [1022]. extended [1074]. extending [176]. extension [471, 566, 574]. Extensions [33, 40, 45, 84, 137, 145, 147, 202, 325, 441, 461, 526, 550, 713, 718, 751, 805, 832, 871, 1081]. Extrema [806]. extremal [913]. Extreme [1058].

facteurs [106]. factor [23, 572]. factorial [22, 851, 1074]. factorials [881, 959]. Factorization [265, 868, 924, 1021]. factorizations [596]. factors [76, 106, 201, 333, 434, 561, 588, 681, 711, 745, 770, 1071]. Faltings [793]. Families [28, 159, 229, 575, 681, 691, 731, 732, 776, 783, 854, 948, 964, 971, 1005, 1060]. family [133, 347, 378, 399, 564, 766, 813, 1047, 1086]. far [566]. Farey [103, 499]. faster [786]. February [26, 289, 517, 819, 1003]. Fekete [816]. Fermat [96, 253, 321, 416, 437, 583, 704, 945]. Fibonacci [37, 101, 158, 371, 424, 432, 532, 1045, 1049, 1120]. fibrations [640]. fibres [882]. field [85, 121, 171, 276, 299, 327, 366, 385, 524, 617, 646, 699, 783, 986, 1075, 1092]. Fields [4, 5, 14, 28, 67, 88, 96, 99, 113, 119, 144, 145, 155, 162, 170, 176, 181, 204, 225, 252, 287, 308, 320, 326, 328, 342, 345, 365, 389, 396, 405, 441, 452, 458, 468, 477, 497, 498, 519, 526, 534, 551, 563, 566, 568, 572, 577, 588, 606, 615, 617, 626, 641, 643, 647, 659, 667, 670, 676, 702, 721, 728, 751, 755, 762, 766, 767, 780, 786, 787, 800, 802, 832, 845, 852, 853, 856, 867, 868, 871, 881, 883, 898, 904, 906, 937, 945, 955, 959, 968, 991, 1016, 1024, 1052, 1063, 1065, 1081, 1092]. fields [84, 216, 453, 942]. fifth [639]. figurate [194]. filtration [696]. Finding [385, 545, 1121]. fine [364]. finie [738]. Finite [19, 53, 119, 202, 204, 257, 299, 333, 358, 365, 389, 405, 426, 491, 551, 553, 573, 588, 626, 699, 718, 738, 767, 774, 788, 791, 845, 942, 968, 1096]. finitely [441]. Finiteness [85, 568, 575]. finitude [1081]. Fintzen [560]. Fiori [787]. First [7, 307, 506, 656, 1104, 1112]. Fitting [550]. fixed [29, 1017, 1020, 1112]. flags [773]. flat [491, 1096]. focus [370]. Ford [146]. Foreword [540]. form [81, 215, 411, 446, 681, 743, 853, 972, 1055, 1120]. formal [385, 524]. forms [13, 66, 95, 103, 167, 174, 178, 179, 196, 197, 207, 208, 271, 275, 279, 282, 286, 300, 307, 313, 322, 357, 361, 367, 396, 401, 421, 422, 451, 472, 496, 497, 529, 541, 542, 544, 565, 575, 590, 593, 599, 601, 610, 622, 624, 628, 636, 643, 647, 649, 664, 683, 685, 716, 717, 749, 757, 769, 770, 779, 782, 811, 814, 831, 846, 876–878, 884, 907, 909, 911, 912, 928, 932, 954, 960, 967, 972, 990, 1008, 1027, 1057, 1066]. formula [7, 42, 67, 216, 278, 339, 364, 508, 574, 593, 608, 625, 628, 665, 848, 916, 1060, 1104]. formulae [301, 785]. formulas [276, 412, 773, 858, 902, 1044]. four [327]. Fourier [66, 124, 224, 307, 561, 628, 736, 911, 914, 928, 954, 1010, 1064, 1107]. fourth [279, 287, 529, 991]. fraction [41, 149, 385, 667, 944, 980]. fractional [55, 688]. fractions [3, 222, 476, 733, 796, 841, 851, 888]. Fredholm [629]. free [245, 257, 395, 633, 671, 1032, 1115, 1119]. Freiman [983]. French [21, 106, 530, 738]. frequencies [915]. friable [379]. Fricke [988]. Frobenius [115, 650, 965, 1083]. full [103, 984]. function [4, 5, 8, 18, 28, 48, 67, 69, 82, 87, 99, 113, 124, 128, 132, 148, 155, 159, 170, 176, 216, 217, 226, 236, 242, 249, 252, 264, 268, 276, 287, 308, 320, 332, 345, 366– 368, 372, 395, 418, 420, 430, 440, 441, 452, 489, 492, 534, 563, 572, 576, 579, 602, 607, 615, 638, 643, 644, 647, 671, 672, 677, 691, 742, 762, 765, 766, 775, 783, 795, 808, 811, 815, 823, 832, 835, 855, 868, 886, 892, 894, 953, 955, 986, 991, 1012, 1020, 1036, 1049, 1061–1063, 1074, 1075, 1084, 1093, 1109]. Functional [523, 536, 576, 811, 1071]. Functions [15, 20, 25, 39, 56, 74, 75, 88, 92, 100, 116, 121, 148, 151, 160, 189, 234, 251, 254, 268, 283, 284, 287, 293, 294, 297, 299, 302, 316, 340, 345, 347, 363, 388, 389, 434, 435, 445, 450, 475, 483, 487, 523, 529, 536, 556, 563, 584, 585, 587, 592, 602, 611, 620, 627, 638, 674, 688, 695, 697, 715, 716, 731, 767, 774, 779, 826, 829, 833, 836, 839, 899, 920, 925, 929, 933, 938, 960, 963, 969, 973, 976, 978, 991, 1006, 1007, 1017, 1031, 1059, 1063, 1064, 1070, 1071, 1091, 1107, 1112, 1113, 1116, 1118, 1119]. functions [198, 700, 977, 1023]. fundamental [32, 357, 616, 834]. Furstenberg [838]. further [165].

G.C.D. [29]. Galois [14, 19, 33, 85, 111, 112, 123, 147, 155, 179, 202, 232, 315, 325, 326, 336, 360, 493, 494, 595, 648, 694, 713, 721, 732, 776, 792, 809, 821, 850, 871, 966, 987, 990, 1054, 1060, 1076, 1077, 1121]. Gamma [76, 588, 1071, 1109]. Gan [559]. Gap [384, 480, 1097]. gaps [338, 1002]. Gauss [344, 672, 676, 881, 959]. GCD [534, 535]. General [16, 35, 80, 184, 333, 628, 850, 896, 903, 987]. generalisation [635]. Generalised [166, 532]. Generalization [32, 102, 239, 373, 463]. Generalizations [944, 1074]. Generalized [37, 64, 156, 160, 253, 321, 432, 445, 485, 569, 583, 612, 772, 804, 853, 936, 1009, 1045, 1079]. Generalizing [725]. generated [441]. generating [244]. generation [646]. generators [193, 359, 538]. generic [286, 448, 527, 559, 645, 1013]. Genocchi [956]. Genus [47, 78, 79, 278, 310, 327, 448, 520, 640, 736, 778, 797]. genus-zero [778]. geodesics [427, 510]. geometric [616, 832]. Geometry [130, 440, 473, 734, 839]. Germain [365, 533]. Gevrey [1109]. given [133, 175, 601, 631, 641, 751]. Glasner [803]. global [324, 386, 616, 617, 641, 662, 724, 780, 797, 825, 832, 867, 946, 955, 1060]. Goldbach [125, 183, 562]. golden [192, 385]. Goldfeld [513]. Gonek [574]. Good [209, 812, 898]. Gordon [227]. Goss [540, 739, 889]. Graded [95, 191, 207]. Graham [270, 499]. graph [223, 1043, 1091]. graphs [46, 63, 252, 854, 1023]. greatest [177, 481]. Greenberg [804, 853]. Gross [122, 559, 902]. grosse [738]. Grosswald [239, 373]. group [11, 45, 60, 85, 103, 111, 155, 171, 202, 205, 223, 233, 240, 359, 408, 458, 491, 538, 694, 721, 732, 743, 776, 788, 799, 829, 845, 850, 908, 916, 986, 987, 1096, 1097]. group-theoretic [721]. groupes [530]. groups [19, 47, 80, 216, 241, 257, 261, 320, 325, 326, 331, 333, 364, 365, 370, 415, 471, 494, 496, 514, 530, 550, 560, 568, 572, 573, 588, 605, 616, 622, 669, 673, 678, 684, 685, 702, 721, 770, 778, 784, 785, 821, 837, 863, 903, 947, 979, 987, 988, 990, 1052, 1077, 1081]. growing [757]. Growth [259, 305, 328, 467, 645, 928, 1052].

Halberstam [23, 379]. half [92, 698, 829, 831, 1005]. half-conifold [1005]. half-integral [698, 829, 831]. Hamming [58]. Hardy [102, 851, 1062]. harmonic [553]. Harris [348]. Hasse [137, 282, 446, 695, 857]. Hauptmoduln [914]. Hausdorff [190, 476, 567, 955, 1108]. having [29, 106, 566, 726, 732]. Hayes [550]. Hecke [15, 32, 88, 104, 120, 266, 347, 834, 911, 913, 921, 1075]. Heegner [64, 1087]. heights [177, 720, 793]. Heilbronn [160]. Heisenberg [1046]. Helleseth [586]. heptagonal [936]. Herbrand [916]. Hermitian [474, 508, 622]. Heron [173, 741]. Heyman [635]. Hida [948]. high [593]. high-weight [593]. Higher [52, 94, 146, 181, 237, 279, 314, 316, 324, 357, 363, 503, 541, 542, 568, 594, 606, 608, 612, 613, 765, 772, 827, 835, 964, 972, 1036, 1044, 1093]. higher-dimensional [568]. Hilbert [108, 185, 246, 262, 275, 425, 529, 702, 782, 784, 805, 824, 829, 990]. Hirzebruch [400]. Hodge [672]. hold [379]. Holomorphic [232, 253, 300, 544, 716, 960, 1057]. holomorphy [780]. homogeneous [783]. homology [396]. Honda [18]. hook [887]. Hopf [202, 871]. Hopf-Galois [871]. Horn [68]. horospheres [425]. House [637]. Hurwitz [268, 368, 602, 691, 763, 1061]. Hybrid [795]. hyperbolas [57, 322, 931]. Hyperbolic [412, 741, 823]. Hyperderivative [552]. hyperderivatives [737]. Hyperelliptic [136, 229, 378, 419, 525, 852]. Hyperfields [144]. hypergeometric [563, 672, 839, 1079]. Hyperkloosterman [830]. hyperplanes [663]. hyperquadratic [222]. hypersurfaces [203, 215]. hypothesis [275, 414, 537, 624, 855].

ideal [43, 171, 175, 233, 359, 538, 1016, 1059]. ideals [550, 822]. idelic [130]. identical [715]. identities [331, 459]. identity [587]. II [14, 69, 116, 130, 284, 326, 338, 496, 520, 541, 616, 767, 788, 857, 957, 966]. III [399, 684]. Iizuka [680]. Ikeda [908]. image [821, 850]. images [317]. imaginary [88, 96, 171, 667, 702, 841, 853, 881, 959, 1024]. implies [158]. Improved [416, 1030]. improvement [115, 206]. increasingly [744]. Indecomposable [162]. indefinite [685]. independence [104, 302, 332, 737, 847, 907, 1008]. independent [564]. index [383, 753]. indices [29, 1045]. indivisible [917, 1117]. induce [114]. Inequalities [30, 215, 849, 938, 1093]. inequality [5, 102, 113, 153, 423, 463, 953, 1001]. Infinite [28, 854, 927]. Infinitely [158, 303, 965, 978]. inhomogeneous [407, 955]. injectivity [705]. Inner [83, 434, 749]. insolvability [722]. Integer [22, 228, 283, 319, 381, 584, 631, 748, 808, 882, 924, 949, 997, 1027, 1078]. integer-valued [584]. integers [1, 2, 90, 106, 109, 154, 162, 168, 219, 269, 295, 338, 354, 379, 522, 571, 632, 637, 692, 711, 726, 871, 893, 932, 1002, 1031, 1055, 1082]. Integral [42, 207, 263, 265, 323, 337, 477, 566, 587, 629, 691, 698, 740, 829, 831, 851, 894, 1000, 1085, 1092]. integrality [1038]. integrals [134, 712, 1116]. integrated [287, 991]. integration [720]. interest [899]. intermediate [526]. interpolated [687, 746]. interpretation [334, 554]. Intersection [34, 78, 130, 310, 922]. Intersections [427, 510, 610, 879, 1098]. intervals [225, 486, 627, 1006, 1012]. Introducing [135]. invariant [170, 402, 446, 646, 693, 846, 898, 1083]. Invariants [240, 282, 349, 386, 595, 788, 825, 881, 947, 959, 1024, 1029, 1121]. Inverse [19, 315, 390, 440]. investigation [464]. involution [761, 903]. involving [8, 55, 82, 87, 104, 128, 372, 894]. irreducibility [990]. Irreducible [910]. irregular [28, 895]. irregularity [250]. isocrystals [652, 657, 658]. Isogeny [46, 598, 854]. isogeny-torsion [854]. isolated [17]. isometries [970]. isomorphisms [342]. iterate [1094]. iterated [84, 712]. iterates [420, 1072]. Iteration [204]. itself [25]. IV [62, 542]. Iwasawa [386, 435, 461, 718].

J [90, 113, 184, 269, 310, 372, 373, 452, 538, 602, 758, 817, 818, 888, 889, 959, 960, 991, 1088, 1104]. Jacobi [207, 226, 242, 307, 636, 757, 831, 846, 877, 906, 954]. Jacobian [437]. Jacobians [378, 525]. Jacquet [834]. January [9, 272, 789, 992]. Jessica [560]. jet [982]. JNT [506, 540]. Joint [438, 450]. Jordan [77, 611]. July [140, 393, 506, 618, 900, 1089]. June [117, 374, 603, 890, 1067]. Jutila [810].

K3 [621, 802, 970, 1103]. Kac [106, 708]. Kaneko [18]. Karatsuba [160]. Kato [105]. Katok [608]. Kedlaya [531]. kernel [278, 380, 679, 1064]. kernels [453]. key [587]. Khintchine [221, 663]. Kida [364]. killed [1096]. Klein [361, 1066]. Klingen [1111]. Kloosterman [322, 794]. Koecher [908]. Kohnen [122]. Kolyvagin [500]. Kronecker [412]. Kummer [461]. Kurokawa [905, 925].

L [367]. L-function [367]. l.c.m [172]. Lagrange [152]. Laguerre [776]. lambda [240]. Landau [574, 958]. Lang [34]. Langlands [340, 560, 666].

Laplace [712]. Laplace-eigenvalue [712]. Laplacians [56]. Large [5, 112, 113, 190, 402, 417, 420, 441, 469, 615, 719, 932, 999, 1002]. largest [23]. Last [416, 704]. Lattès [699]. Lattice [157, 466, 474, 613, 744, 772, 773, 791, 949, 970]. lattices [43, 62, 370, 473, 577, 913, 1114]. Laurent [385]. law [91]. laws [181]. layered [258]. LCM [638]. least [214, 571]. Lebesgue [220]. Legendre [667]. Lehmer [644]. lemma [834]. length [712, 727, 1020]. lengths [887]. Lerch [249]. level [345, 347, 361, 472, 542, 543, 636, 656, 736, 766, 824, 831, 956, 1057, 1063]. levels [95, 189]. Lewin [690]. Li [969]. Lie [19, 191, 659, 1083]. lift [908, 1009]. lifting [493]. lifts [61, 469, 880, 905, 922, 925, 1083]. like [254, 433, 476]. limit [412, 571, 634]. limits [590]. Lindemann [982]. line [30, 335, 474, 576, 634, 652, 715, 978]. Linear [31, 80, 268, 285, 302, 333, 524, 569, 602, 717, 847, 907, 987]. linearized [987, 1077]. linearly [564]. Liouville [165]. Littlewood [377, 927]. Ljunggren [220]. LLL [860]. Local [11, 99, 155, 169, 181, 324, 340, 386, 421, 434, 461, 559, 560, 572, 606, 616, 617, 662, 724, 751, 797, 836, 946, 1071, 1121]. local-global [616, 724, 797]. localization [813]. locally [336, 1038]. Log [473, 953, 1078, 1119]. Log-concavity [953, 1078]. Log-free [1119]. log-unit [473]. logarithm [868]. Logarithmic [240, 563]. Logarithmic-type [563]. Logarithms [92, 501]. logic [1048]. Long [589]. Long-range [589]. Longest [369]. loops [927]. Low [62, 145, 151, 1084]. low-discrepancy [1084]. low-lying [151]. Lower [44, 123, 188, 347, 363, 416, 591, 704, 754, 1116]. Lübeck [19]. Lubin [40, 209]. lying [151].

Maass [61, 103, 908, 911, 912]. MacMahon [599]. Mahler [81, 94, 104, 283, 404, 996]. Main [435, 916, 948, 964]. manifolds [738]. Manin [679, 714, 1000]. many [35, 158, 303, 744, 965, 978]. mapping [25]. maps [84, 501, 699, 705, 747, 756, 842, 1094]. March [49, 311, 843, 1014]. Markoff [119, 323, 451, 1000]. Markoff-type [1000]. Markov [166]. Martin [138]. Mass [1044]. Massey [1054]. matrices [265, 508, 552, 650, 803, 1005]. matrix [237, 265, 831, 1103]. max [1097]. maximal [46, 365, 390, 398, 792, 866, 1028]. Maximum [327, 480]. May [97, 350, 580, 874, 1050]. Mazur [18, 73]. Mean [136, 192, 251, 447, 638, 689, 734, 765, 773, 848, 892, 1006, 1017, 1062, 1118]. mean-value [765]. means [100]. measure [81, 154, 404, 996]. measures [94]. mediants [675]. Medvedovsky [531]. meet [262]. memorial [540]. Meromorphic [716, 779, 960]. meromorphy [254]. Mertens [594]. Metacommutation [24]. metaplectic [224, 293, 770, 1009, 1071]. method [54, 810, 813, 1006, 1099]. metric [221, 407]. Metrical [167]. Meyer [983]. middle [280]. middle-third [280]. minima [937]. Minimal [58, 200, 277, 359, 390, 392, 538, 664, 893]. Minkowski [135]. minus [266, 461]. missing [354, 725]. mixed [363]. Möbius [148, 149, 892, 1046]. mock [430, 483]. Mod [470, 472, 1076]. Mod- [1076]. model [341, 969]. models [31, 80, 165, 836, 1121]. Moderate [281]. Modular [57, 78, 103, 185, 189, 197, 245, 246, 271, 275, 291, 303, 307, 310, 322, 334, 337, 357, 367, 396, 425, 427, 438, 469, 472, 496, 497, 500, 510, 520, 529, 541– 544, 575, 593, 599, 643, 646, 647, 678, 685, 712, 716, 727, 779, 782, 784, 811, 829, 838, 884, 907, 913, 920, 931, 947, 960, 972, 990, 1008, 1053, 1057]. modularity [96, 108, 470]. module [645, 648, 713, 809]. modules [143, 169, 235, 285, 441, 545–549, 564, 646, 650, 674, 718, 883, 1022, 1047, 1091]. moduli [5, 113, 142, 176, 255, 615, 840]. modulo [59, 214, 381, 589, 735, 1031, 1079]. moment [7, 20, 39, 287, 826, 991, 1104, 1112]. Moments [88, 146, 279, 363, 376, 475, 613, 670, 815, 835, 878, 963, 1036, 1059]. monochromatic [522]. monodromie [738]. monodromy [11, 738, 1005]. monogeneity [477]. Monogenic [304, 468, 818]. monogenity [527, 728]. monoids [745]. Monomials [126, 309, 989, 1021]. Monotonicity [86]. Mordell [34, 678, 1042]. Morse [1048]. motivated [1095]. motives [413, 536, 551]. motivic [191, 554, 696]. moving [215]. MSTD [612]. mu [240]. multi [58, 555, 718, 1084]. multi-base [58]. multi-dimensional [1084]. multi-poly-Bernoulli [555]. multi-variable [718]. Multidimensional [297]. multilinear [865]. multinorm [1056]. multinorm-one [1056]. multipartition [418, 492]. Multiple [198, 223, 433, 459, 523, 567, 571, 683, 687, 696, 715, 746, 747, 895, 986]. multiples [139, 856]. multiplication [46, 91, 295, 552, 802]. Multiplicative [59, 148, 180, 225, 297, 397, 592, 638, 879, 1058, 1113]. multiplicatively [632]. multiplicativity [745]. Multiplicities [508, 669]. multiplicity [417]. multiplier [829]. multivariable [638, 996]. multizeta [554]. Mumford [714].

Nagao [229, 798]. Nagell [220]. Nathanson [277, 1099]. natural [133, 600]. near [297, 426, 529, 634]. near-primitive [426]. Nearly [300, 964]. need [617]. nef [263]. negative [1045]. Néron [111]. newform [380, 419, 943]. newforms [649, 824, 859, 905, 921, 1010]. Newton [200, 864]. Niho [586]. nilmanifold [1046]. Noetherian [671]. Non [14, 75, 81, 122, 158, 178, 250, 294, 326, 406, 468, 477, 516, 529, 559, 670, 673, 681, 728, 778, 785, 832, 841, 948, 965, 982, 997, 1007, 1016]. non-abelian [673]. non-Archimedean [982]. Non-CM [965]. non-coincident [294]. Non-commutative [516]. non-compact [785]. non-congruent [681]. non-critical [529]. non-Euclidean [841]. Non-Galois [14, 326]. non-generic [559]. non-geometric [832]. non-integer [997]. Non-monogenic [468]. non-normal [670]. non-ordinary [778]. Non-Pólya [1016]. Non-real [250]. non-split [122]. non-squarefree [477]. non-tempered [81]. non-triviality [948]. Non-vanishing [75, 178, 406, 1007]. noncongruence [753]. noncorrelated [387]. nondegenerate [508]. nonnegative [319]. Norm [137, 606, 857, 954, 1041]. normal [233, 282, 446, 561, 566, 670, 871, 895]. normality [135]. normalization [935]. normalized [913]. norms [13, 402, 969, 1028]. note [68, 80, 136, 148, 296, 411, 658, 710, 778, 821, 852, 934, 1077, 1098]. Notes [657].

November [230, 456, 729, 961]. Number [2, 14, 84, 88, 90, 96, 113, 145, 184, 199, 218, 241, 269, 310, 321, 326–328, 342, 372, 373, 382, 389, 446, 452, 471, 477, 497, 513, 526, 538, 561, 591, 602, 626, 638, 641, 659, 670, 689, 710, 711, 728, 758, 787, 817, 818, 882, 888, 889, 899, 906, 921, 937, 945, 959, 960, 991, 1020, 1032, 1040, 1052, 1072, 1073, 1081, 1087, 1088, 1104]. number-theoretic [899]. numbers [36, 37, 53, 78, 82, 123, 133, 166, 188, 194, 208, 238, 258, 296, 310, 321, 383, 397, 436, 437, 439, 444, 451, 455, 460, 519, 521, 525, 555, 598, 600, 669, 681, 684, 734, 777, 824, 839, 895, 936, 956, 970, 1019, 1056, 1069, 1102, 1113]. Numerical [348, 377, 958]. Numerically [1102].

obstruction [1000]. Obstructions [165, 320]. octic [108, 1005]. October [211, 442, 706, 950]. Odd [228, 365, 401, 418, 492, 596, 681, 736, 863]. odd-regular [401]. often [631]. Ohno [687]. Ohno-type [687]. oldforms [649]. One [94, 96, 335, 345, 347, 360, 617, 640, 673, 766, 797, 857, 1005, 1041, 1047, 1056, 1063]. One-level [347, 766, 1063]. one-parameter [1005, 1047]. only [781, 932]. Ono [953]. open [821]. operators [103, 268, 317, 602, 656, 939, 1028]. Oppenheim [286]. Optimal [193, 431, 731, 1011]. Optimality [849]. Optimizing [1039]. orbits [34, 306, 845, 870, 1018, 1047, 1108]. Order [261, 347, 383, 503, 564, 1017, 1093, 1112]. orders [59, 431, 455, 636, 827, 973, 1011]. ordinary [65, 778, 852, 964]. orthogonal [19, 684, 685, 732, 761, 947]. orthogonality [700]. oscillation [236]. other [958]. overconvergence [1086]. overconvergent [652, 658]. overpartitions [86, 227, 495, 504].

P. [465]. packings [38, 208]. Padé [368, 839]. Pages [9, 26, 49, 71, 97, 117, 140, 163, 186, 211, 230, 247, 272, 289, 311, 329, 350, 374, 393, 409, 428, 442, 456, 478, 517, 557, 580, 603, 618, 660, 706, 729, 759, 789, 819, 843, 861, 874, 890, 900, 918, 940, 950, 961, 974, 992, 1003, 1014, 1033, 1050, 1067, 1089, 1105]. pair [52, 447, 454, 623, 817, 1036, 1084]. Pairing [578, 679]. Pairs [114, 349, 392, 578, 1025]. palindromes [369]. Pan [635]. parabolic [412, 622]. parameter [1005, 1047]. parametric [839]. Paramodular [95, 565]. Parity [227, 338, 483, 513, 609, 957]. part [14, 496, 894]. Partial [782, 980]. partially [397]. Participants [506]. partition [132, 418, 430, 599, 764, 795, 938, 953, 1029, 1049, 1093]. partitions [228, 244, 424, 599, 611, 813, 917, 930, 957, 985, 1078, 1117]. Parts [131, 228, 691, 775, 917, 1117]. pattern [387]. patterns [165]. paucity [610, 808]. PCF [694]. Pell [37, 432, 1085]. Perfect [62, 82, 210, 596, 639, 1114]. Period [149, 275, 448, 585, 737]. periodic [699, 701, 842, 998]. Periods [201, 276, 285, 343, 400, 547, 548, 569, 672, 770, 825, 907]. Petersson [7, 1104]. Pezzo [205, 640]. phenomena [237, 838]. phenomenon [274, 610]. physical [860]. Piatetski [127, 358, 931, 1035]. Picard [982]. Pillai [366, 710]. Pisot [534]. place [150]. Plancherel [508]. plane [182, 813]. Plausible [48]. plus [461]. plus/minus [461]. Pohst [463]. Poincaré [406]. point [65, 75, 466, 613, 731, 859, 942, 1005]. Points [4, 22, 35, 64, 139, 157, 203, 263, 265, 303, 323, 327, 337, 356, 381, 405, 425, 441, 452, 461, 481, 500, 521, 564, 699, 772, 791, 842, 1000, 1018, 1103]. Poissonian [52, 503, 623, 827]. polarized [365]. poles [189, 250, 836]. poly [555]. Pólya [45, 153, 326, 414, 458, 498, 1016]. polygon [864]. polygonal [321]. polygons [200]. polylogarithm [198]. Polynomial [84, 194, 264, 381, 404, 501, 527, 570, 633, 691, 803, 976, 1043, 1062, 1082, 1085]. polynomials [94, 107, 131, 185, 204, 213, 246, 275, 358, 384, 423, 480, 682, 695, 719, 732, 739, 748, 751, 776, 781, 783, 786, 806, 816, 865, 880, 889, 910, 968, 987, 996, 1021, 1025, 1058, 1077, 1115]. popular [199]. position [485]. positive [200, 278, 492, 496, 528, 663]. possible [416, 704]. Potentially [336, 469]. pour [106]. power [5, 113, 178, 385, 417, 524, 552, 631, 700, 835, 845, 922]. powerful [1113]. Powers [55, 210, 296, 548, 639, 655, 676, 735, 771, 897]. Practical [36]. Prasad [559]. Preface [507]. preimages [1072]. premiers [106]. prescribed [218, 283, 325, 598, 984]. primary [781]. Prime [6, 23, 106, 107, 152, 159, 175, 178, 183, 225, 261, 297, 360, 365, 447, 561, 600, 677, 681, 689, 711, 758, 762, 870, 956, 965, 987, 1010, 1032, 1040, 1073]. prime-order [261]. Primes [23, 24, 28, 69, 126, 127, 129, 138, 158, 179, 195, 243, 338, 411, 422, 486, 522, 533, 592, 594, 610, 735, 778, 781, 787, 828, 867, 873, 885, 897, 980, 1035, 1120]. Primitive [69, 89, 170, 214, 299, 308, 425, 426, 870, 1112]. Primitively [421, 814]. principal [415, 450, 867]. principle [137, 324, 384, 616, 662, 797, 857, 946]. pro [526, 1081]. pro- [526, 1081]. Probabilistic [341, 768]. Problem [19, 34, 55, 125, 251, 277, 315, 338, 366, 453, 471, 515, 562, 808, 924, 971, 1045, 1097]. problems [283, 390, 755, 922]. Product [276, 363, 389, 398, 438, 671, 673, 886, 1025, 1046, 1062, 1118]. product-one [673]. products [134, 195, 244, 301, 412, 597, 698, 933, 976, 1054]. progress [165]. progression [182, 210, 601]. progressions [369, 445, 460]. projections [597]. projective [56, 652, 783]. Proof [192, 355, 492, 653]. properties [86, 132, 258, 388, 391, 487, 629, 721, 756, 809, 943, 1082]. property [16, 174, 184, 803, 876]. proportions [921]. prosolvable [616]. Pseudo [491, 781, 913]. Pseudo-normalized [913]. pseudo-polynomials [781]. Pseudo-rigid [491]. pullbacks [925]. pure [304, 477, 818]. Purely [709]. Pythagorean [578].

Quadratic [20, 43, 84, 88, 96, 115, 162, 174, 179, 182, 208, 222, 345, 396, 401, 403, 421, 422, 444, 451, 455, 484, 519, 577, 601, 610, 664, 667, 676, 694, 702, 756, 769, 814, 816, 841, 853, 856, 876, 881, 883, 892, 898, 904, 910, 932, 959, 967, 973, 1016, 1024, 1027, 1035, 1116]. quadratics [911]. quadrics [979]. quadrilateral [576]. quantities [743, 958]. Quartic [99, 405, 475, 755, 885, 1065, 1080, 1085]. Quasi [434, 779, 997, 1008]. quasi-ergodic [997]. Quasi-inner [434]. quasi-modular [779, 1008]. quaternary [814]. quaternion [61, 431, 1011]. quaternionic [271, 367]. questions [498]. quinary [876]. quotient [245].

Quotients [69, 77, 361, 504, 609, 815, 939, 957, 980, 1066, 1070, 1075].

Rademacher [244]. Rallis [834]. Ramanujan [102, 227, 252, 361, 644, 785, 956, 1039, 1066, 1087]. Ramanujan-style [956]. Ramanujan-type [1087]. Ramification [65, 315, 526]. ramified [360, 417, 713, 1081]. Ramírez [869]. random [172, 398, 773, 1082, 1108, 1115]. range [589, 1009]. Rank [229, 335, 357, 502, 541, 542, 548, 577, 645, 646, 648, 665, 685, 702, 769, 885, 964, 972, 1013, 1044, 1078]. rank-2 [1013]. Rankin [74, 197, 686, 948, 971, 1053]. Ranks [86, 341, 378, 430, 678, 815, 1060]. Rarefied [1048]. rates [1086]. ratio [385]. Rational [3, 4, 203, 287, 299, 328, 349, 405, 422, 425, 452, 521, 693, 799, 800, 842, 888, 898, 926, 986, 991, 1094]. rationality [533, 904, 1042]. rationally [1052]. rationals [895]. Ratios [620]. ray [530, 822]. rayons [530]. rd [1041]. Re [294]. Real [30, 43, 46, 85, 109, 136, 139, 162, 250, 389, 396, 423, 458, 519, 577, 715, 806, 856, 883, 895, 916, 1020, 1061]. real-rooted [423]. Realizing [200]. reals [41]. reciprocal [1021]. reciprocals [691]. reciprocity [181, 1060]. Recurrence [132, 172, 218, 665]. recurrences [288]. Recursively [1019]. reduced [460]. Reducible [107, 470, 595]. reduction [105, 209, 461, 521, 525, 564, 645, 709, 784, 852, 898]. Reductions [1, 90, 143, 256, 383, 546]. reductive [320, 745, 821]. refined [493]. refinements [353]. reflexive [682]. Reformulating [927]. regarding [332]. region [395, 807]. Regular [196, 401, 494, 769, 833, 957, 1094]. regularity [174, 876]. related [251, 279, 534, 755, 872, 958, 987, 1001, 1086, 1097, 1108]. relating [331, 455]. relation [6, 93, 440, 687, 758]. Relations [397, 404, 523, 554, 746, 930, 1109]. Relative [45, 512, 654]. remainder [368]. Remarks [105, 686]. Renormalization [134, 768]. Renyi [476]. Renyi-like [476]. repeated [321]. repeller [1065]. Report [353]. representability [981]. Representation [58, 76, 123, 179, 208, 284, 397, 522, 648, 671, 774, 788, 947]. representations [19, 85, 112, 134, 174, 333, 336, 360, 402, 470, 493, 514, 587, 595, 664, 666, 749, 761, 771, 821, 825, 850, 1038, 1060, 1076]. representative [236]. represented [358, 1027]. Representing [601, 856, 932]. Reprint [546]. repulsion [786]. repunits [371]. resembling [148]. residually [595]. Residue [460, 606, 1091, 1097, 1117]. residues [218, 484, 697, 816]. resolution [745, 1101]. resonance [878]. respect [836]. restricted [244, 301, 346, 526, 560, 922, 953, 1012, 1026, 1100]. restriction [762]. result [44, 407, 422, 575, 653]. resultantal [674]. results [198, 213, 495, 568, 656]. revisited [830]. Riemann [48, 124, 157, 275, 332, 414, 420, 489, 537, 607, 765, 855]. Riesz [100]. Rigid [11, 491, 666]. Rikuna [527]. ring [53, 645, 871, 1097]. rings [95, 143, 207, 237, 281, 482, 546, 650, 838, 841, 966]. Rogers [227, 734]. Romik [226]. Root [69, 214, 437, 534, 669, 786, 824, 882]. Rooted [423, 674, 747]. Roots [244, 426, 719, 750, 806, 968, 980]. rotational [777]. Roth [1035]. Roth-type [1035]. rough [1102]. rounded [43]. RSA [370]. Rubin [73]. Ruelle [823]. Rülling [653]. Run [1020]. Run-length [1020]. running [115]. Ruzsa [983].

S [624]. Saito [905, 925]. Salem [238, 970]. same [774, 937]. sampled [571]. Sárközy [129, 484, 838]. Sarnak [192, 608]. Satake [291]. satisfy [1073]. satisfying [397]. Sato [511, 798, 1036, 1103]. sauvage [738]. savings [13]. scaffolds [33, 713]. scheme [671, 845]. schemes [449, 491, 833, 1096]. Schinzel [69, 297]. Schinzel-type [69, 297]. Schmidt [599, 783, 896]. Schreier [200]. second [20, 416, 704]. Secondary [963]. sections [474]. sector [726]. sectors [486]. Selberg [74, 450, 536, 634, 686, 700, 948, 971, 1006]. selectivity [431, 1011]. self [1094]. self-maps [1094]. Selmer [267, 364, 496, 669, 1060]. semi [215, 521]. semi-abelian [521]. semi-decomposable [215]. semiabelian [1094]. semigroups [910]. semistable [33, 525]. September [186, 428, 660, 940]. sequence [29, 257, 371, 499, 1045]. Sequences [57, 89, 132, 170, 172, 180, 182, 213, 358, 384, 387, 432, 502, 532, 589, 623, 673, 675, 788, 931, 944, 981, 1036, 1084]. Series [16, 22, 103, 104, 120, 134, 184, 223, 309, 314, 331, 342, 354, 382, 385, 406, 412, 415, 417, 447, 488, 524, 585, 597, 683, 686, 701, 736, 815, 851, 906, 908, 952, 967, 986, 1079, 1087, 1088, 1109]. Serre [96]. set [206, 280, 283, 308, 781, 828]. Sets [60, 125, 129, 133, 154, 262, 398, 426, 462, 612, 664, 668, 717, 774, 788, 893, 983, 1027, 1073, 1099, 1108]. setup [1110]. seven [210, 531, 977]. several [1101]. Severi [111]. sextic [328, 405, 477]. Shafarevich [471]. Shapiro [127, 358, 931, 1035]. Sharifi [362]. shears [51]. Sheats [739, 889]. shells [466]. Shifted [23, 129, 515, 592, 808, 810, 1064, 1095, 1107, 1110, 1118]. Shimura [317, 863, 995, 1098]. Shintani [1065]. Shioda [1013]. Short [21, 159, 279, 486, 627, 697, 744, 1006, 1012]. short-sum [697]. shortest [453]. Shparlinski [635]. shuffle [523, 554]. shuffle-type [523]. sides [560]. Sidon [668, 717, 1099]. Siegel [48, 66, 161, 216, 235, 307, 734, 736, 952, 954, 958, 967, 1088]. Sierpiński [723]. sieve [5, 113, 615, 1001]. sigma [929]. Sign [165, 614]. signature [583, 600, 945, 947, 967]. Signatures [318]. signs [160]. simple [19, 24, 54, 202, 365, 408]. simplicial [972]. Simply [895]. Simultaneous [179, 722, 842, 949, 1007]. single [206]. singular [142, 447, 495, 504, 840]. singularities [65]. site [150]. sixth [826]. Skabelund [792]. Skałba [869]. Skolem [399, 1045]. slope [389]. slopes [474]. Slow [101, 288, 757]. Small [126, 261, 338, 370, 381, 486, 688, 837, 894, 1084]. small-exponent [370]. smallest [775]. Smallness [793]. Smarandache [264]. smooth [182, 295]. Solution [750]. solutions [319, 416, 626, 629, 704, 710, 752]. Solving [786, 1045]. Some [19, 59, 108, 198, 258, 298, 301, 338, 353, 458, 459, 465, 466, 498, 535, 568, 667, 686, 778, 899, 1103, 1112]. sommes [21]. Sophie [365, 533]. space [61, 92, 508, 544, 757, 969, 1057]. spaces [56, 266]. spacing [274]. spanning [866]. Sparse [781, 786, 1073]. Special [147, 178, 466, 547, 549, 672, 701, 833, 899, 1023, 1091]. specialization [705]. Specializations [1013]. specified [812]. spectra [63]. Spectral [56, 855, 1023]. Spectrum [476, 524, 935, 1113]. speed [305]. sphere [38]. spheres [56, 146, 926]. spin [6, 415, 758, 903]. spinor [431, 1011]. split [122, 178, 836]. splitting [867, 1121]. spoof [596]. sporadic [408]. Square [121, 245, 615, 633, 682, 689, 835, 848, 878, 980, 1017, 1032, 1115]. Square-free [245, 633, 1032, 1115]. Square-reflexive [682]. squarefree [477, 636, 1002]. squares [174, 280, 522, 628, 856, 892, 1012, 1026, 1069]. stability [659, 994]. stable [223]. standard [285, 367]. Stark [902]. statistic [764]. statistics [1029, 1078]. Stein [73]. Steinberg [396, 883]. Stern [675]. Stieltjes [1037]. stochastic [1084]. strategy [899]. stretched [791]. string [1095]. strip [1070]. Strong [174, 267, 876]. stronger [110]. strongly [62]. structure [119, 232, 336, 713, 774, 809]. structures [126, 202, 656, 879]. study [179, 691, 860]. Sturm [647]. Sturm-type [647]. Sturmian [201, 944]. style [956]. Subconvexity [13, 587, 742, 795]. subcovers [792]. subgeneral [485]. subgroup [516, 598, 812]. subgroups [225, 241, 450, 520, 638, 753, 883, 1075]. sublattices [773]. submodules [718]. subrings [591, 780]. subschemes [485]. subsequence [257]. Subset [924, 966]. subset-sum [924]. Subspace [215, 485, 783]. successive [937]. sufficiently [35]. Sum [8, 69, 128, 257, 280, 372, 391, 515, 628, 639, 688, 697, 788, 894, 911, 924, 1012, 1032, 1080, 1097, 1110]. sum-of-divisors [69]. sum-of-squares [1012]. Sums [21, 60, 156, 190, 210, 225, 252, 257, 279, 298, 313, 316, 322, 344, 380, 433, 455, 522, 535, 552, 553, 561, 573, 586, 627, 676, 695, 725, 750, 767, 794, 801, 810, 830, 856, 864, 878, 909, 931, 936, 943, 966, 1026, 1028, 1048, 1064, 1069, 1095, 1107]. sumsets [154, 292, 725]. Sun [352, 353, 1026]. sup [13]. sup-norms [13]. Super [17]. Super-isolated [17]. Supercuspidal [749, 761]. supercuspidals [1111]. supernormal [764]. supersingular [461, 493]. supplements [390]. surface [205, 438, 970]. surfaces [65, 111, 130, 157, 246, 323, 335, 425, 449, 482, 621, 640, 705, 802, 823, 882, 1000, 1103]. Surjectivity [648]. survey [509, 511]. Swinnerton [67, 872]. Swinnerton-Dyer [67, 872]. symbol [175, 606, 805]. symbols [920]. symmetric [121, 349, 529, 637, 700, 771, 835]. symplectic [347, 850]. system [1035]. systems [11, 105, 167, 286, 626, 678, 829]. Systolic [727].

table [295]. taken [23]. Tamagawa [525, 598, 684]. tame [453, 805]. tamely [1081]. Tate [18, 40, 209, 282, 446, 471, 536, 798, 1036, 1103]. tau [644]. tau-function [644]. taxonomy [38]. Taylor [242]. Tempered [81, 1114]. Tensor [169, 548]. term [210, 217, 241, 447, 689, 751, 848]. terms [29, 172, 347, 927, 963, 981, 1042, 1084]. Ternary [196, 401, 601]. Test [74, 115, 731]. th [175, 534, 570, 922]. their [29, 104, 132, 154, 173, 294, 600, 678, 702, 726, 779, 881, 931, 959, 1010]. themes [511]. Theor [758]. theorem [32, 106, 122, 129, 206, 215, 221, 235, 236, 286, 321, 344, 416, 440, 450, 485, 486, 493, 516, 528, 571, 634, 663, 689, 697, 704, 708, 734, 765, 783, 787, 838, 982, 983, 1035, 1040, 1073]. Théorème [106]. theorems [59, 69, 167, 183, 297, 490, 588, 896]. theoretic [179, 721, 899]. Theory [47, 90, 113, 130, 184, 269, 310, 372, 373, 407, 435, 451, 452, 461, 538, 602, 643, 805, 817, 818, 888, 889, 959, 960, 991, 1048, 1088, 1092, 1095, 1104]. there [158]. thereof [691]. Theta [48, 92, 226, 430, 483, 514, 736, 829, 933, 967, 985, 1009]. Theta-type [985]. third [280]. Thompson [408]. Three [30, 379, 639, 685, 712, 840, 887, 989, 1101]. threefolds [108, 654, 1005]. Thue [1048]. time [115]. Tits [1075]. topological [193, 1029]. topology [335]. tori [855, 857, 923, 1056]. Torsion [3, 40, 328, 441, 491, 564, 598, 671, 693, 709, 812, 854, 888, 1052]. Torsions [884]. torsors [616]. torus [1041]. total [726, 838]. totally [85, 178, 389, 713, 1065]. totient [69, 77, 82, 87, 254, 611]. towers [334, 659]. trace [296]. traces [142, 724, 912, 965]. transcendental [25, 283]. transfer [103]. Transformation [42, 149]. transforms [734]. translates [106]. tree [747, 1075]. trees [674]. triangle [785]. triangles [173, 349, 356, 741, 817]. triangular [196, 727, 864, 1069]. trigonometric [743]. trinomials [728]. Triple [313, 389, 886, 909]. triples [114, 621]. Triquadratic [800]. triviality [948]. truncated [144]. tuple [571]. tuples [297, 632]. Turán [414, 1093]. twice [7, 1104]. Twin [69, 183, 873]. Twin-prime [183]. Twisted [323, 363, 553, 655, 823, 908, 978, 986]. Twisting [255]. Twists [73, 79, 178, 403, 444, 669, 766, 885, 892, 954, 976]. Two [34, 195, 307, 331, 345, 346, 371, 521, 628, 849, 864, 887, 902, 921, 979]. two-division [521]. type [19, 69, 220, 227, 297, 345, 361, 407, 523, 563, 599, 617, 647, 687, 710, 785, 791, 858, 908, 953, 985, 1000, 1035, 1054, 1066, 1087, 1103]. Type-I [345]. types [749]. Typically [693].

u042d [487]. ubiquity [956]. Ultraproduct [509]. unbounded [753]. unconditional [115, 492]. Undecidable [388, 629]. Underapproximation [796]. Unexpected [917]. unified [899]. Uniform [306, 343, 630, 1100]. uniformly [571]. unimodal [502]. Unique [573, 817, 915]. uniqueness [16, 184]. unit [438, 473, 637, 1097]. unitary [82, 415, 514, 863]. units [14, 902, 947]. unity [244, 719, 750]. univariate [131]. Universal [421, 577, 814, 929, 936, 1069, 1116]. universality [450, 664]. Univoque [895]. Unlikely [879, 922, 1098]. unnormalized [684]. Unramified [145, 514, 568, 1041]. upper [92, 175]. using [916, 1053].

V [357]. Vaaler [407]. valuation [481, 721]. Value [15, 136, 734, 765, 773, 1006]. Value-distribution [15]. valued [144, 406, 584, 748, 821, 928]. Values [23, 44, 104, 107, 120, 121, 131, 178, 190, 194, 198, 249, 251, 271, 284, 299, 302, 418, 420, 432, 433, 449, 459, 475, 529, 547, 549, 553, 554, 579, 582, 633, 638, 662, 672, 686, 687, 696, 701, 703, 716, 719, 746, 747, 833, 847, 943, 948, 960, 999, 1007, 1008, 1017, 1023, 1058, 1109, 1112, 1115, 1118]. Vandermonde [552]. Vanishing [75, 133, 178, 205, 406, 653, 731, 859, 1007]. variable [94, 718]. variables [152, 863, 864, 1101]. Variance [466, 627, 1012, 1049]. variant [235, 366, 1040]. Variants [221]. Variation [525]. Variations [352, 414, 511, 644]. Variétés [738]. varieties [4, 17, 40, 265, 276, 291, 343, 365, 437, 452, 674, 678, 709, 783, 793, 863, 984, 1076, 1092, 1098]. variety [1094]. vector [406, 453, 928]. vector-valued [406, 928]. vectors [74, 666, 744, 866, 1111]. Venkatesh [348]. verification [348, 377]. version [67, 528, 708, 783]. versions [464]. vertical [640]. very [35, 786]. VI [972]. via [48, 54, 229, 412, 448, 587, 805, 925, 983, 1048]. Vinogradov [153, 486]. violating [797]. visibility [772, 949]. Vologodsky [720]. volume [540].

Waldschmidt [35]. walks [101]. Waring [125]. ways [48]. weak [320, 757, 853, 912]. weakly [544, 716, 960, 1057]. Weber [471]. Wei [352, 353]. Weierstrass [81, 929]. weight [39, 58, 469, 497, 593, 698, 829, 831]. Weighted [177, 436, 675, 708, 1110]. weights [402, 608]. Weil [67, 216, 560, 586, 678, 947]. Weil-étale [67]. Weil-restricted [560]. well [43]. well-rounded [43]. Weyl [54, 190, 223, 714, 846, 986, 1028, 1080]. which [37, 219, 702, 1108]. Whittaker [31, 80, 1009]. whose [1025]. Wieferich [158]. wild [278, 738]. Wildly [417]. Wiles [1060]. Wintner [490]. without [1052]. words [201, 532, 625].

XIII [599].

Yau [108, 654, 1005]. Yui [988].

Zagier [122, 400, 455]. Zariski [845]. Zeckendorf [2, 269]. zero [78, 257, 274, 310, 324, 395, 520, 778, 788, 823, 958, 1097, 1111, 1119]. zero-cycles [324]. zerofree [395]. zero-spacing [274]. zero-sum [257, 788, 1097]. Zeroes [116, 161, 271]. Zeros [151, 217, 294, 382, 404, 562, 576, 585, 607, 807, 865, 978, 1061, 1063]. Zeta [42, 48, 56, 94, 99, 124, 198, 217, 249, 254, 268, 368, 395, 420, 433, 449, 450, 459, 489, 523, 536, 553, 576, 602, 607, 671, 687, 691, 696, 697, 715, 746, 763, 765, 823, 855, 929, 973, 1023, 1061, 1071]. zeta-function [249, 395, 420]. zeta-functions [523, 715]. Zhi [352, 353]. Zhi-Wei [352, 353]. Zumkeller [258].

References

Debry:2016:RAI

[1] Christophe Debry and Antonella Perucca. Reductions of algebraic integers. Journal of Number Theory, 167(??):259–283, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

      tronic). URL http://www.sciencedirect.com/science/article/
      pii/S0022314X16300488. See addendum [90].

Chang:2018:ANZ

[2] Sungkon Chang. Average number of Zeckendorf integers. Journal of Number Theory, 186(??):452–472, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

      sciencedirect.com/science/article/pii/S0022314X17304055. See

corrigendum [269].

Daowsud:2018:CFR

[3] Katthaleeya Daowsud and Thomas A. Schmidt. Continued fractions for rational torsion. Journal of Number Theory, 189(??):115–130, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304602. See corrigendum [888].

Salami:2019:RPC

[4] Sajad Salami. The rational points on certain Abelian varieties over function fields. Journal of Number Theory, 195(??):330–337, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301811. See corrigendum [452].

Baier:2019:LSI

[5] Stephan Baier and Rajneesh Kumar Singh. Large sieve inequality with power moduli for function fields. Journal of Number Theory, 196(??): 1–13, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302725. See erratum [113].

McMeekin:2019:APS

[6] Christine McMeekin. On the asymptotics of a prime spin relation. Journal of Number Theory, 200(??):407–426, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300307. See

corrigendum [758].

Sun:2019:DFM

[7] Haiwei Sun and Yangbo Ye. Double first moment for L(\frac{1}{2}, Sym^{2}f × g) by applying Petersson’s formula twice. Journal of Number Theory, 202(??):141–159, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300538. See corrigendum [1104].

Bordelles:2019:SIE

[8] Olivier Bordellès, Lixia Dai, Randell Heyman, Hao Pan, and Igor E. Shparlinski. On a sum involving the Euler function. Journal of Number Theory, 202(??):278–297, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300459. See

corrigendum [372].

Anonymous:2020:PAA

[9] Anonymous. Pages A1–A2, 1–320 (January 2020). Journal of Number Theory, 206(??):A1–, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBa

[10] Anonymous. Editorial Board. Journal of Number Theory, 206(??): ifc, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930304X.

Katz:2020:RLS

[11] Nicholas M. Katz, Antonio Rojas-León, and Pham Huu Tiep. Rigid local systems with monodromy group the Conway group Co_{3}. Journal of Number Theory, 206(??):1–23, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302380.

Arenas:2020:BDA

[12] Manuel Arenas and Luis Arenas-Carmona. Branches on division algebras. Journal of Number Theory, 206(??):24–45, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302082.

Gillman:2020:ESS

[13] Nate Gillman. Explicit subconvexity savings for sup-norms of cusp forms on PGL_{n}(R). Journal of Number Theory, 206(??):46–61, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302094.

Louboutin:2020:NGC

[14] Stéphane R. Louboutin. Non-Galois cubic number fields with exceptional units. Part II. Journal of Number Theory, 206(??):62–80, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302100.

Akbary:2020:VDC

[15] Amir Akbary and Alia Hamieh. Value-distribution of cubic Hecke L-functions. Journal of Number Theory, 206(??):81–122, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19302124.

Dixit:2020:UPG

[16] Anup B. Dixit. A uniqueness property of general Dirichlet series. Journal of Number Theory, 206(??):123–137, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302148. See

corrigendum [184].

Scholl:2020:SIA

[17] Travis Scholl. Super-isolated abelian varieties. Journal of Number Theory, 206(??):138–168, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930215X.

Guerzhoy:2020:HKC

[18] P. Guerzhoy. Honda–Kaneko congruences and the Mazur–Tate padic σ-function. Journal of Number Theory, 206(??):169–181, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302161.

Zenteno:2020:LCR

[19] Adrián Zenteno. Lübeck’s classification of representations of finite simple groups of Lie type and the inverse Galois problem for some orthogonal groups. Journal of Number Theory, 206(??):182–193, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302173.

Sono:2020:SMQ

[20] Keiju Sono. The second moment of quadratic Dirichlet L-functions. Journal of Number Theory, 206(??):194–230, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302185.

Labihi:2020:ECS

[21] O. Labihi and A. Raouj. Estimation de certaines sommes courtes. (French) [Estimation of certain short sums]. Journal of Number Theory, 206(??):231–249, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302197.

Seppala:2020:EFS

[22] Louna Seppälä. Euler’s factorial series at algebraic integer points. Journal of Number Theory, 206(??):250–281, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302203.

Wu:2020:EHC

[23] Jie Wu. Elliott–Halberstam conjecture and values taken by the largest prime factor of shifted primes. Journal of Number Theory, 206(??):282– 295, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302379.

Chari:2020:MPC

[24] Sara Chari. Metacommutation of primes in central simple algebras. Journal of Number Theory, 206(??):296–309, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302392.

Lelis:2020:TEF

[25] Jean Lelis and Diego Marques. On transcendental entire functions mapping Q into itself. Journal of Number Theory, 206(??):310–319, January 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302409.

Anonymous:2020:PF

[26] Anonymous. Pages 1–472 (February 2020). Journal of Number Theory, 207(??):1–472, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBb

[27] Anonymous. Editorial Board. Journal of Number Theory, 207(??):ifc, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303233.

Lee:2020:IFI

[28] Jungyun Lee and Yoonjin Lee. Infinite families of irregular primes in cyclotomic function fields. Journal of Number Theory, 207(??):1–21, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302762.

Kim:2020:DTE

[29] Seoyoung Kim. The density of the terms in an elliptic divisibility sequence having a fixed G.C.D. with their indices. Journal of Number Theory, 207(??):22–41, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302367.

Barnard:2020:TCI

[30] Richard C. Barnard and Stefan Steinerberger. Three convolution inequalities on the real line with connections to additive combinatorics. Journal of Number Theory, 207(??):42–55, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302549.

Sayag:2020:DLM

[31] Eitan Sayag and Mahendra Kumar Verma. On disjointness of linear models and degenerate Whittaker models. Journal of Number Theory, 207(??):56–82, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930143X.

Panda:2020:GTH

[32] Corina B. Panda. A generalization of a theorem of Hecke for SL_{2}(F_{p}) to fundamental discriminants. Journal of Number Theory, 207(??):83– 109, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301490.

Keating:2020:GSS

[33] Kevin Keating. Galois scaffolds and semistable extensions. Journal of Number Theory, 207(??):110–121, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302550.

Rout:2020:DML

[34] Sudhansu Sekhar Rout. The dynamical Mordell–Lang problem for intersection of two orbits. Journal of Number Theory, 207(??):122–137, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302562.

Chang:2020:DCW

[35] Yu-Lin Chang and Shin-Yao Jow. Demailly’s conjecture on Waldschmidt constants for sufficiently many very general points in P^{n}. Journal of Number Theory, 207(??):138–144, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302574.

Leonetti:2020:PNA

[36] Paolo Leonetti and Carlo Sanna. Practical numbers among the binomial coefficients. Journal of Number Theory, 207(??):145–155, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302586.

Ddamulira:2020:CPE

[37] Mahadi Ddamulira and Florian Luca. On the x-coordinates of Pell equations which are k-generalized Fibonacci numbers. Journal of Number Theory, 207(??):156–195, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302598.

Chait-Roth:2020:TCS

[38] Devora Chait-Roth, Alisa Cui, and Zachary Stier. A taxonomy of crystallographic sphere packings. Journal of Number Theory, 207(??):196– 246, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302604.

Frolenkov:2020:CMA

[39] Dmitry Frolenkov. The cubic moment of automorphic L-functions in the weight aspect. Journal of Number Theory, 207(??):247–281, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302628.

Ozeki:2020:TAV

[40] Yoshiyasu Ozeki. Torsion of abelian varieties and Lubin–Tate extensions. Journal of Number Theory, 207(??):282–293, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930263X.

Mehmetaj:2020:CFE

[41] Erblin Mehmetaj. On the r-continued fraction expansions of reals. Journal of Number Theory, 207(??):294–314, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302653.

Espinoza:2020:CZI

[42] Milton Espinoza. On certain zeta integral: Transformation formula. Journal of Number Theory, 207(??):315–348, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302665.

Srinivasan:2020:CCW

[43] Anitha Srinivasan. A complete classification of well-rounded real quadratic ideal lattices. Journal of Number Theory, 207(??):349–355, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302677.

Zhai:2020:LBR

[44] Shuai Zhai. A lower bound result for the central L-values of elliptic curves. Journal of Number Theory, 207(??):356–366, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19302689.

Maarefparvar:2020:RPG

[45] Abbas Maarefparvar and Ali Rajaei. Relative Pólya group and Pólya dihedral extensions of Q. Journal of Number Theory, 207(??):367–384, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302690.

Ionica:2020:IGM

[46] Sorina Ionica and Emmanuel Thomé. Isogeny graphs with maximal real multiplication. Journal of Number Theory, 207(??):385–422, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302410.

Gras:2020:GTC

[47] Georges Gras. Genus theory and ε-conjectures on p-class groups. Journal of Number Theory, 207(??):423–459, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302616.

Matiyasevich:2020:PWC

[48] Yu. Matiyasevich. Plausible ways for calculating the Riemann zeta function via the Riemann–Siegel theta function. Journal of Number Theory, 207(??):460–471, February 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302744.

Anonymous:2020:PMa

[49] Anonymous. Pages 1–440 (March 2020). Journal of Number Theory, 208 (??):1–440, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBc

[50] Anonymous. Editorial Board. Journal of Number Theory, 208(??): ifc, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303610.

Kelmer:2020:EES

[51] Dubi Kelmer and Alex Kontorovich. Exponents for the equidistribution of shears and applications. Journal of Number Theory, 208(??): 1–46, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303208.

Steinerberger:2020:PPC

[52] Stefan Steinerberger. Poissonian pair correlation in higher dimensions. Journal of Number Theory, 208(??):47–58, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302641.

Rosen:2020:FAR

[53] Julian Rosen. A finite analogue of the ring of algebraic numbers. Journal of Number Theory, 208(??):59–71, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302707.

Aggarwal:2020:WBA

[54] Keshav Aggarwal. Weyl bound for GL(2) in t-aspect via a simple delta method. Journal of Number Theory, 208(??):72–100, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19302719.

Yu:2020:BAP

[55] Gang Yu. On a binary additive problem involving fractional powers. Journal of Number Theory, 208(??):101–119, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302720.

Hajli:2020:SZF

[56] Mounir Hajli. On the spectral zeta functions of the Laplacians on the projective complex spaces and on the n-spheres. Journal of Number Theory, 208(??):120–147, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302732.

Technau:2020:MHB

[57] Marc Technau. Modular hyperbolas and Beatty sequences. Journal of Number Theory, 208(??):148–167, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302756.

Krenn:2020:MHW

[58] Daniel Krenn, Vorapong Suppakitpaisarn, and Stephan Wagner. On the minimal Hamming weight of a multi-base representation. Journal of Number Theory, 208(??):168–179, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302768.

Kim:2020:STM

[59] Sungjin Kim. Some theorems on multiplicative orders modulo p on average. Journal of Number Theory, 208(??):180–207, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930277X.

Gao:2020:SSA

[60] Weidong Gao, Meiling Huang, Wanzhen Hui, Yuanlin Li, Chao Liu, and Jiangtao Peng. Sums of sets of abelian group elements. Journal of Number Theory, 208(??):208–229, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302793.

Wagh:2020:MSL

[61] Siddhesh Wagh. Maass space for lifts to GL(2) over a division quaternion algebra. Journal of Number Theory, 208(??):230–261, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930280X.

Hu:2020:LDS

[62] Sihuang Hu and Gabriele Nebe. Low dimensional strongly perfect lattices IV: the dual strongly perfect lattices of dimension 16. Journal of Number Theory, 208(??):262–294, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302811.

Monius:2020:ADS

[63] Katja Mönius. The algebraic degree of spectra of circulant graphs. Journal of Number Theory, 208(??):295–304, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302823.

Ota:2020:BHP

[64] Kazuto Ota. Big Heegner points and generalized Heegner cycles. Journal of Number Theory, 208(??):305–334, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19302859.

Kim:2020:RCA

[65] Dohyeong Kim. Ramification in the cohomology of algebraic surfaces arising from ordinary double point singularities. Journal of Number Theory, 208(??):335–345, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302860.

Martin:2020:DSC

[66] Yves Martin. On degree 2 Siegel cusp forms and its Fourier coefficients. Journal of Number Theory, 208(??):346–366, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302926.

Geisser:2020:WEV

[67] Thomas H. Geisser and Takashi Suzuki. A Weil-étale version of the Birch and Swinnerton-Dyer formula over function fields. Journal of Number Theory, 208(??):367–389, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302938.

Li:2020:NBH

[68] Weixiong Li. A note on the Bateman–Horn conjecture. Journal of Number Theory, 208(??):390–399, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302781.

Garcia:2020:PRB

[69] Stephan Ramon Garcia, Florian Luca, Kye Shi, and Gabe Udell. Primitive root bias for twin primes II: Schinzel-type theorems for totient quotients and the sum-of-divisors function. Journal of Number Theory, 208(??):400–417, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302835.

Gallegos-Ruiz:2020:DE

[70] H. R. Gallegos-Ruiz, N. Katsipis, Sz. Tengely, and M. Ulas. On the Diophantine equation \binom{n}{k} = \binom{m}{l} + d. Journal of Number Theory, 208(??):418–440, March 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302872.

Anonymous:2020:PAb

[71] Anonymous. Pages 1–530 (April 2020). Journal of Number Theory, 209 (??):1–530, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBd

[72] Anonymous. Editorial Board. Journal of Number Theory, 209(??): ifc, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304202.

Diamantis:2020:ATC

[73] Nikolaos Diamantis, Jeffrey Hoffstein, Eren Mehmet Kıral, and Min Lee. Additive twists and a conjecture by Mazur, Rubin and Stein. Journal of Number Theory, 209(??):1–36, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19304160.

Booker:2020:TVR

[74] Andrew R. Booker, M. Krishnamurthy, and Min Lee. Test vectors for Rankin–Selberg L-functions. Journal of Number Theory, 209(??):37– 48, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302902.

Jobrack:2020:NVD

[75] Matthew Jobrack. Non-vanishing of the derivative of L-functions at the central point. Journal of Number Theory, 209(??):49–82, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19302963.

Chen:2020:GFA

[76] Shih-Yu Chen, Yao Cheng, and Isao Ishikawa. Gamma factors for the Asai representation of GL_{2}. Journal of Number Theory, 209(??):83– 146, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302884.

Moree:2020:JTQ

[77] Pieter Moree, Sumaia Saad Eddin, Alisa Sedunova, and Yuta Suzuki. Jordan totient quotients. Journal of Number Theory, 209(??):147– 166, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930294X.

Murakami:2020:INM

[78] Yuya Murakami. Intersection numbers of modular correspondences for genus zero modular curves. Journal of Number Theory, 209(??):167– 194, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302951. See corrigendum [310].

Cardona:2020:TGC

[79] Gabriel Cardona and Joan-Carles Lario. Twists of the genus 2 curve Y^{2} = X^{6}+1. Journal of Number Theory, 209(??):195–211, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302975.

Mitra:2020:NDW

[80] Arnab Mitra. A note on degenerate Whittaker models for general linear groups. Journal of Number Theory, 209(??):212–224, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19302987.

Giard:2020:MMN

[81] Antoine Giard. Mahler measure of a non-tempered Weierstrass form. Journal of Number Theory, 209(??):225–245, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302999.

Yamada:2020:APN

[82] Tomohiro Yamada. An analog of perfect numbers involving the unitary totient function. Journal of Number Theory, 209(??):246–266, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303002.

Kishore:2020:IC

[83] Krishna Kishore. Inner cohomology of GL_{n} a. Journal of Number Theory, 209(??):267–288, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303014.

Yamamoto:2020:IEN

[84] Kota Yamamoto. On iterated extensions of number fields arising from quadratic polynomial maps. Journal of Number Theory, 209(??):289– 311, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303026.

Choi:2020:FCR

[85] Dohoon Choi and Suh Hyun Choi. Finiteness for crystalline representations of the absolute Galois group of a totally real field. Journal of Number Theory, 209(??):312–329, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19303166.

Xiong:2020:MPR

[86] Huan Xiong and Wenston J. T. Zang. Monotonicity properties for ranks of overpartitions. Journal of Number Theory, 209(??):330– 346, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930321X.

Saunders:2020:DEI

[87] J. C. Saunders. Diophantine equations involving the Euler totient function. Journal of Number Theory, 209(??):347–358, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19303336.

Gao:2020:MQH

[88] Peng Gao and Liangyi Zhao. Moments of quadratic Hecke L-functions of imaginary quadratic number fields. Journal of Number Theory, 209(??):359–377, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303348.

Verzobio:2020:PDS

[89] Matteo Verzobio. Primitive divisors of sequences associated to elliptic curves. Journal of Number Theory, 209(??):378–390, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930335X.

Perucca:2020:ARA

[90] Antonella Perucca, Pietro Sgobba, and Sebastiano Tronto. Addendum to: Reductions of algebraic integers [J. Number Theory 167 (2016) 259–283]. Journal of Number Theory, 209(??):391–395, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303361. See [1].

Eum:2020:CLC

[91] Ick Sun Eum, Ho Yun Jung, Ja Kyung Koo, and Dong Hwa Shin. Composition law and complex multiplication. Journal of Number Theory, 209(??):396–420, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303373.

Ito:2020:LTF

[92] Hiroshi Ito. Logarithms of theta functions on the upper half space. Journal of Number Theory, 209(??):421–450, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303385.

Filip:2020:RBA

[93] Ferdinánd Filip, Alexandr Jankov, and Jan Šustek. On relation between asymptotic and Abel densities. Journal of Number Theory, 209(??):451–466, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303397.

Kawamura:2020:HMM

[94] Satoshi Kawamura. On higher Mahler measures and zeta Mahler measures for one variable polynomials. Journal of Number Theory, 209(??): 467–482, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303403.

Williams:2020:GRP

[95] Brandon Williams. Graded rings of paramodular forms of levels 5 and 7. Journal of Number Theory, 209(??):483–515, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302896.

Turcas:2020:SMC

[96] George C. Ţurcaş. On Serre’s modularity conjecture and Fermat’s equation over quadratic imaginary fields of class number one. Journal of Number Theory, 209(??):516–530, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302914.

Anonymous:2020:PMb

[97] Anonymous. Pages 1–488 (May 2020). Journal of Number Theory, 210 (??):1–488, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBe

[98] Anonymous. Editorial Board. Journal of Number Theory, 210(??): ifc, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300020.

Hough:2020:LZF

[99] Robert D. Hough. The local zeta function in enumerating quartic fields. Journal of Number Theory, 210(??):1–131, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300251.

Duke:2020:RMC

[100] William Duke and Ha Nam Nguyen. Riesz means of certain arithmetic functions. Journal of Number Theory, 210(??):132–141, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303798.

Chung:2020:SFW

[101] Fan Chung, Ron Graham, and Sam Spiro. Slow Fibonacci walks. Journal of Number Theory, 210(??):142–170, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303488.

Pollack:2020:GHR

[102] Paul Pollack. A generalization of the Hardy–Ramanujan inequality and applications. Journal of Number Theory, 210(??):171–182, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303762.

Bonanno:2020:SEM

[103] Claudio Bonanno and Stefano Isola. Series expansions for Maass forms on the full modular group from the Farey transfer operators. Journal of Number Theory, 210(??):183–230, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303415.

Tanuma:2020:AIV

[104] Yusuke Tanuma. Algebraic independence of the values of certain series and their derivatives involving the Hecke–Mahler series. Journal of Number Theory, 210(??):231–248, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19303427.

Kim:2020:RKE

[105] Chan-Ho Kim and Kentaro Nakamura. Remarks on Kato’s Euler systems for elliptic curves with additive reduction. Journal of Number Theory, 210(??):249–279, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303439.

Goudout:2020:TEK

[106] Élie Goudout. Théorème d’Erdős–Kac pour les translatés d’entiers ayant k facteurs premiers. (French) [Erdős–Kac theorem for translates of integers having k prime factors]. Journal of Number Theory, 210(??):280–291, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303440.

Franze:2020:APV

[107] C. S. Franze and P. H. Kao. Almost-prime values of reducible polynomials at prime arguments. Journal of Number Theory, 210(??):292– 312, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303452.

Cynk:2020:HMS

[108] Sławomir Cynk, Matthias Schütt, and Duco van Straten. Hilbert modularity of some double octic Calabi–Yau threefolds. Journal of Number Theory, 210(??):313–332, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303476.

Koleda:2020:DRA

[109] Denis V. Koleda. Distribution of real algebraic integers. Journal of Number Theory, 210(??):333–372, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1930349X.

Kravitz:2020:SCB

[110] Noah Kravitz and Ashwin Sah. A stronger connection between the Erdős–Burgess and Davenport constants. Journal of Number Theory, 210(??):373–388, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303506.

Duan:2020:GAN

[111] Lian Duan. Galois action on the Néron–Severi group of Dwork surfaces. Journal of Number Theory, 210(??):389–415, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930352X.

Kadets:2020:LAG

[112] Borys Kadets. Large arboreal Galois representations. Journal of Number Theory, 210(??):416–430, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303531.

Baier:2020:ELS

[113] Stephan Baier and Rajneesh Kumar Singh. Erratum to “Large sieve inequality with power moduli for function fields” [J. Number Theory 196 (2019) 1–13]. Journal of Number Theory, 210(??):431– 432, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303774. See [5].

Cipu:2020:DPI

[114] Mihai Cipu, Alan Filipin, and Yasutsugu Fujita. Diophantine pairs that induce certain Diophantine triples. Journal of Number Theory, 210(??):433–475, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303518.

Grantham:2020:UIR

[115] Jon Grantham. An unconditional improvement to the running time of the quadratic Frobenius test. Journal of Number Theory, 210(??): 476–480, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303786.

Bordignon:2020:EBE

[116] Matteo Bordignon. Explicit bounds on exceptional zeroes of Dirichlet L-functions II. Journal of Number Theory, 210(??):481–487, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1930383X.

Anonymous:2020:PJa

[117] Anonymous. Pages 1–544 (June 2020). Journal of Number Theory, 211 (??):1–544, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBf

[118] Anonymous. Editorial Board. Journal of Number Theory, 211(??): ifc, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300482.

Cerbu:2020:CSM

[119] Alois Cerbu, Elijah Gunther, Michael Magee, and Luke Peilen. The cycle structure of a Markoff automorphism over finite fields. Journal of Number Theory, 211(??):1–27, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303543.

Khuri-Makdisi:2020:VSH

[120] Kamal Khuri-Makdisi, Winfried Kohnen, and Wissam Raji. Values of L-series of Hecke eigenforms. Journal of Number Theory, 211(??): 28–42, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303865.

Sachdeva:2020:CVF

[121] Gunja Sachdeva. Critical values of L-functions for GL_{3} × GL_{1} and symmetric square L-functions for GL_{2} over a CM field. Journal of Number Theory, 211(??):43–74, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303853.

Kohen:2020:GKZ

[122] Daniel Kohen and Nicolás Sirolli. A Gross–Kohnen–Zagier theorem for non-split Cartan curves. Journal of Number Theory, 211(??):75– 94, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303804.

Ohshita:2020:ALB

[123] Tatsuya Ohshita. Asymptotic lower bound of class numbers along a Galois representation. Journal of Number Theory, 211(??):95– 112, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303567.

Elaissaoui:2020:FER

[124] Lahoucine Elaissaoui and Zine El Abidine Guennoun. Fourier expansion of the Riemann zeta function and applications. Journal of Number Theory, 211(??):113–138, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303580.

Feng:2020:ESB

[125] Zhenzhen Feng, Zhixin Liu, and Jing Ma. On exceptional sets of biquadratic Waring–Goldbach problem. Journal of Number Theory, 211(??):139–154, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303592.

Jung:2020:DSA

[126] Myunghyun Jung and Donggyun Kim. Dynamic structures of p-adic monomials for small primes p. Journal of Number Theory, 211(??):155– 183, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303737.

Li:2020:DAP

[127] Taiyu Li and Huafeng Liu. Diophantine approximation over Piatetski– Shapiro primes. Journal of Number Theory, 211(??):184–198, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303749.

Zhai:2020:SIE

[128] Wenguang Zhai. On a sum involving the Euler function. Journal of Number Theory, 211(??):199–219, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19303750.

Wang:2020:TSD

[129] Ruoyi Wang. On a theorem of Sárközy for difference sets and shifted primes. Journal of Number Theory, 211(??):220–234, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19303816.

Czerniawska:2020:AGA

[130] Weronika Czerniawska and Paolo Dolce. Adelic geometry on arithmetic surfaces II: Completed adeles and idelic Arakelov intersection theory. Journal of Number Theory, 211(??):235–296, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19303828.

Moreschi:2020:PVU

[131] Maurizio Moreschi. S-parts of values of univariate polynomials. Journal of Number Theory, 211(??):297–321, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303841.

Zmija:2020:RSC

[132] Błażej Żmija. Recurrence sequences connected with the m-ary partition function and their divisibility properties. Journal of Number Theory, 211(??):322–370, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303555.

Filipow:2020:DSN

[133] Rafał Filipów and Jacek Tryba. Densities for sets of natural numbers vanishing on a given family. Journal of Number Theory, 211(??):371– 382, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303877.

Edwards:2020:RIP

[134] Samuel C. Edwards. Renormalization of integrals of products of Eisenstein series and analytic continuation of representations. Journal of Number Theory, 211(??):383–454, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19303889.

Dajani:2020:IMN

[135] Karma Dajani, Mathijs R. de Lepper, and E. Arthur Robinson. Introducing Minkowski normality. Journal of Number Theory, 211(??): 455–476, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303890.

Jung:2020:NMV

[136] Hwanyup Jung. A note on the mean value of L(\frac{1}{2}, χ) in the real hyperelliptic ensemble. Journal of Number Theory, 211(??):477–499, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303907.

Macedo:2020:HNP

[137] André Macedo. The Hasse norm principle for A_{n}-extensions. Journal of Number Theory, 211(??):500–512, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303920.

Lichtman:2020:APB

[138] Jared Duker Lichtman. Almost primes and the Banks–Martin conjecture. Journal of Number Theory, 211(??):513–529, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19304068.

Cowan:2020:DMR

[139] Alex Cowan. The distribution of multiples of real points on an elliptic curve. Journal of Number Theory, 211(??):530–544, June 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20300184.

Anonymous:2020:PJb

[140] Anonymous. Pages 1–482 (July 2020). Journal of Number Theory, 212 (??):1–482, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBg

[141] Anonymous. Editorial Board. Journal of Number Theory, 212(??): ifc, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300767.

Ellers:2020:EBT

[142] Havi Ellers, Meagan Kenney, Riad Masri, and Wei-Lun Tsai. Effective bounds for traces of singular moduli. Journal of Number Theory, 212(??):1–17, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300196.

Garai:2020:ERR

[143] Sumita Garai and Mihran Papikian. Endomorphism rings of reductions of Drinfeld modules. Journal of Number Theory, 212(??):18– 39, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300848.

Lee:2020:HTD

[144] Junguk Lee. Hyperfields, truncated DVRs, and valued fields. Journal of Number Theory, 212(??):40–71, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19303919.

Konig:2020:UEL

[145] Joachim König, Danny Neftin, and Jack Sonn. Unramified extensions over low degree number fields. Journal of Number Theory, 212(??): 72–87, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303932.

Haynes:2020:HMD

[146] Alan Haynes and Kayleigh Measures. Higher moments of distances between consecutive Ford spheres. Journal of Number Theory, 212(??): 88–104, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303944.

Ray:2020:CCS

[147] Anwesh Ray. Constructing certain special analytic Galois extensions. Journal of Number Theory, 212(??):105–112, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19303956.

Aymone:2020:NMF

[148] Marco Aymone. A note on multiplicative functions resembling the Möbius function. Journal of Number Theory, 212(??):113–121, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1930397X.

Rada:2020:BPC

[149] Hanka Řada and Štěpán Starosta. Bounds on the period of the continued fraction after a Möbius transformation. Journal of Number Theory, 212(??):122–172, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19303993.

Sagnier:2020:ASC

[150] Aurélien Sagnier. An arithmetic site at the complex place. Journal of Number Theory, 212(??):173–202, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19304007.

Cho:2020:DCL

[151] Peter J. Cho and Jeongho Park. Dirichlet characters and lowlying zeros of L-functions. Journal of Number Theory, 212(??):203– 232, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304019.

Ching:2020:LEA

[152] Tak Wing Ching. Lagrange’s equation with almost-prime variables. Journal of Number Theory, 212(??):233–264, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19304020.

Kerr:2020:CPV

[153] Bryce Kerr. On the constant in the Pólya–Vinogradov inequality. Journal of Number Theory, 212(??):265–284, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19304032.

Bienvenu:2020:DMS

[154] Pierre-Yves Bienvenu and François Hennecart. On the density or measure of sets and their sumsets in the integers or the circle. Journal of Number Theory, 212(??):285–310, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19304044.

Kluners:2020:CDL

[155] Jürgen Klüners and Raphael Müller. The conductor density of local function fields with abelian Galois group. Journal of Number Theory, 212(??):311–322, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930407X.

Dillon:2020:AGD

[156] Travis Dillon and Stephanie Gaston. An average of generalized Dedekind sums. Journal of Number Theory, 212(??):323–338, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19304081.

Alsina:2020:CPL

[157] Montserrat Alsina and Dimitrios Chatzakos. CM-points and lattice counting on arithmetic compact Riemann surfaces. Journal of Number Theory, 212(??):339–353, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19304093.

Peng:2020:AIT

[158] Wayne Peng. ABC implies there are infinitely many non–Fibonacci– Wieferich primes. Journal of Number Theory, 212(??):354–375, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1930410X.

Giri:2020:SAD

[159] Sumit Giri. Short average distribution of a prime counting function over families of elliptic curves. Journal of Number Theory, 212(??): 376–408, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304111.

Das:2020:DSK

[160] Mithun Kumar Das and Sudhir Pujahari. Distribution of signs of Karatsuba’s and generalized Davenport–Heilbronn Z-functions. Journal of Number Theory, 212(??):409–447, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19304123.

Morrill:2020:EBS

[161] Thomas Morrill and Tim Trudgian. An elementary bound on Siegel zeroes. Journal of Number Theory, 212(??):448–457, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19303968.

Tinkova:2020:IIR

[162] Magdaléna Tinková and Paul Voutier. Indecomposable integers in real quadratic fields. Journal of Number Theory, 212(??):458–482, July 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304056.

Anonymous:2020:PAc

[163] Anonymous. Pages 1–498 (August 2020). Journal of Number Theory, 213 (??):1–498, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBh

[164] Anonymous. Editorial Board. Journal of Number Theory, 213(??): ifc, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300962.

Sawin:2020:DML

[165] Will Sawin. Dynamical models for Liouville and obstructions to further progress on sign patterns. Journal of Number Theory, 213(??):1– 15, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300469.

Karpenkov:2020:GMN

[166] Oleg Karpenkov and Matty van Son. Generalised Markov numbers. Journal of Number Theory, 213(??):16–66, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300433.

Hussain:2020:MTS

[167] Mumtaz Hussain, Simon Kristensen, and David Simmons. Metrical theorems on systems of affine forms. Journal of Number Theory, 213(??):67–100, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19304147.

Biswas:2020:ACI

[168] Arindam Biswas and Jyoti Prakash Saha. Asymptotic complements in the integers. Journal of Number Theory, 213(??):101–115, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304159.

Hedayatzadeh:2020:TCD

[169] S. Mohammad Hadi Hedayatzadeh. Tensor constructions on z-divisible local Anderson modules. Journal of Number Theory, 213(??):116– 151, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19304172.

Naskrecki:2020:PDE

[170] Bartosz Naskrȩcki and Marco Streng. Primitive divisors of elliptic divisibility sequences over function fields with constant j-invariant. Journal of Number Theory, 213(??):152–186, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2030010X.

Francirek:2020:AIC

[171] Pavel Francírek. Annihilators of the ideal class group of an imaginary cyclic field. Journal of Number Theory, 213(??):187–220, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300111.

Sanna:2020:CMR

[172] Carlo Sanna. On the l.c.m. of random terms of binary recurrence sequences. Journal of Number Theory, 213(??):221–231, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300123.

Halbeisen:2020:HTT

[173] Lorenz Halbeisen and Norbert Hungerbühler. Heron triangles and their elliptic curves. Journal of Number Theory, 213(??):232–253, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300135.

Kim:2020:QFS

[174] Kyoungmin Kim and Byeong-Kweon Oh. Quadratic forms with a strong regularity property on the representations of squares. Journal of Number Theory, 213(??):254–270, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300159.

Grenie:2020:CUB

[175] Loïc Grenié and Giuseppe Molteni. Conditional upper bound for the k-th prime ideal with given Artin symbol. Journal of Number Theory, 213(??):271–284, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300160.

Eisenstein:2020:EAC

[176] Eugene Eisenstein, Lalit K. Jain, and Wentang Kuo. On extending Artin’s conjecture to composite moduli in function fields. Journal of Number Theory, 213(??):285–318, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300172.

Beshaj:2020:WGC

[177] L. Beshaj, J. Gutierrez, and T. Shaska. Weighted greatest common divisors and weighted heights. Journal of Number Theory, 213(??):319– 346, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300202.

Kwon:2020:NVS

[178] Jaesung Kwon and Hae-Sang Sun. Non-vanishing of special L-values of cusp forms on GL(2) with totally split prime power twists. Journal of Number Theory, 213(??):347–369, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300214.

Horiba:2020:GTS

[179] Hiroto Horiba, Masanari Kida, and Genki Koda. Galois theoretic study on simultaneous representation of primes by binary quadratic forms. Journal of Number Theory, 213(??):370–387, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20300226.

Li:2020:CMA

[180] Shuo Li. On completely multiplicative automatic sequences. Journal of Number Theory, 213(??):388–399, August 2020. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300238.

Florez:2020:ERL

[181] Jorge Flórez. Explicit reciprocity laws for higher local fields. Journal of Number Theory, 213(??):400–444, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2030024X.

Badr:2020:QPS

[182] Eslam Badr and Mohammad Sadek. On quadratic progression sequences on smooth plane curves. Journal of Number Theory, 213(??):445– 452, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300263.

Paran:2020:TPG

[183] Elad Paran. Twin-prime and Goldbach theorems for Z[[x]]. Journal of Number Theory, 213(??):453–461, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300275.

Dixit:2020:CUP

[184] Anup B. Dixit. Corrigendum to “A uniqueness property of general Dirichlet series” [J. Number Theory 206 (2020) 123–137]. Journal of Number Theory, 213(??):462–463, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300615. See [16].

Martindale:2020:HMP

[185] Chloe Martindale. Hilbert modular polynomials. Journal of Number Theory, 213(??):464–498, August 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300743.

Anonymous:2020:PS

[186] Anonymous. Pages 1–452 (September 2020). Journal of Number Theory, 214(??):1–452, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBi

[187] Anonymous. Editorial Board. Journal of Number Theory, 214(??):ifc, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301463.

Griffin:2020:ECL

[188] Michael Griffin and Ken Ono. Elliptic curves and lower bounds for class numbers. Journal of Number Theory, 214(??):1–12, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301311.

Jenkins:2020:CCM

[189] Paul Jenkins and Ryan Keck. Congruences for coefficients of modular functions in levels 3, 5, and 7 with poles at 0. Journal of Number Theory, 214(??):13–26, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301268.

Chen:2020:HDL

[190] Changhao Chen and Igor E. Shparlinski. Hausdorff dimension of the large values of Weyl sums. Journal of Number Theory, 214(??):27–37, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301438.

Li:2020:DGM

[191] Jiangtao Li. Depth-graded Motivic Lie algebra. Journal of Number Theory, 214(??):38–55, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20301384.

Mozzochi:2020:PSG

[192] C. J. Mozzochi. A proof of Sarnak’s golden mean conjecture. Journal of Number Theory, 214(??):56–62, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300640.

Stier:2020:OTG

[193] Zachary Stier. Optimal topological generators of U(1). Journal of Number Theory, 214(??):63–78, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300627.

Hajdu:2020:PVF

[194] Lajos Hajdu and Nóra Varga. Polynomial values of figurate numbers. Journal of Number Theory, 214(??):79–99, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20301359.

Eddin:2020:DPT

[195] Sumaia Saad Eddin and Yuta Suzuki. On the distribution of products of two primes. Journal of Number Theory, 214(??):100–136, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301347.

Kim:2020:RTT

[196] Mingyu Kim and Byeong-Kweon Oh. Regular ternary triangular forms. Journal of Number Theory, 214(??):137–169, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20301360.

Beyerl:2020:RCB

[197] Jeffrey Beyerl. Rankin–Cohen brackets of eigenforms and modular forms. Journal of Number Theory, 214(??):170–176, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20301414.

Xu:2020:SRM

[198] Ce Xu. Some results on multiple polylogarithm functions and alternating multiple zeta values. Journal of Number Theory, 214(??):177–201, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301372.

Huicochea:2020:NPD

[199] Mario Huicochea. On the number of popular differences in Z/pZ. Journal of Number Theory, 214(??):202–239, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301293.

Booher:2020:RAS

[200] Jeremy Booher and Rachel Pries. Realizing Artin–Schreier covers of curves with minimal Newton polygons in positive characteristic. Journal of Number Theory, 214(??):240–250, September 2020. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301207.

Peltomaki:2020:APF

[201] Jarkko Peltomäki. Abelian periods of factors of Sturmian words. Journal of Number Theory, 214(??):251–285, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2030127X.

Tsang:2020:HGS

[202] Cindy (Sin Yi) Tsang. Hopf–Galois structures on finite extensions with almost simple Galois group. Journal of Number Theory, 214 (??):286–311, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301244.

Hu:2020:CRP

[203] L. Q. Hu. Counting rational points on biprojective hypersurfaces of bidegree (1, 2). Journal of Number Theory, 214(??):312–325, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301256.

Ren:2020:IPF

[204] Rufei Ren. Iteration of polynomials AX^{d}+C over finite fields. Journal of Number Theory, 214(??):326–347, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300937.

Riman:2020:VBG

[205] Manar Riman. Vanishing of the Brauer group of a del Pezzo surface of degree 4. Journal of Number Theory, 214(??):348–359, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20300913.

Grynkiewicz:2020:SSI

[206] David J. Grynkiewicz. A single set improvement to the 3k − 4 theorem. Journal of Number Theory, 214(??):360–381, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030086X.

Gritsenko:2020:GRI

[207] Valery Gritsenko and Haowu Wang. Graded rings of integral Jacobi forms. Journal of Number Theory, 214(??):382–398, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20300925.

Toma:2020:ERN

[208] Radu Toma. Estimates for representation numbers of binary quadratic forms and Apollonian circle packings. Journal of Number Theory, 214 (??):399–413, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300883.

Tsushima:2020:GRA

[209] Takahiro Tsushima. Good reduction of affinoids in the Lubin–Tate curve in even equal characteristic, I. Journal of Number Theory, 214 (??):414–439, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300949.

Argaez-Garcia:2020:PPS

[210] Alejandro Argáez-García and Vandita Patel. On perfect powers that are sums of cubes of a seven term arithmetic progression. Journal of Number Theory, 214(??):440–451, September 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301323.

Anonymous:2020:PO

[211] Anonymous. Pages 1–362 (October 2020). Journal of Number Theory, 215(??):1–362, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBj

[212] Anonymous. Editorial Board. Journal of Number Theory, 215(??): ifc, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301797.

Baker:2020:ERS

[213] Simon Baker. Equidistribution results for sequences of polynomials. Journal of Number Theory, 215(??):1–19, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20300366.

Kerr:2020:LPR

[214] Bryce Kerr, Kevin J. McGown, and Tim Trudgian. The least primitive root modulo p ^{2}. Journal of Number Theory, 215(??):20–27, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300354.

Ji:2020:STM

[215] Qingchun Ji, Qiming Yan, and Guangsheng Yu. Subspace theorem for moving hypersurfaces and semi-decomposable form inequalities. Journal of Number Theory, 215(??):28–51, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300342.

Xiong:2020:SWF

[216] Wei Xiong. On the Siegel–Weil formula for classical groups over function fields. Journal of Number Theory, 215(??):52–97, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20300408.

Paul:2020:ETZ

[217] Biplab Paul and Ayyadurai Sankaranarayanan. On the error term and zeros of the Dedekind zeta function. Journal of Number Theory, 215(??):98–119, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300652.

Dubickas:2020:RPN

[218] Artūras Dubickas and Aivaras Novikas. Recurrence with prescribed number of residues. Journal of Number Theory, 215(??):120–137, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300378.

Kobayashi:2020:IW

[219] Mits Kobayashi and Tim Trudgian. On integers n for which σ(2n + 1) ≥ σ(2n). Journal of Number Theory, 215(??):138–148, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20300664.

Dabrowski:2020:CLL

[220] Andrzej Da̧browski, Nursena Günhan, and Gökhan Soydan. On a class of Lebesgue–Ljunggren–Nagell type equations. Journal of Number Theory, 215(??):149–159, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300330.

Kaziulyte:2020:VKT

[221] Laima Kaziulytė. Variants of Khintchine’s theorem in metric Diophantine approximation. Journal of Number Theory, 215(??):160–170, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030038X.

Ayadi:2020:QAH

[222] Khalil Ayadi and Tomohiro Ooto. On quadratic approximation for hyperquadratic continued fractions. Journal of Number Theory, 215 (??):171–185, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300676.

Cai:2020:CGD

[223] Yuanqing Cai. On the crystal graph description of the stable Weyl group multiple Dirichlet series. Journal of Number Theory, 215(??):186– 215, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300391.

Karasiewicz:2020:FCM

[224] Edmund Karasiewicz. The Fourier coefficients of a metaplectic Eisenstein distribution on the double cover of SL(3) over Q. Journal of Number Theory, 215(??):216–260, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300445.

DiBenedetto:2020:NEE

[225] Daniel Di Benedetto, Moubariz Z. Garaev, Victor C. Garcia, Diego Gonzalez-Sanchez, Igor E. Shparlinski, and Carlos A. Trujillo. New estimates for exponential sums over multiplicative subgroups and intervals in prime fields. Journal of Number Theory, 215(??):261–274, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300639.

Wakhare:2020:RCJ

[226] Tanay Wakhare. Romik’s conjecture for the Jacobi theta function. Journal of Number Theory, 215(??):275–296, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20300421.

Sang:2020:PCR

[227] Doris D. M. Sang, Diane Y. H. Shi, and Ae Ja Yee. Parity considerations in Rogers–Ramanujan–Gordon type overpartitions. Journal of Number Theory, 215(??):297–320, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2030069X.

Ray:2020:AIP

[228] Chiranjit Ray and Rupam Barman. On Andrews’ integer partitions with even parts below odd parts. Journal of Number Theory, 215 (??):321–338, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300457.

Hammonds:2020:RBF

[229] Trajan Hammonds, Seoyoung Kim, Benjamin Logsdon, Álvaro Lozano-Robledo, and Steven J. Miller. Rank and bias in families of hyperelliptic curves via Nagao’s conjecture. Journal of Number Theory, 215(??):339–361, October 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301335.

Anonymous:2020:PN

[230] Anonymous. Pages 1–460 (November 2020). Journal of Number Theory, 216(??):1–460, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBk

[231] Anonymous. Editorial Board. Journal of Number Theory, 216(??):ifc, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302055.

Bleher:2020:GSH

[232] Frauke M. Bleher, Ted Chinburg, and Aristides Kontogeorgis. Galois structure of the holomorphic differentials of curves. Journal of Number Theory, 216(??):1–68, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20301426.

Schoof:2020:ICG

[233] René Schoof. On the ideal class group of the normal closure of Q( ^{p} \sqrt{n}). Journal of Number Theory, 216(??):69–82, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20301281.

Morimoto:2020:FR

[234] Kazuki Morimoto and David Soudry. On L-functions for U_{2n+1} × Res_{E/F} GL_{m}(m > n). Journal of Number Theory, 216(??):83–141, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030130X.

Coccia:2020:VST

[235] Simone Coccia and Dragos Ghioca. A variant of Siegel’s theorem for Drinfeld modules. Journal of Number Theory, 216(??):142–156, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301190.

Dai:2020:OTA

[236] Lixia Dai and Hao Pan. An oscillation theorem on the additive representative function over N. Journal of Number Theory, 216(??):157–173, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300688.

The:2020:EPH

[237] Nguyen Van The and Le Anh Vinh. Expanding phenomena over higher dimensional matrix rings. Journal of Number Theory, 216(??):174–191, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300706.

Gotze:2020:DSN

[238] Friedrich Götze and Anna Gusakova. On the distribution of Salem numbers. Journal of Number Theory, 216(??):192–215, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20300718.

Banerjee:2020:GCG

[239] Pradipto Banerjee and Ranjan Bera. On a generalization of a conjecture of Grosswald. Journal of Number Theory, 216(??):216–241, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030072X. See corrigendum [373].

Villanueva-Gutierrez:2020:MLI

[240] José-Ibrahim Villanueva-Gutiérrez. On the mu and lambda invariants of the logarithmic class group. Journal of Number Theory, 216 (??):242–263, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300731.

Sui:2020:ETC

[241] Yankun Sui and Dan Liu. On the error term concerning the number of subgroups of the groups Z_{m} × Z_{n} with mn ≤ x. Journal of Number Theory, 216(??):264–279, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300895.

Wakhare:2020:TCJ

[242] Tanay Wakhare and Christophe Vignat. Taylor coefficients of the Jacobi θ_{3}(q) function. Journal of Number Theory, 216(??):280–306, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300858.

Ji:2020:PBC

[243] Caleb Ji, Joshua Kazdan, and Vaughan McDonald. Primes with Beatty and Chebotarev conditions. Journal of Number Theory, 216(??):307– 334, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20300901.

OSullivan:2020:RCE

[244] Cormac O’Sullivan. Rademacher’s conjecture and expansions at roots of unity of products generating restricted partitions. Journal of Number Theory, 216(??):335–379, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20301232.

Bars:2020:BQM

[245] Francesc Bars, Josep González, and Mohamed Kamel. Bielliptic quotient modular curves with N square-free. Journal of Number Theory, 216(??):380–402, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20300871.

Milio:2020:MPH

[246] Enea Milio and Damien Robert. Modular polynomials on Hilbert surfaces. Journal of Number Theory, 216(??):403–459, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301402.

Anonymous:2020:PD

[247] Anonymous. Pages 1–494 (December 2020). Journal of Number Theory, 217(??):1–494, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2020:EBl

[248] Anonymous. Editorial Board. Journal of Number Theory, 217(??):ifc, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302420.

Murty:2020:CVL

[249] M. Ram Murty and Siddhi Pathak. Convolution of values of the Lerch zeta-function. Journal of Number Theory, 217(??):1–22, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030041X.

Lowry-Duda:2020:NRP

[250] David Lowry-Duda. Non-real poles and irregularity of distribution. Journal of Number Theory, 217(??):23–35, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030158X.

Elma:2020:PRD

[251] Ertan Elma. On a problem related to discrete mean values of Dirichlet L-functions. Journal of Number Theory, 217(??):36–43, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301712.

Sardari:2020:RGE

[252] Naser T. Sardari and Masoud Zargar. Ramanujan graphs and exponential sums over function fields. Journal of Number Theory, 217(??):44–77, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301645.

Hidalgo:2020:HDG

[253] Rubén A. Hidalgo. Holomorphic differentials of generalized Fermat curves. Journal of Number Theory, 217(??):78–101, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301694.

Essouabri:2020:DMC

[254] Driss Essouabri and Oswaldo Velásquez Castañón. Domain of meromorphy of a class of Bateman-like totient zeta functions. Journal of Number Theory, 217(??):102–141, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19303981.

Bedert:2020:TM

[255] Benjamin Bedert, George Cooper, Thomas Oliver, and Pengcheng Zhang. Twisting moduli for GL(2). Journal of Number Theory, 217(??):142–162, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301219.

Jarossay:2020:DRA

[256] David Jarossay. Depth reductions for associators. Journal of Number Theory, 217(??):163–192, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20301396.

Peng:2020:SSZ

[257] Jiangtao Peng, Yuanlin Li, Chao Liu, and Meiling Huang. On subsequence sums of a zero-sum free sequence over finite abelian groups. Journal of Number Theory, 217(??):193–217, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030144X.

Mahanta:2020:SPZ

[258] Pankaj Jyoti Mahanta, Manjil P. Saikia, and Daniel Yaqubi. Some properties of Zumkeller numbers and k-layered numbers. Journal of Number Theory, 217(??):218–236, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20301591.

Ambi:2020:GCC

[259] Chaitanya Ambi. On the growth of cuspidal cohomology of GL(2) and GL(3). Journal of Number Theory, 217(??):237–255, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301633.

Altassan:2020:E

[260] Alaa Altassan and Florian Luca. On the equation ∑_{j=1}^{k} jF_{j}^{p} = F_{n}^{q}. Journal of Number Theory, 217(??):256–277, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20301578.

Lev:2020:SDP

[261] Vsevolod F. Lev and Ilya D. Shkredov. Small doubling in prime-order groups: From 2.4 to 2.6. Journal of Number Theory, 217(??):278–291, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301621.

Hegyvari:2020:HCM

[262] Norbert Hegyvári and Péter Pál Pach. Hilbert cubes meet arithmetic sets. Journal of Number Theory, 217(??):292–300, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301736.

Heier:2020:DIP

[263] Gordon Heier and Aaron Levin. On the degeneracy of integral points and entire curves in the complement of nef effective divisors. Journal of Number Theory, 217(??):301–319, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301700.

Li:2020:PAS

[264] Xiumei Li and Min Sha. Polynomial analogue of the Smarandache function. Journal of Number Theory, 217(??):320–339, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301724.

Jordan:2020:IPV

[265] Bruce W. Jordan and Yevgeny Zaytman. Integral points on varieties defined by matrix factorization into elementary matrices. Journal of Number Theory, 217(??):340–352, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301748.

Choi:2020:CHE

[266] SoYoung Choi, Chang Heon Kim, and Kyung Seung Lee. Congruences for Hecke eigenvalues in minus spaces. Journal of Number Theory, 217(??):353–375, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301670.

Chiu:2020:SSC

[267] Ching-Heng Chiu. Strong Selmer companion elliptic curves. Journal of Number Theory, 217(??):376–421, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302006.

Prado:2020:LOH

[268] Bernardo Bianco Prado and Kim Klinger-Logan. Linear operators, the Hurwitz zeta function and Dirichlet L-functions. Journal of Number Theory, 217(??):422–442, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301657. See

corrigendum [602].

Chang:2020:CAN

[269] Sungkon Chang. Corrigendum to “Average number of Zeckendorf integers” [J. Number Theory 186 (2018) 452–472]. Journal of Number Theory, 217(??):443–444, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2030216X. See [2].

Tengely:2020:DEE

[270] Szabolcs Tengely, Maciej Ulas, and Jakub Zygadło. On a Diophantine equation of Erdős and Graham. Journal of Number Theory, 217 (??):445–459, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301608.

Martin:2020:ZQM

[271] Kimball Martin and Jordan Wiebe. Zeroes of quaternionic modular forms and central L-values. Journal of Number Theory, 217(??):460– 494, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030161X.

Anonymous:2021:PJa

[272] Anonymous. Pages 1–402 (January 2021). Journal of Number Theory, 218(??):1–402, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBa

[273] Anonymous. Editorial Board. Journal of Number Theory, 218(??): ifc, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302730.

Watkins:2021:CDZ

[274] Mark Watkins. Comments on Deuring’s zero-spacing phenomenon. Journal of Number Theory, 218(??):1–43, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302249.

Babei:2021:RHP

[275] Angelica Babei, Larry Rolen, and Ian Wagner. The Riemann hypothesis for period polynomials of Hilbert modular forms. Journal of Number Theory, 218(??):44–61, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20302195.

Hartl:2021:PFP

[276] Urs Hartl and Rajneesh Kumar Singh. Product formulas for periods of CM abelian varieties and the function field analog. Journal of Number Theory, 218(??):62–151, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20302547.

Sun:2021:PNM

[277] Cui-Fang Sun. On a problem of Nathanson on minimal asymptotic bases. Journal of Number Theory, 218(??):152–160, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302225.

Asensouyis:2021:GFP

[278] Hassan Asensouyis, Jilali Assim, and Youness Mazigh. A genus formula for the positive étale wild kernel. Journal of Number Theory, 218(??):161–179, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302262.

Ernvall-Hytonen:2021:FHM

[279] Anne-Maria Ernvall-Hytönen and Esa V. Vesalainen. On fourth and higher moments of short exponential sums related to cusp forms. Journal of Number Theory, 218(??):180–208, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302286.

Wang:2021:SSM

[280] Zhiqiang Wang, Kan Jiang, Wenxia Li, and Bing Zhao. On the sum of squares of the middle-third Cantor set. Journal of Number Theory, 218(??):209–222, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302237.

Anh:2021:MER

[281] Dao Nguyen Van Anh, Le Quang Ham, Doowon Koh, Mozhgan Mirzaei, Hossein Mojarrad, and Thang Pham. Moderate expanders over rings. Journal of Number Theory, 218(??):223–233, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302250.

Morton:2021:HITa

[282] Patrick Morton. On the Hasse invariants of the Tate normal forms E_{5} and E_{7}. Journal of Number Theory, 218(??):234–271, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302304.

Marques:2021:EST

[283] Diego Marques and Carlos Gustavo Moreira. On the exceptional set of transcendental functions with integer coefficients in a prescribed set: the Problems A and C of Mahler. Journal of Number Theory, 218(??):272–287, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302274.

Jiang:2021:VRF

[284] Xing-Wang Jiang, Csaba Sándor, and Quan-Hui Yang. On the values of representation functions II. Journal of Number Theory, 218 (??):288–301, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302298.

Suzuki:2021:CSM

[285] Miyu Suzuki. Classification of standard modules with linear periods. Journal of Number Theory, 218(??):302–310, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302183.

Bandi:2021:GEO

[286] Prasuna Bandi, Anish Ghosh, and Jiyoung Han. A generic effective Oppenheim theorem for systems of forms. Journal of Number Theory, 218(??):311–333, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302201.

Djankovic:2021:IFM

[287] Goran Djanković, Dragan – Dokić, Nikola Lelas, and Ilija Vrećica. The integrated fourth moment of Dirichlet L-functions over rational function fields. Journal of Number Theory, 218(??):334–369, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302213. See corrigendum [991].

Spiro:2021:SR

[288] Sam Spiro. Slow recurrences. Journal of Number Theory, 218(??):370– 401, January 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301669.

Anonymous:2021:PF

[289] Anonymous. Pages 1–466 (February 2021). Journal of Number Theory, 219(??):1–466, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBb

[290] Anonymous. Editorial Board. Journal of Number Theory, 219(??):ifc, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303024.

Haberli:2021:SCA

[291] Simon Häberli. Satake compactification of analytic Drinfeld modular varieties. Journal of Number Theory, 219(??):1–92, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302985.

Nathanson:2021:CS

[292] Melvyn B. Nathanson. Chromatic sumsets. Journal of Number Theory, 219(??):93–108, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302869.

Mercuri:2021:AMF

[293] Salvatore Mercuri. Algebraicity of metaplectic L-functions. Journal of Number Theory, 219(??):109–161, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302924.

Kirila:2021:CFT

[294] Scott Kirila. Combinations of L-functions and their non-coincident zeros for re( s )>1. Journal of Number Theory, 219(??):162–171, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302912.

Mehdizadeh:2021:MTS

[295] Marzieh Mehdizadeh. The multiplication table for smooth integers. Journal of Number Theory, 219(??):172–197, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302997.

Philippon:2021:NTP

[296] Patrice Philippon and Purusottam Rath. A note on trace of powers of algebraic numbers. Journal of Number Theory, 219(??):198–211, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030295X.

Garcia:2021:MST

[297] Stephan Ramon Garcia, Gabe Udell, and Jiahui Yu. Multidimensional Schinzel-type theorems for multiplicative functions near prime tuples. Journal of Number Theory, 219(??):212–227, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302894.

Bag:2021:BSD

[298] Nilanjan Bag and Igor E. Shparlinski. Bounds of some double exponential sums. Journal of Number Theory, 219(??):228–236, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303000.

Cohen:2021:PVR

[299] Stephen D. Cohen, Hariom Sharma, and Rajendra Sharma. Primitive values of rational functions at primitive elements of a finite field. Journal of Number Theory, 219(??):237–246, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302973.

Horinaga:2021:NHA

[300] Shuji Horinaga. Nearly holomorphic automorphic forms on SL _{2}. Journal of Number Theory, 219(??):247–282, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302961.

Goswami:2021:SFC

[301] Ankush Goswami and Venkata Raghu Tej Pantangi. Some formulae for coefficients in restricted q-products. Journal of Number Theory, 219(??):283–299, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302882.

Lepetit:2021:LIV

[302] Gabriel Lepetit. On the linear independence of values of G-functions. Journal of Number Theory, 219(??):300–343, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302857.

Jeon:2021:MCI

[303] Daeyeol Jeon. Modular curves with infinitely many cubic points. Journal of Number Theory, 219(??):344–355, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302870.

Aygin:2021:MPC

[304] Zafer Selcuk Aygin and Khoa D. Nguyen. Monogenic pure cubics. Journal of Number Theory, 219(??):356–367, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030281X.

Shang:2021:GSD

[305] Lei Shang and Min Wu. On the growth speed of digits in Engel expansions. Journal of Number Theory, 219(??):368–385, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302948.

Chen:2021:UDA

[306] Changhao Chen, Xiaohua Wang, and Shengyou Wen. On uniform distribution of αβ-orbits. Journal of Number Theory, 219(??):386–403, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030264X.

Manickam:2021:FFJ

[307] M. Manickam. On the first Fourier–Jacobi coefficient of Siegel modular forms of degree two. Journal of Number Theory, 219(??):404–411, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302584.

Gomez-Colunga:2021:EPS

[308] Andrés Gómez-Colunga, Charlotte Kavaler, Nathan McNew, and Mirilla Zhu. On the Erdős primitive set conjecture in function fields. Journal of Number Theory, 219(??):412–444, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302821.

Griffin:2021:MES

[309] Trevor Griffin, Nathan Kenshur, Abigail Price, Bradshaw Vandenberg-Daves, Hui Xue, and Daozhou Zhu. Monomials of Eisenstein series. Journal of Number Theory, 219(??):445–459, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302833.

Murakami:2021:CIN

[310] Yuya Murakami. Corrigendum to “Intersection numbers of modular correspondences for genus zero modular curves” [J. Number Theory 209 (April 2020) 167–194]. Journal of Number Theory, 219(??):460–465, February 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302560. See [78].

Anonymous:2021:PMa

[311] Anonymous. Pages 1–346 (March 2021). Journal of Number Theory, 220 (??):1–346, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBc

[312] Anonymous. Editorial Board. Journal of Number Theory, 220(??): ifc, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303140.

Hulse:2021:TCS

[313] Thomas A. Hulse, Chan Ieong Kuan, David Lowry-Duda, and Alexander Walker. Triple correlation sums of coefficients of cusp forms. Journal of Number Theory, 220(??):1–18, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302596.

Young:2021:ASH

[314] Paul Thomas Young. Absolute series for higher Euler constants. Journal of Number Theory, 220(??):19–33, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302158.

Pollak:2021:RIG

[315] Benjamin Pollak. Ramification in the inverse Galois problem. Journal of Number Theory, 220(??):34–60, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301979.

Hu:2021:SHD

[316] Guangwei Hu and Guangshi Lü. Sums of higher divisor functions. Journal of Number Theory, 220(??):61–74, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20302602.

Boylan:2021:OCS

[317] Matthew Boylan. U_{p}-operators and congruences for Shimura images. Journal of Number Theory, 220(??):75–93, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302626.

Dabrowski:2021:SDC

[318] Andrzej Da̧browski and Jacek Pomykała. Signatures of Dirichlet characters and elliptic curves. Journal of Number Theory, 220(??):94– 106, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302614.

Kebli:2021:NIS

[319] Salima Kebli, Omar Kihel, Jesse Larone, and Florian Luca. On the nonnegative integer solutions to the equation F_{n}±F_{m} = y^{a}. Journal of Number Theory, 220(??):107–127, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20302559.

Tian:2021:OWA

[320] Yisheng Tian. Obstructions to weak approximation for reductive groups over p-adic function fields. Journal of Number Theory, 220(??):128– 162, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302638.

Banerjee:2021:FPN

[321] Soumyarup Banerjee, Manav Batavia, Ben Kane, Muratzhan Kyranbay, Dayoon Park, Sagnik Saha, Hiu Chun So, and Piyush Varyani. Fermat’s polygonal number theorem for repeated generalized polygonal numbers. Journal of Number Theory, 220(??):163–181, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302031.

Shkredov:2021:MHB

[322] I. D. Shkredov. Modular hyperbolas and bilinear forms of Kloosterman sums. Journal of Number Theory, 220(??):182–211, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20302018.

Chen:2021:IPT

[323] Sheng Chen. Integral points on twisted Markoff surfaces. Journal of Number Theory, 220(??):212–234, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20301992.

Haas:2021:LGP

[324] Johann Haas and Morten Lüders. A local to global principle for higher zero-cycles. Journal of Number Theory, 220(??):235–265, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20301980.

Kim:2021:GEP

[325] Kwang-Seob Kim and Joachim König. On Galois extensions with prescribed decomposition groups. Journal of Number Theory, 220(??):266– 294, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301967.

Chabert:2021:PFP

[326] Jean-Luc Chabert and Emmanuel Halberstadt. From Pólya fields to Pólya groups (II) non-Galois number fields. Journal of Number Theory, 220(??):295–319, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301955.

Howe:2021:MNP

[327] Everett W. Howe. The maximum number of points on a curve of genus eight over the field of four elements. Journal of Number Theory, 220(??):320–329, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20302572.

Guzvic:2021:TGR

[328] Tomislav Gužvić. Torsion growth of rational elliptic curves in sextic number fields. Journal of Number Theory, 220(??):330–345, March 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302936.

Anonymous:2021:PAa

[329] Anonymous. Pages 1–504 (April 2021). Journal of Number Theory, 221 (??):1–504, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBd

[330] Anonymous. Editorial Board. Journal of Number Theory, 221(??): ifc, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303711.

Ginzburg:2021:TIR

[331] David Ginzburg and David Soudry. Two identities relating Eisenstein series on classical groups. Journal of Number Theory, 221(??): 1–108, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2030322X.

Han:2021:ADI

[332] Qi Han and Jingbo Liu. Algebraic differential independence regarding the Riemann ζ-function and the Euler Γ-function. Journal of Number Theory, 221(??):109–121, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20300147.

Ye:2021:EFR

[333] Rongqing Ye and Elad Zelingher. Epsilon factors of representations of finite general linear groups. Journal of Number Theory, 221(??):122– 142, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301931.

Chen:2021:MIB

[334] Rui Chen, Zhuo Chen, and Chuangqiang Hu. A modular interpretation of BBGS towers. Journal of Number Theory, 221(??):143– 173, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301943.

Bannai:2021:ESR

[335] Shinzo Bannai and Hiro o Tokunaga. Elliptic surfaces of rank one and the topology of cubic-line arrangements. Journal of Number Theory, 221(??):174–189, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2030189X.

Patankar:2021:SLP

[336] Vijay M. Patankar and C. S. Rajan. On the structure of locally potentially equivalent Galois representations. Journal of Number Theory, 221(??):190–210, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301682.

Cai:2021:IPM

[337] Yulin Cai. Integral points on the modular curves X_{0}(p). Journal of Number Theory, 221(??):211–221, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2030192X.

Goldston:2021:SGB

[338] Daniel A. Goldston, Sidney W. Graham, Apoorva Panidapu, Janos Pintz, Jordan Schettler, and Cem Y. Yıldırım. Small gaps between almost primes, the parity problem, and some conjectures of Erdős on consecutive integers II. Journal of Number Theory, 221(??):222– 231, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301888.

Wang:2021:AAF

[339] Biao Wang. Analogues of Alladi’s formula. Journal of Number Theory, 221(??):232–246, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301906.

Henniart:2021:ACF

[340] Guy Henniart and Luis Lomelí. Asai cube L-functions and the local Langlands correspondence. Journal of Number Theory, 221(??):247– 269, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301761.

Lozano-Robledo:2021:PMD

[341] Álvaro Lozano-Robledo. A probabilistic model for the distribution of ranks of elliptic curves over Q. Journal of Number Theory, 221(??): 270–338, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301773.

Smit:2021:SIN

[342] Harry Smit. L-series and isomorphisms of number fields. Journal of Number Theory, 221(??):339–357, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2030175X.

Huang:2021:UBP

[343] Keping Huang. Uniform bounds for periods of endomorphisms of varieties. Journal of Number Theory, 221(??):358–364, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303656.

Nien:2021:CTG

[344] Chufeng Nien and Lei Zhang. Converse theorem of Gauss sums. Journal of Number Theory, 221(??):365–388, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303644.

Bui:2021:TCO

[345] Hung M. Bui, Alexandra Florea, and Jonathan P. Keating. Type-I contributions to the one and two level densities of quadratic Dirichlet Lfunctions over function fields. Journal of Number Theory, 221(??):389– 423, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303590.

vanFrankenhuijsen:2021:TRA

[346] Machiel van Frankenhuijsen. Two restricted ABC conjectures. Journal of Number Theory, 221(??):424–446, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303632.

Waxman:2021:LOT

[347] Ezra Waxman. Lower order terms for the one-level density of a symplectic family of Hecke L-functions. Journal of Number Theory, 221(??): 447–483, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303589.

Marcil:2021:NVC

[348] David Marcil. Numerical verification of a conjecture of Harris and Venkatesh. Journal of Number Theory, 221(??):484–495, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302845.

Andrica:2021:PRT

[349] Dorin Andrica and George C. Ţurcaş. Pairs of rational triangles with equal symmetric invariants. Journal of Number Theory, 221(??):496– 504, April 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303681.

Anonymous:2021:PMb

[350] Anonymous. Pages 1–428 (May 2021). Journal of Number Theory, 222 (??):1–428, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBe

[351] Anonymous. Editorial Board. Journal of Number Theory, 222(??): ifc, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000299.

Machiavelo:2021:ZWS

[352] António Machiavelo and Nikolaos Tsopanidis. Zhi-Wei Sun’s 1–3– 5 conjecture and variations. Journal of Number Theory, 222(??):1– 20, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303218.

Machiavelo:2021:RZW

[353] António Machiavelo, Rogério Reis, and Nikolaos Tsopanidis. Report on Zhi-Wei Sun’s 1–3–5 conjecture and some of its refinements. Journal of Number Theory, 222(??):21–29, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2100007X.

Nathanson:2021:DSI

[354] Melvyn B. Nathanson. Dirichlet series of integers with missing digits. Journal of Number Theory, 222(??):30–37, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303243.

Ye:2021:PCC

[355] Dongxi Ye. Proof of a conjecture of Cooper. Journal of Number Theory, 222(??):38–47, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303255.

Laflamme:2021:CPA

[356] Jeanne Laflamme and Matilde Lalín. On Ceva points of (almost) equilateral triangles. Journal of Number Theory, 222(??):48–74, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000147.

Gekeler:2021:DMF

[357] Ernst-Ulrich Gekeler. On Drinfeld modular forms of higher rank V: the behavior of distinguished forms on the fundamental domain. Journal of Number Theory, 222(??):75–114, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303322.

Saito:2021:DFS

[358] Kota Saito and Yuuya Yoshida. Distributions of finite sequences represented by polynomials in Piatetski–Shapiro sequences. Journal of Number Theory, 222(??):115–156, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21000123.

Kim:2021:MGI

[359] Henry H. Kim. Minimal generators of the ideal class group. Journal of Number Theory, 222(??):157–167, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000093. See

corrigendum [538].

Ray:2021:CGR

[360] Anwesh Ray. Constructing Galois representations ramified at one prime. Journal of Number Theory, 222(??):168–180, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2100010X.

Huber:2021:RTC

[361] Timothy Huber and Dongxi Ye. Ramanujan type congruences for quotients of level 7 Klein forms. Journal of Number Theory, 222(??):181– 203, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303280.

Wang:2021:TCS

[362] Jun Wang. Towards a conjecture of Sharifi. Journal of Number Theory, 222(??):204–232, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2100024X.

Chen:2021:LBH

[363] Guohua Chen and Xiaoguang He. Lower bound for higher moments of the mixed product of twisted L-functions. Journal of Number Theory, 222(??):233–248, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000159.

Kundu:2021:AKF

[364] Debanjana Kundu. An analogue of Kida’s formula for fine Selmer groups of elliptic curves. Journal of Number Theory, 222(??):249– 261, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000160.

Hwang:2021:ASP

[365] WonTae Hwang and Kyunghwan Song. Absolutely simple polarized abelian varieties of odd Sophie Germain prime dimension over finite fields with maximal automorphism groups. Journal of Number Theory, 222(??):262–277, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000135.

Fuchs:2021:FFV

[366] Clemens Fuchs and Sebastian Heintze. A function field variant of Pillai’s problem. Journal of Number Theory, 222(??):278–292, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303334.

Bouganis:2021:SFA

[367] Thanasis Bouganis. On the standard l-function attached to quaternionic modular forms. Journal of Number Theory, 222(??):293–345, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303693.

Prevost:2021:DCR

[368] M. Prévost and T. Rivoal. Diagonal convergence of the remainder Padé approximants for the Hurwitz zeta function. Journal of Number Theory, 222(??):346–361, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303620.

Pongsriiam:2021:LAP

[369] Prapanpong Pongsriiam. Longest arithmetic progressions of palindromes. Journal of Number Theory, 222(??):362–375, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303577.

Miller:2021:CLF

[370] Stephen D. Miller, Bhargav Narayanan, and Ramarathnam Venkatesan. Coppersmith’s lattices and “focus groups”: an attack on small-exponent RSA. Journal of Number Theory, 222(??):376–392, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000226.

Gomez:2021:TRF

[371] Carlos A. Gómez, Jhonny C. Gómez, and Florian Luca. Two b-repunits in the Fibonacci sequence. Journal of Number Theory, 222(??):393– 422, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303309.

Zhai:2021:CSI

[372] Wenguang Zhai. Corrigendum to “On a sum involving the Euler function” [J. Number Theory 202 (2019) 278–297]. Journal of Number Theory, 222(??):423–425, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303267. See [8].

Banerjee:2021:CGC

[373] Pradipto Banerjee and Ranjan Bera. Corrigendum to “On a generalization of a conjecture of Grosswald” [J. Number Theory 216 (2020) 216–241]. Journal of Number Theory, 222(??):426–428, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303231. See [239].

Anonymous:2021:PJb

[374] Anonymous. Pages 1–370 (June 2021). Journal of Number Theory, 223 (??):1–370, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBf

[375] Anonymous. Editorial Board. Journal of Number Theory, 223(??): ifc, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000627.

Bailey:2021:MMZ

[376] E. C. Bailey and J. P. Keating. On the moments of the moments of ζ(1/2 + it). Journal of Number Theory, 223(??):79–100, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000184.

Languasco:2021:NVL

[377] Alessandro Languasco. Numerical verification of Littlewood’s bounds for |L(1, χ)|. Journal of Number Theory, 223(??):12–34, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000251.

Jedrzejak:2021:RFH

[378] Tomasz Jȩdrzejak. Ranks in the family of hyperelliptic Jacobians of y^{2} = x^{5} + ax. Journal of Number Theory, 223(??):35–52, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303425.

Wang:2021:TCH

[379] Zhiwei Wang. Three conjectures on P^{+}(n) and P^{+}(n + 1) hold under the Elliott–Halberstam conjecture for friable integers. Journal of Number Theory, 223(??):1–11, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000196.

Nguyen:2021:KND

[380] Evuilynn Nguyen, Juan J. Ramirez, and Matthew P. Young. The kernel of newform Dedekind sums. Journal of Number Theory, 223(??): 53–63, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303292.

Kerr:2021:PPC

[381] Bryce Kerr and Ali Mohammadi. Points on polynomial curves in small boxes modulo an integer. Journal of Number Theory, 223(??): 64–78, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303310.

Choi:2021:NZC

[382] SoYoung Choi and Bo-Hae Im. The number of zeros of certain combinations of the Eisenstein series for Γ_{0}^{+}(2). Journal of Number Theory, 223(??):101–131, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303395.

Sgobba:2021:DOI

[383] Pietro Sgobba. On the distribution of the order and index for the reductions of algebraic numbers. Journal of Number Theory, 223(??):132– 152, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303371.

Choi:2021:GPD

[384] Stephen Choi, Peter Cho-Ho Lam, and Daniel Tarnu. Gap principle of divisibility sequences of polynomials. Journal of Number Theory, 223(??):153–167, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303346.

Hofer:2021:FBC

[385] Roswitha Hofer. Finding both, the continued fraction and the Laurent series expansion of golden ratio analogs in the field of formal power series. Journal of Number Theory, 223(??):168–194, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303358.

Kleine:2021:LGB

[386] Sören Kleine. On local and global bounds for Iwasawa λ-invariants. Journal of Number Theory, 223(??):195–228, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2030336X.

Konieczny:2021:ACN

[387] Jakub Konieczny. Algorithmic classification of noncorrelated binary pattern sequences. Journal of Number Theory, 223(??):229–254, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303383.

Ferguson:2021:EFU

[388] Timothy Ferguson. Entire functions with undecidable arithmetic properties. Journal of Number Theory, 223(??):255–266, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303450.

Molina:2021:FST

[389] Santiago Molina. Finite slope triple product p-adic L-functions over totally real number fields. Journal of Number Theory, 223(??):267– 306, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303401.

Alon:2021:IPM

[390] Noga Alon, Noah Kravitz, and Matt Larson. Inverse problems for minimal complements and maximal supplements. Journal of Number Theory, 223(??):307–324, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303437.

Amdeberhan:2021:APS

[391] Tewodros Amdeberhan, Victor H. Moll, Vaishavi Sharma, and Diego Villamizar. Arithmetic properties of the sum of divisors. Journal of Number Theory, 223(??):325–349, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303449.

Biswas:2021:ACM

[392] Arindam Biswas and Jyoti Prakash Saha. On additive co-minimal pairs. Journal of Number Theory, 223(??):350–370, June 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303413.

Anonymous:2021:PJc

[393] Anonymous. Pages 1–368 (July 2021). Journal of Number Theory, 224 (??):1–368, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBg

[394] Anonymous. Editorial Board. Journal of Number Theory, 224(??): ifc, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000895.

Lee:2021:EZF

[395] Ethan S. Lee. On an explicit zero-free region for the Dedekind zeta-function. Journal of Number Theory, 224(??):307–322, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000238.

Ash:2021:SHM

[396] Avner Ash and Dan Yasaki. Steinberg homology, modular forms, and real quadratic fields. Journal of Number Theory, 224(??):323– 367, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000214.

Eum:2021:RNS

[397] Ick Sun Eum. On the representation numbers satisfying partially multiplicative relations. Journal of Number Theory, 224(??):1–12, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000421.

Mastrostefano:2021:MPS

[398] Daniele Mastrostefano. On maximal product sets of random sets. Journal of Number Theory, 224(??):13–40, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000536.

Hajdu:2021:SCC

[399] L. Hajdu, F. Luca, and R. Tijdeman. Skolem’s conjecture confirmed for a family of exponential equations, III. Journal of Number Theory, 224(??):41–49, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000500.

Su:2021:EPH

[400] Feng Su. Estimate of periods on Hirzebruch–Zagier cycles. Journal of Number Theory, 224(??):50–66, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000597.

Kim:2021:DOR

[401] Mingyu Kim. Diagonal odd-regular ternary quadratic forms. Journal of Number Theory, 224(??):67–94, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000494.

Assaf:2021:EIN

[402] Eran Assaf. Existence of invariant norms in p-adic representations of GL_{2}(F) of large weights. Journal of Number Theory, 224(??):95– 141, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000573.

Choi:2021:QT

[403] Junhwa Choi and Yongxiong Li. Quadratic twists of X_{0}(14). Journal of Number Theory, 224(??):142–164, July 2021. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000524.

Ounaies:2021:RBZ

[404] M. Ounaies, G. Rhin, and J.-M. Sac-Épée. On the relations between the zeros of a polynomial and its Mahler measure. Journal of Number Theory, 224(??):165–190, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000445.

Oliveira:2021:RPC

[405] José Alves Oliveira. Rational points on cubic, quartic and sextic curves over finite fields. Journal of Number Theory, 224(??):191– 216, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000561.

Zunar:2021:NVV

[406] Sonja Žunar. Non-vanishing of vector-valued Poincaré series. Journal of Number Theory, 224(??):217–242, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000457.

Yu:2021:MTI

[407] Han Yu. On the metric theory of inhomogeneous Diophantine approximation: an Erdős–Vaaler type result. Journal of Number Theory, 224(??):243–273, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000482.

Khaqan:2021:ECT

[408] Maryam Khaqan. Elliptic curves and Thompson’s sporadic simple group. Journal of Number Theory, 224(??):274–306, July 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000469.

Anonymous:2021:PAb

[409] Anonymous. Pages 1–408 (August 2021). Journal of Number Theory, 225 (??):1–408, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBh

[410] Anonymous. Editorial Board. Journal of Number Theory, 225(??): ifc, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001177.

Song:2021:NPF

[411] Yanbo Song. A note on primes of the form ⌊αp + β⌋. Journal of Number Theory, 225(??):1–17, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2100041X.

vonPippich:2021:KLF

[412] Anna-Maria von Pippich, Markus Schwagenscheidt, and Fabian Völz. Kronecker limit formulas for parabolic, hyperbolic and elliptic Eisenstein series via Borcherds products. Journal of Number Theory, 225(??): 18–58, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000512.

Grishkov:2021:AM

[413] A. Grishkov and D. Logachev. h^{1} ≠ h_{1} for Anderson t-motives. Journal of Number Theory, 225(??):59–89, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000585.

Alkan:2021:VCP

[414] Emre Alkan. Variations on criteria of Pólya and Turán for the Riemann hypothesis. Journal of Number Theory, 225(??):90–124, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000433.

Ban:2021:GUP

[415] Dubravka Ban, Kwangho Choiy, and David Goldberg. R-groups for unitary principal series of Spin groups. Journal of Number Theory, 225(??):125–150, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000603.

Mihailescu:2021:ILB

[416] Preda Mihăilescu. Improved lower bounds for possible solutions in the Second Case of the Fermat Last Theorem and in the Catalan Equation. Journal of Number Theory, 225(??):151–173, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2100055X.

Nordqvist:2021:WRP

[417] Jonas Nordqvist. Wildly ramified power series with large multiplicity. Journal of Number Theory, 225(??):174–197, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000548.

Chen:2021:DOV

[418] Shi-Chao Chen. On the density of the odd values of the partition function and the t-multipartition function. Journal of Number Theory, 225(??):198–213, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000743.

Dembner:2021:HCN

[419] Spencer Dembner and Vanshika Jain. Hyperelliptic curves and newform coefficients. Journal of Number Theory, 225(??):214–239, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000779.

Chirre:2021:LVA

[420] Andrés Chirre and Kamalakshya Mahatab. Large values of the argument of the Riemann zeta-function and its iterates. Journal of Number Theory, 225(??):240–259, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000822.

Earnest:2021:LCU

[421] A. G. Earnest and B. L. K. Gunawardana. Local criteria for universal and primitively universal quadratic forms. Journal of Number Theory, 225(??):260–280, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000780.

Chan:2021:ERE

[422] Wai Kiu Chan, Haochen Gao, and Han Li. Explicit result on equivalence of rational quadratic forms avoiding primes. Journal of Number Theory, 225(??):281–293, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000809.

Guo:2021:ICR

[423] Jeremy J. F. Guo. An inequality for coefficients of the real-rooted polynomials. Journal of Number Theory, 225(??):294–309, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2100086X.

Chow:2021:FP

[424] Sam Chow and Tom Slattery. On Fibonacci partitions. Journal of Number Theory, 225(??):310–326, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000846.

Luethi:2021:PRP

[425] Manuel Luethi. Primitive rational points on expanding horospheres in Hilbert modular surfaces. Journal of Number Theory, 225(??):327– 359, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000834.

Agrawal:2021:FSC

[426] Komal Agrawal and Paul Pollack. Finite sets containing nearprimitive roots. Journal of Number Theory, 225(??):360–373, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000792.

Rickards:2021:CIC

[427] James Rickards. Computing intersections of closed geodesics on the modular curve. Journal of Number Theory, 225(??):374–408, August 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303668.

Anonymous:2021:PS

[428] Anonymous. Pages 1–372 (September 2021). Journal of Number Theory, 226(??):1–372, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBi

[429] Anonymous. Editorial Board. Journal of Number Theory, 226(??):ifc, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001268.

Gomez:2021:BCM

[430] Kevin Gomez and Eric Zhu. Bounds for coefficients of the f(q) mock theta function and applications to partition ranks. Journal of Number Theory, 226(??):1–23, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21000470.

Peng:2021:OSS

[431] Deke Peng and Jiangwei Xue. Optimal spinor selectivity for quaternion Bass orders. Journal of Number Theory, 226(??):24–50, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000755.

Bravo:2021:CVG

[432] Jhon J. Bravo, Jose L. Herrera, and Florian Luca. Common values of generalized Fibonacci and Pell sequences. Journal of Number Theory, 226(??):51–71, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000767.

Akhilesh:2021:MZV

[433] P. Akhilesh. Multiple zeta values and multiple Apéry-like sums. Journal of Number Theory, 226(??):72–138, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000810.

Connes:2021:QIF

[434] Alain Connes and Caterina Consani. Quasi-inner functions and local factors. Journal of Number Theory, 226(??):139–167, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000858.

Kim:2021:CAA

[435] Byoung Du Kim. Congruences of algebraic p-adic L-functions and the Main Conjecture of Iwasawa Theory. Journal of Number Theory, 226 (??):168–212, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000871.

Gao:2021:AWC

[436] Yibo Gao and Andrew Gu. Arithmetic of weighted Catalan numbers. Journal of Number Theory, 226(??):213–242, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000986.

Shu:2021:RNJ

[437] Jie Shu. Root numbers for the Jacobian varieties of Fermat curves. Journal of Number Theory, 226(??):243–270, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21001037.

Jana:2021:JEP

[438] Subhajit Jana. Joint equidistribution on the product of the circle and the unit cotangent bundle of the modular surface. Journal of Number Theory, 226(??):271–283, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001013.

Burrell:2021:DEN

[439] Stuart A. Burrell and Han Yu. Digit expansions of numbers in different bases. Journal of Number Theory, 226(??):284–306, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000408.

Matsuzawa:2021:DRI

[440] Yohsuke Matsuzawa and Joseph H. Silverman. The distribution relation and inverse function theorem in arithmetic geometry. Journal of Number Theory, 226(??):307–357, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001086.

Asayama:2021:TPD

[441] Takuya Asayama. Torsion points of Drinfeld modules over large algebraic extensions of finitely generated function fields. Journal of Number Theory, 226(??):358–372, September 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001062.

Anonymous:2021:PO

[442] Anonymous. Pages 1–370 (October 2021). Journal of Number Theory, 227(??):1–370, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBj

[443] Anonymous. Editorial Board. Journal of Number Theory, 227(??): ifc, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001323.

Griffin:2021:QTE

[444] Michael Griffin, Ken Ono, and Wei-Lun Tsai. Quadratic twists of elliptic curves and class numbers. Journal of Number Theory, 227(??): 1–29, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000950.

Nguyen:2021:GDF

[445] David T. Nguyen. Generalized divisor functions in arithmetic progressions: I. Journal of Number Theory, 227(??):30–93, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001141.

Morton:2021:HITb

[446] Patrick Morton. The Hasse invariant of the Tate normal form E_{5} and the class number of Q(−5l). Journal of Number Theory, 227(??):94– 143, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001001.

Goldston:2021:ETC

[447] D. A. Goldston and Ade Irma Suriajaya. The error term in the Cesàro mean of the prime pair singular series. Journal of Number Theory, 227(??):144–157, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000974.

Ma:2021:PGG

[448] Qixiao Ma. Period of the generic genus g curve via degeneration. Journal of Number Theory, 227(??):158–165, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001128.

Sato:2021:ECA

[449] Kanetomo Sato. Étale cohomology of arithmetic schemes and zeta values of arithmetic surfaces. Journal of Number Theory, 227(??):166– 234, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100113X.

Mishou:2021:JUT

[450] Hidehiko Mishou. Joint universality theorem of Selberg zeta functions for principal congruence subgroups. Journal of Number Theory, 227(??):235–264, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001050.

Reutenauer:2021:QNF

[451] Christophe Reutenauer. On quadratic numbers and forms, and Markoff theory. Journal of Number Theory, 227(??):265–305, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000998.

Salami:2021:CRP

[452] Sajad Salami. Corrigendum to “The rational points on certain Abelian varieties over function fields” [J. Number Theory 195 (2019) 330–337]. Journal of Number Theory, 227(??):306–307, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001025. See [4].

Zhang:2021:SVP

[453] Long Zhang, Kejian Xu, Zhaopeng Dai, and Chaochao Sun. The shortest vector problem and tame kernels of cyclotomic fields. Journal of Number Theory, 227(??):308–329, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001153.

Adedji:2021:DP

[454] Kouèssi Norbert Adédji, Bo He, Ákos Pintér, and Alain Togbé. On the Diophantine pair {a, 3a}. Journal of Number Theory, 227(??):330– 351, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001049.

Mizuno:2021:CRC

[455] Yoshinori Mizuno. Congruences relating class numbers of quadratic orders and Zagier’s sums. Journal of Number Theory, 227(??):352– 370, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001104.

Anonymous:2021:PN

[456] Anonymous. Pages 1–390 (November 2021). Journal of Number Theory, 228(??):1–390, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBk

[457] Anonymous. Editorial Board. Journal of Number Theory, 228(??):ifc, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002407.

Maarefparvar:2021:PGS

[458] Abbas Maarefparvar. Pólya group in some real biquadratic fields. Journal of Number Theory, 228(??):1–7, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001505.

Shen:2021:SIM

[459] Zhongyan Shen and Hua He. Some identities for multiple alternating zeta values. Journal of Number Theory, 228(??):8–21, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001426.

Banks:2021:APC

[460] William Banks. Arithmetic progressions of Carmichael numbers in a reduced residue class. Journal of Number Theory, 228(??):22–28, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21000962.

Kim:2021:CPM

[461] Byoung Du (BD) Kim. Constructing plus/minus local points over Iwasawa Kummer extensions, and Iwasawa theory for supersingular reduction. Journal of Number Theory, 228(??):29–58, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001566.

Baker:2021:NDB

[462] Simon Baker. New dimension bounds for αβ sets. Journal of Number Theory, 228(??):59–72, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2100158X.

Battistoni:2021:GPI

[463] Francesco Battistoni and Giuseppe Molteni. Generalization of a Pohst’s inequality. Journal of Number Theory, 228(??):73–86, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001542.

Francis:2021:IEV

[464] Forrest J. Francis. An investigation into explicit versions of Burgess’ bound. Journal of Number Theory, 228(??):87–107, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001116.

Allouche:2021:SCP

[465] J.-P. Allouche, G.-N. Han, and J. Shallit. On some conjectures of P. Barry. Journal of Number Theory, 228(??):108–132, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001098.

Jiang:2021:VLP

[466] Tao Jiang. Variance of lattice point counting in some special shells in R^{d}. Journal of Number Theory, 228(??):133–161, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21001074.

Bhagwat:2021:GCC

[467] Chandrasheel Bhagwat and Sudipa Mondal. On the growth of cuspidal cohomology of GL _{4}. Journal of Number Theory, 228(??):162–173, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001578.

Smith:2021:NMD

[468] Hanson Smith. Non-monogenic division fields of elliptic curves. Journal of Number Theory, 228(??):174–187, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001359.

Blanco-Chacon:2021:PDM

[469] Iván Blanco-Chacón and Luis Dieulefait. Potentially diagonalizable modular lifts of large weight. Journal of Number Theory, 228(??):188–207, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001244.

Yoshikawa:2021:RMR

[470] Sho Yoshikawa. Reducible mod 105 representations and modularity of elliptic curves. Journal of Number Theory, 228(??):208–218, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001451.

Gras:2021:TSG

[471] Georges Gras. Tate–Shafarevich groups in the cyclotomic Ẑ-extension and Weber’s class number problem. Journal of Number Theory, 228 (??):219–252, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001608.

Dalal:2021:MCD

[472] Tarun Dalal and Narasimha Kumar. On mod p congruences for Drinfeld modular forms of level pm. Journal of Number Theory, 228 (??):253–275, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2100161X.

Tellez:2021:GBC

[473] Fernando Azpeitia Tellez, Christopher Powell, and Shahed Sharif. Geometry of biquadratic and cyclic cubic log-unit lattices. Journal of Number Theory, 228(??):276–293, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001463.

Chinburg:2021:SLS

[474] T. Chinburg, Q. Guignard, and C. Soulé. On the slopes of the lattice of sections of Hermitian line bundles. Journal of Number Theory, 228(??):294–341, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001475.

Gao:2021:MCV

[475] Peng Gao and Liangyi Zhao. Moments of central values of quartic Dirichlet L-functions. Journal of Number Theory, 228(??):342–358, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001591.

Ghenciu:2021:HDS

[476] Andrei E. Ghenciu and Sara Munday. The Hausdorff dimension spectrum of Renyi-like continued fractions. Journal of Number Theory, 228 (??):359–374, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001372.

Fadil:2021:IBM

[477] Lhoussain El Fadil. On integral bases and monogeneity of pure sextic number fields with non-squarefree coefficients. Journal of Number Theory, 228(??):375–389, November 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001360.

Anonymous:2021:PD

[478] Anonymous. Pages 1–498 (December 2021). Journal of Number Theory, 229(??):1–498, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2021:EBl

[479] Anonymous. Editorial Board. Journal of Number Theory, 229(??):ifc, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002493.

Al-Kateeb:2021:MGC

[480] Ala’a Al-Kateeb, Mary Ambrosino, Hoon Hong, and Eunjeong Lee. Maximum gap in cyclotomic polynomials. Journal of Number Theory, 229(??):1–15, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001530.

Voutier:2021:GCV

[481] Paul Voutier and Minoru Yabuta. The greatest common valuation of φ_{n} and ψ_{n}^{2} at points on elliptic curves. Journal of Number Theory, 229(??):16–38, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001529.

Fieker:2021:CER

[482] Claus Fieker, Tommy Hofmann, and Sogo Pierre Sanon. On the computation of the endomorphism rings of abelian surfaces. Journal of Number Theory, 229(??):39–52, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001669.

Wang:2021:PCM

[483] Liuquan Wang. Parity of coefficients of mock theta functions. Journal of Number Theory, 229(??):53–99, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001657.

Chen:2021:CSQ

[484] Yong-Gao Chen and Xiao-Hui Yan. A conjecture of Sárközy on quadratic residues. Journal of Number Theory, 229(??):100–124, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001402.

He:2021:GST

[485] Yan He and Min Ru. A generalized subspace theorem for closed subschemes in subgeneral position. Journal of Number Theory, 229 (??):125–141, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001396.

Khale:2021:BVT

[486] Tanmay Khale, Cooper O’Kuhn, Apoorva Panidapu, Alec Sun, and Shengtong Zhang. A Bombieri–Vinogradov Theorem for primes in short intervals and small sectors. Journal of Number Theory, 229 (??):142–167, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001414.

Ferguson:2021:APU

[487] Timothy Ferguson. Algebraic properties of [u042d]-functions. Journal of Number Theory, 229(??):168–178, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001487.

Chamizo:2021:CCD

[488] Fernando Chamizo and Bruno Martin. The convergence of certain Diophantine series. Journal of Number Theory, 229(??):179–198, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001633.

Matiyasevich:2021:CCC

[489] Yu. Matiyasevich. Continuous crop circles drawn by Riemann’s zeta function. Journal of Number Theory, 229(??):199–217, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001670.

Drmota:2021:EEW

[490] Michael Drmota and Johann Verwee. Effective Erdős–Wintner theorems for digital expansions. Journal of Number Theory, 229(??):218– 260, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100144X.

Hoshi:2021:PRT

[491] Yuichiro Hoshi. Pseudo-rigid p-torsion finite flat commutative group schemes. Journal of Number Theory, 229(??):261–276, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001682.

Zanello:2021:DPO

[492] Fabrizio Zanello. Deducing the positive odd density of p ( n ) from that of a multipartition function: an unconditional proof. Journal of Number Theory, 229(??):277–281, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001785.

Ray:2021:RLT

[493] Anwesh Ray. A refined lifting theorem for supersingular Galois representations. Journal of Number Theory, 229(??):282–313, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100189X.

Konig:2021:CGR

[494] Joachim König. Covering groups of M_{22} as regular Galois groups over Q. Journal of Number Theory, 229(??):314–327, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21001840.

Singh:2021:NDR

[495] Ajit Singh and Rupam Barman. New density results and congruences for Andrews’ singular overpartitions. Journal of Number Theory, 229 (??):328–341, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001943.

Hatley:2021:CAS

[496] Jeffrey Hatley and Antonio Lei. Comparing anticyclotomic Selmer groups of positive coranks for congruent modular forms — Part II. Journal of Number Theory, 229(??):342–363, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001864.

Kojima:2021:MFW

[497] Hisashi Kojima and Hiroshi Sakata. On the modular forms of weight 1/2 over algebraic number fields. Journal of Number Theory, 229 (??):364–385, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001499.

Tougma:2021:SQB

[498] Charles Wend-Waoga Tougma. Some questions on biquadratic Pólya fields. Journal of Number Theory, 229(??):386–398, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303462.

Wang:2021:FSG

[499] Liuquan Wang. Farey sequence and Graham’s conjectures. Journal of Number Theory, 229(??):399–404, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303474.

Hatton:2021:KDM

[500] Richard Hatton. Kolyvagin derivatives of modular points on elliptic curves. Journal of Number Theory, 229(??):405–431, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20303498.

Reis:2021:ACP

[501] Lucas Reis. Arithmetic constraints of polynomial maps through discrete logarithms. Journal of Number Theory, 229(??):432–443, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303516.

Bringmann:2021:ADR

[502] Kathrin Bringmann, Chris Jennings-Shaffer, and Karl Mahlburg. The asymptotic distribution of the rank for unimodal sequences. Journal of Number Theory, 229(??):444–462, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303504.

Cohen:2021:PCH

[503] Alex Cohen. Poissonian correlation of higher order differences. Journal of Number Theory, 229(??):463–486, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303541.

Singh:2021:CEQ

[504] Ajit Singh and Rupam Barman. Certain eta-quotients and arithmetic density of Andrews’ singular overpartitions. Journal of Number Theory, 229(??):487–498, December 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303528.

Anonymous:2022:EBa

[505] Anonymous. Editorial Board. Journal of Number Theory, 230(??): ifc, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003115.

Anonymous:2022:PFJ

[506] Anonymous. Participants at the first JNT Biennial Conference, July 22– 26, 2019. Journal of Number Theory, 230(??):1–2, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002912.

Goldfeld:2022:P

[507] Dorian Goldfeld, Ambrus Pál, Federico Pellarin, and Lejla Smajlović. Preface. Journal of Number Theory, 230(??):3–4, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002882.

Beuzart-Plessis:2022:MPF

[508] Raphaël Beuzart-Plessis. Multiplicities and Plancherel formula for the space of nondegenerate Hermitian matrices. Journal of Number Theory, 230(??):5–63, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001815.

Cadoret:2022:UCE

[509] Anna Cadoret. Ultraproduct coefficients in étale cohomology — a survey. Journal of Number Theory, 230(??):64–88, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21001876.

Darmon:2022:AIM

[510] Henri Darmon and Jan Vonk. Arithmetic intersections of modular geodesics. Journal of Number Theory, 230(??):89–111, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21000202.

Li:2022:VTS

[511] Wen-Wei Li. Variations on themes of Sato: a survey. Journal of Number Theory, 230(??):112–145, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X20302900.

Dimitrov:2022:CRB

[512] Vesselin Dimitrov, Ziyang Gao, and Philipp Habegger. A consequence of the relative Bogomolov conjecture. Journal of Number Theory, 230(??):146–160, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001621.

Burungale:2022:EPG

[513] Ashay Burungale and Ye Tian. The even parity Goldfeld conjecture: Congruent number elliptic curves. Journal of Number Theory, 230(??):161–195, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001773.

Liu:2022:TCA

[514] Yifeng Liu. Theta correspondence for almost unramified representations of unitary groups. Journal of Number Theory, 230(??):196–224, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001554.

Munshi:2022:SCS

[515] Ritabrata Munshi. On a shifted convolution sum problem. Journal of Number Theory, 230(??):225–232, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21000172.

Yafaev:2022:NCA

[516] Andrei Yafaev. Non-commutative analytic subgroup theorem. Journal of Number Theory, 230(??):233–237, January 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001517.

Anonymous:2022:PF

[517] Anonymous. Pages 1–494 (February 2022). Journal of Number Theory, 231(??):1–494, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBb

[518] Anonymous. Editorial Board. Journal of Number Theory, 231(??):ifc, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003371.

Byeon:2022:CNR

[519] Dongho Byeon and Jigu Kim. Class numbers of real quadratic fields. Journal of Number Theory, 231(??):1–47, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X20303486.

Cho:2022:MEC

[520] Bumkyu Cho. Modular equations for congruence subgroups of genus zero (II). Journal of Number Theory, 231(??):48–79, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030353X.

Schroer:2022:ECR

[521] Stefan Schröer. Elliptic curves over the rational numbers with semiabelian reduction and two-division points. Journal of Number Theory, 231(??):80–101, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303565.

Mallesham:2022:RIM

[522] Kummari Mallesham, Gyan Prakash, and D. S. Ramana. Representation of integers as monochromatic sums of squares of primes. Journal of Number Theory, 231(??):102–119, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303619.

Komiyama:2022:STF

[523] Nao Komiyama. On shuffle-type functional relations of desingularized multiple zeta-functions. Journal of Number Theory, 231(??):120–138, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303553.

Azzouz:2022:SLD

[524] Tinhinane A. Azzouz. Spectrum of a linear differential equation over a field of formal power series. Journal of Number Theory, 231(??):139– 157, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20303607.

Betts:2022:VTN

[525] L. Alexander Betts. Variation of Tamagawa numbers of Jacobians of hyperelliptic curves with semistable reduction. Journal of Number Theory, 231(??):158–213, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21000019.

ElHabibi:2022:PEN

[526] Abdelaziz El Habibi and Yasushi Mizusawa. On pro-p-extensions of number fields with restricted ramification over intermediate Z_{p}-extensions. Journal of Number Theory, 231(??):214–238, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000081.

Sekigawa:2022:RGC

[527] Ryutaro Sekigawa. Rikuna’s generic cyclic polynomial and the monogenity. Journal of Number Theory, 231(??):239–250, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21000111.

Wakabayashi:2022:EVB

[528] Yasuhiro Wakabayashi. An effective version of Belyĭ’s theorem in positive characteristic. Journal of Number Theory, 231(??):251–268, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001797.

Chen:2022:ANC

[529] Shih-Yu Chen. Algebraicity of the near central non-critical values of symmetric fourth L-functions for Hilbert modular forms. Journal of Number Theory, 231(??):269–315, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001803.

Jaulent:2022:CAG

[530] Jean-François Jaulent. Capitulation abélienne des groupes de classes de rayons. (French) [Abelian capitulation of ray class groups]. Journal of Number Theory, 231(??):316–332, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001839.

Taylor:2022:SCK

[531] Noah Taylor. On seven conjectures of Kedlaya and Medvedovsky. Journal of Number Theory, 231(??):333–348, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001827.

Hare:2022:GFS

[532] Kevin Hare and J. C. Saunders. Generalised Fibonacci sequences constructed from balanced words. Journal of Number Theory, 231 (??):349–377, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001931.

Lim:2022:RSG

[533] Donghyeok Lim. On p-rationality of Q(ζ_{2ℓ+1})^{+} for Sophie Germain primes ℓ. Journal of Number Theory, 231(??):378–400, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2100192X.

Guo:2022:PRC

[534] Ji Guo, Chia-Liang Sun, and Julie Tzu-Yueh Wang. On Pisot’s d-th root conjecture for function fields and related GCD estimates. Journal of Number Theory, 231(??):401–432, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21001918.

Shkredov:2022:SAG

[535] Ilya D. Shkredov. On some applications of GCD sums to arithmetic combinatorics. Journal of Number Theory, 231(??):433–450, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001906.

Koyama:2022:FES

[536] Shin ya Koyama and Nobushige Kurokawa. Functional equations for Selberg zeta functions with Tate motives. Journal of Number Theory, 231(??):451–463, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001955.

Simonic:2022:EEU

[537] Aleksander Simonič. On explicit estimates for S(t), S_{1}(t), and ζ(1/2+it) under the Riemann Hypothesis. Journal of Number Theory, 231(??):464– 491, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001980.

Kim:2022:CMG

[538] Henry H. Kim. Corrigendum to “Minimal generators of the ideal class group” [J. Number Theory 222 (2021) 157–167]. Journal of Number Theory, 231(??):492–494, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21001888. See [359].

Anonymous:2022:EBc

[539] Anonymous. Editorial Board. Journal of Number Theory, 232(??): ifc, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003504.

Pal:2022:FDG

[540] Ambrus Pál and Federico Pellarin. Foreword to David Goss’ JNT memorial volume. Journal of Number Theory, 232(??):1–3, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003474.

Gekeler:2022:DMFa

[541] Ernst-Ulrich Gekeler. On Drinfeld modular forms of higher rank II. Journal of Number Theory, 232(??):4–32, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303494.

Gekeler:2022:DMFb

[542] Ernst-Ulrich Gekeler. On Drinfeld modular forms of higher rank IV: Modular forms with level. Journal of Number Theory, 232(??):33– 74, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301593.

Hattori:2022:CDM

[543] Shin Hattori. On the compactification of the Drinfeld modular curve of level Γ_{1}^{Δ}(n). Journal of Number Theory, 232(??):75–100, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X20302316.

Choi:2022:BSW

[544] SoYoung Choi. A basis for the space of weakly holomorphic Drinfeld modular forms for GL _{2} ( A ). Journal of Number Theory, 232(??):101–117, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301258.

Kuhn:2022:FED

[545] Nikolas Kuhn and Richard Pink. Finding endomorphisms of Drinfeld modules. Journal of Number Theory, 232(??):118–154, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001852.

Garai:2022:RER

[546] Sumita Garai and Mihran Papikian. Reprint of: Endomorphism rings of reductions of Drinfeld modules. Journal of Number Theory, 232(??): 155–176, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003486.

Maurischat:2022:PMS

[547] Andreas Maurischat. Periods of t-modules as special values. Journal of Number Theory, 232(??):177–203, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1830297X.

Green:2022:TPR

[548] Nathan Green. Tensor powers of rank 1 Drinfeld modules and periods. Journal of Number Theory, 232(??):204–241, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301210.

Fang:2022:ESV

[549] Jiangxue Fang. Equivariant special L-values of abelian t-modules. Journal of Number Theory, 232(??):242–260, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303408.

Bandini:2022:FIC

[550] Andrea Bandini, Francesc Bars, and Edoardo Coscelli. Fitting ideals of class groups in Carlitz–Hayes cyclotomic extensions. Journal of Number Theory, 232(??):261–282, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19300228.

Rad:2022:CMF

[551] Esmail Arasteh Rad and Urs Hartl. Category of C-motives over finite fields. Journal of Number Theory, 232(??):283–316, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2030202X.

Papanikolas:2022:HPS

[552] Matthew A. Papanikolas. Hyperderivative power sums, Vandermonde matrices, and Carlitz multiplication coefficients. Journal of Number Theory, 232(??):317–354, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2030367X.

Pellarin:2022:THS

[553] F. Pellarin and R. Perkins. On twisted A-harmonic sums and Carlitz finite zeta values. Journal of Number Theory, 232(??):355–378, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303299.

Huang:2022:MIS

[554] Wei-Cheng Huang. A t-motivic interpretation of shuffle relations for multizeta values. Journal of Number Theory, 232(??):379–405, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301891.

Harada:2022:MPB

[555] Ryotaro Harada. On multi-poly-Bernoulli–Carlitz numbers. Journal of Number Theory, 232(??):406–422, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1830341X.

Jeong:2022:EF

[556] Sangtae Jeong. Ergodic functions over Z_{p}. Journal of Number Theory, 232(??):423–479, March 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001645.

Anonymous:2022:PA

[557] Anonymous. Pages 1-498 (April 2022). Journal of Number Theory, 233 (??):1–498, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBd

[558] Anonymous. Editorial Board. Journal of Number Theory, 233(??): ifc, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003565.

Haan:2022:LGG

[559] Jaeho Haan. The local Gan–Gross–Prasad conjecture for U(n + 1) × U(n): a non-generic case. Journal of Number Theory, 233(??):1– 23, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002006.

Aubert:2022:CDB

[560] Anne-Marie Aubert and Roger Plymen. Comparison of the depths on both sides of the local Langlands correspondence for Weil-restricted groups (with an appendix by Jessica Fintzen). Journal of Number Theory, 233(??):24–58, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21001992.

Murty:2022:NNP

[561] M. Ram Murty, V. Kumar Murty, and Sudhir Pujahari. On the normal number of prime factors of sums of Fourier coefficients of eigenforms. Journal of Number Theory, 233(??):59–77, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002018.

Friedlander:2022:EZG

[562] J. B. Friedlander, D. A. Goldston, H. Iwaniec, and A. I. Suriajaya. Exceptional zeros and the Goldbach problem. Journal of Number Theory, 233(??):78–86, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002031.

Hasegawa:2022:LTE

[563] Takehiro Hasegawa. Logarithmic-type and exponential-type hypergeometric functions for function fields. Journal of Number Theory, 233(??): 87–111, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100202X.

Ghioca:2022:OTR

[564] Dragos Ghioca and Igor E. Shparlinski. Order of torsion for reduction of linearly independent points for a family of Drinfeld modules. Journal of Number Theory, 233(??):112–125, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002055.

Roy:2022:PFC

[565] Manami Roy. Paramodular forms coming from elliptic curves. Journal of Number Theory, 233(??):126–157, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002067.

DelCorso:2022:HFE

[566] Ilaria Del Corso, Fabio Ferri, and Davide Lombardo. How far is an extension of p-adic fields from having a normal integral basis? Journal of Number Theory, 233(??):158–197, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21002079.

Zou:2022:HDM

[567] Yuru Zou, Jian Lu, and Vilmos Komornik. Hausdorff dimension of multiple expansions. Journal of Number Theory, 233(??):198– 227, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002080.

Rapinchuk:2022:SFR

[568] Andrei S. Rapinchuk and Igor A. Rapinchuk. Some finiteness results for algebraic groups and unramified cohomology over higher-dimensional fields. Journal of Number Theory, 233(??):228–260, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002092.

Lu:2022:GLP

[569] Hengfei Lu. The generalized linear periods. Journal of Number Theory, 233(??):261–284, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002134.

Gonzalez:2022:DP

[570] Josep González. On the p-th division polynomial. Journal of Number Theory, 233(??):285–300, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21002146.

Buraczewski:2022:CLT

[571] Dariusz Buraczewski, Alexander Iksanov, and Alexander Marynych. Central limit theorem for the least common multiple of a uniformly sampled m-tuple of integers. Journal of Number Theory, 233(??):301– 336, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002109.

Kakuhama:2022:LDG

[572] Hirotaka Kakuhama. On the local doubling γ-factor for classical groups over function fields. Journal of Number Theory, 233(??):337– 369, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002122.

Leung:2022:USD

[573] Ka Hin Leung and Bernhard Schmidt. Unique sums and differences in finite Abelian groups. Journal of Number Theory, 233(??):370– 388, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002110.

Aryan:2022:ELG

[574] Farzad Aryan. On an extension of the Landau–Gonek formula. Journal of Number Theory, 233(??):389–404, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002158.

Serban:2022:FRA

[575] Vlad Serban. A finiteness result for p-adic families of Bianchi modular forms. Journal of Number Theory, 233(??):405–431, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2100216X.

Nakamura:2022:FEZ

[576] Takashi Nakamura. Functional equation and zeros on the critical line of the quadrilateral zeta function. Journal of Number Theory, 233(??): 432–455, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002183.

Kim:2022:RQF

[577] Byeong Moon Kim, Myung-Hwan Kim, and Dayoon Park. Real quadratic fields admitting universal lattices of rank 7. Journal of Number Theory, 233(??):456–466, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002195.

Halbeisen:2022:PPP

[578] Lorenz Halbeisen and Norbert Hungerbühler. Pairing Pythagorean pairs. Journal of Number Theory, 233(??):467–480, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002171.

Broll:2022:MEV

[579] Christopher Broll, Michael P. Knapp, Jessica A. Kuiper, Paulo H. A. Rodrigues, and Daiane Veras. More exact values of the function Γ^{∗}(k). Journal of Number Theory, 233(??):481–497, April 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002043.

Anonymous:2022:PM

[580] Anonymous. Pages 1–502 (May 2022). Journal of Number Theory, 234 (??):1–502, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBe

[581] Anonymous. Editorial Board. Journal of Number Theory, 234(??): ifc, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000154.

Moerman:2022:VC

[582] Boaz Moerman. L-values for conductor 32. Journal of Number Theory, 234(??):1–30, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003279.

Chalupka:2022:CGF

[583] Karolina Chałupka, Andrzej Da̧browski, and Gökhan Soydan. On a class of generalized Fermat equations of signature (2, 2n, 3). Journal of Number Theory, 234(??):153–178, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002201.

ODesky:2022:FIV

[584] Andrew O’Desky. Functions with integer-valued divided differences. Journal of Number Theory, 234(??):179–199, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002237.

Choi:2022:ZPF

[585] SoYoung Choi and Bo-Hae Im. On the zeros of period functions associated to the Eisenstein series for Γ_{0}^{+}(N). Journal of Number Theory, 234(??):200–239, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002213.

Nguyen:2022:WSC

[586] Liem Nguyen. On Weil sums, conjectures of Helleseth, and Niho exponents. Journal of Number Theory, 234(??):240–258, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002225.

Schumacher:2022:SFK

[587] Raphael Schumacher. Subconvexity for GL_{3}(R) L-functions: the key identity via integral representations. Journal of Number Theory, 234(??): 259–284, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002249.

Liu:2022:GFC

[588] Baiying Liu and Qing Zhang. Gamma factors and converse theorems for classical groups over finite fields. Journal of Number Theory, 234(??):285–332, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002250.

Lutsko:2022:LRC

[589] Christopher Lutsko. Long-range correlations of sequences modulo 1. Journal of Number Theory, 234(??):333–348, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002274.

Hanson:2022:CFA

[590] Michael Hanson and Marie Jameson. Cusp forms as p-adic limits. Journal of Number Theory, 234(??):349–362, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002262.

Isham:2022:LBN

[591] Kelly Isham. Lower bounds for the number of subrings in Z^{n}. Journal of Number Theory, 234(??):363–390, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002286.

Sachpazis:2022:MFS

[592] Stelios Sachpazis. Multiplicative functions on shifted primes. Journal of Number Theory, 234(??):391–403, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002298.

Thackeray:2022:BFH

[593] Henry (Maya) Robert Thackeray. A BSD formula for high-weight modular forms. Journal of Number Theory, 234(??):404–447, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21002304.

Bayless:2022:HMC

[594] Jonathan Bayless, Paul Kinlaw, and Jared Duker Lichtman. Higher Mertens constants for almost primes. Journal of Number Theory, 234(??): 448–475, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002316.

Kundu:2022:AIR

[595] Debanjana Kundu and Anwesh Ray. Anticyclotomic μ-invariants of residually reducible Galois representations. Journal of Number Theory, 234(??):476–498, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002328.

Andersen:2022:OSP

[596] Nickolas Andersen, Spencer Durham, Michael J. Griffin, Jonathan Hales, Paul Jenkins, Ryan Keck, Hankun Ko, Grant Molnar, Eric Moss, Pace P. Nielsen, Kyle Niendorf, Vandy Tombs, Merrill Warnick, and Dongsheng Wu. Odd, spoof perfect factorizations. Journal of Number Theory, 234(??):31–47, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002845.

Frendreiss:2022:CPP

[597] William Frendreiss, Jennifer Gao, Austin Lei, Amy Woodall, Hui Xue, and Daozhou Zhu. Cuspidal projections of products of Eisenstein series. Journal of Number Theory, 234(??):48–73, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002997.

Trbovic:2022:TNE

[598] Antonela Trbović. Tamagawa numbers of elliptic curves with prescribed torsion subgroup or isogeny. Journal of Number Theory, 234(??):74– 94, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003218.

Andrews:2022:MPA

[599] George E. Andrews and Peter Paule. MacMahon’s partition analysis XIII: Schmidt type partitions and modular forms. Journal of Number Theory, 234(??):95–119, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003243.

Kolossvary:2022:DBN

[600] István B. Kolossváry and István T. Kolossváry. Distance between natural numbers based on their prime signature. Journal of Number Theory, 234(??):120–139, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003267.

Hejda:2022:TQF

[601] Tomáš Hejda and Vítězslav Kala. Ternary quadratic forms representing a given arithmetic progression. Journal of Number Theory, 234(??): 140–152, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003310.

Prado:2022:CLO

[602] Bernardo Bianco Prado and Kim Klinger-Logan. Corrigendum to “Linear operators, the Hurwitz zeta function and Dirichlet L-functions” [J. Number Theory 217 (2020) 422–442]. Journal of Number Theory, 234(??):499–502, May 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002675. See [268].

Anonymous:2022:PJa

[603] Anonymous. Pages 1–502 (June 2022). Journal of Number Theory, 235 (??):1–502, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBf

[604] Anonymous. Editorial Board. Journal of Number Theory, 235(??): ifc, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000312.

Ginzburg:2022:DDC

[605] David Ginzburg and David Soudry. Double descent in classical groups. Journal of Number Theory, 235(??):1–156, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2100305X.

Florez:2022:NRS

[606] Jorge Flórez. The norm residue symbol for higher local fields. Journal of Number Theory, 235(??):157–218, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002341.

Hasanalizade:2022:CZR

[607] Elchin Hasanalizade, Quanli Shen, and Peng-Jie Wong. Counting zeros of the Riemann zeta function. Journal of Number Theory, 235(??): 219–241, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100233X.

Imamoglu:2022:KSF

[608] Ö. Imamoğlu, A. Lägeler, and Á. Tóth. The Katok–Sarnak formula for higher weights. Journal of Number Theory, 235(??):242– 274, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002353.

Keith:2022:PCC

[609] William J. Keith and Fabrizio Zanello. Parity of the coefficients of certain eta-quotients. Journal of Number Theory, 235(??):275– 304, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002365.

Sedunova:2022:IBQ

[610] Alisa Sedunova. Intersections of binary quadratic forms in primes and the paucity phenomenon. Journal of Number Theory, 235(??):305– 327, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002377.

Bubboloni:2022:CPJ

[611] Daniela Bubboloni and Florian Luca. Coprime partitions and Jordan totient functions. Journal of Number Theory, 235(??):328–357, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002389.

Kim:2022:CGM

[612] Elena Kim and Steven J. Miller. Constructions of generalized MSTD sets in higher dimensions. Journal of Number Theory, 235(??):358– 381, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002572.

Wu:2022:HML

[613] Xiaorun Wu. Higher moments for lattice point discrepancy of convex domains and annuli. Journal of Number Theory, 235(??):382– 400, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002584.

Liu:2022:SC

[614] Dan Liu and Yankun Sui. Sign changes of Δ_{(1)}(x). Journal of Number Theory, 235(??):401–419, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002596.

Baier:2022:LSS

[615] Stephan Baier and Rajneesh Kumar Singh. The large sieve with square moduli in function fields. Journal of Number Theory, 235(??):420– 464, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002602.

Saidi:2022:LGP

[616] Mohamed Saïdi. A local-global principle for torsors under geometric prosolvable fundamental groups II. Journal of Number Theory, 235(??):465–483, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002614.

Chipchakov:2022:FDO

[617] Ivan D. Chipchakov. Fields of dimension one algebraic over a global or local field need not be of type C_{1}. Journal of Number Theory, 235(??):484–501, June 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002638.

Anonymous:2022:PJb

[618] Anonymous. Pages 1–492 (July 2022). Journal of Number Theory, 236 (??):1–492, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBg

[619] Anonymous. Editorial Board. Journal of Number Theory, 236(??): ifc, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000397.

Hochfilzer:2022:RAF

[620] Leonhard Hochfilzer and Thomas Oliver. Ratios of Artin L-functions. Journal of Number Theory, 236(??):1–40, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002626.

Kazalicki:2022:DTK

[621] Matija Kazalicki and Bartosz Naskrȩcki. Diophantine triples and K3 surfaces. Journal of Number Theory, 236(??):41–70, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2100264X.

Karabulut:2022:BHF

[622] Cihan Karabulut. From binary Hermitian forms to parabolic cocycles of Euclidean Bianchi groups. Journal of Number Theory, 236(??):71– 115, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002651.

Weiss:2022:SAP

[623] Christian Weiß and Thomas Skill. Sequences with almost Poissonian pair correlations. Journal of Number Theory, 236(??):116– 127, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002663.

Ye:2022:BTH

[624] Yangbo Ye. Bounds toward Hypothesis S for cusp forms. Journal of Number Theory, 236(??):128–143, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21002687.

Akiyama:2022:AFB

[625] Shigeki Akiyama. Asymptotic formula for balanced words. Journal of Number Theory, 236(??):144–159, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002699.

Perez:2022:NSS

[626] Mariana Pérez and Melina Privitelli. On the number of solutions of systems of certain diagonal equations over finite fields. Journal of Number Theory, 236(??):160–187, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002705.

Hochfilzer:2022:VSS

[627] Leonhard Hochfilzer. Variance of sums in short intervals and L -functions in F_{q}[t]. Journal of Number Theory, 236(??):188–213, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002717.

Xu:2022:GAF

[628] Chenran Xu. General asymptotic formula of Fourier coefficients of cusp forms over sum of two squares. Journal of Number Theory, 236(??): 214–229, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002729.

Ferguson:2022:UAP

[629] Timothy Ferguson. Undecidable arithmetic properties of solutions of Fredholm integral equations. Journal of Number Theory, 236(??):230– 244, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002730.

Languasco:2022:UEE

[630] Alessandro Languasco and Timothy S. Trudgian. Uniform effective estimates for |L(1, χ)|. Journal of Number Theory, 236(??):245– 260, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002742.

Hilberdink:2022:HOP

[631] Titus Hilberdink. How often is d(n) a power of a given integer? Journal of Number Theory, 236(??):261–279, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002754.

Vukusic:2022:CTM

[632] Ingrid Vukusic and Volker Ziegler. Consecutive tuples of multiplicatively dependent integers. Journal of Number Theory, 236(??):280– 307, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002766.

Zhou:2022:SFV

[633] Guang-Liang Zhou and Yuchen Ding. On the square-free values of the polynomial x^{2} + y^{2} + z^{2} + k. Journal of Number Theory, 236(??):308– 322, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100278X.

Lee:2022:AES

[634] Yoonbok Lee. An asymptotic expansion of Selberg’s central limit theorem near the critical line. Journal of Number Theory, 236(??):323– 333, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21002778.

Ma:2022:GBD

[635] J. Ma, J. Wu, and F. Zhao. On a generalisation of Bordellès– Dai–Heyman–Pan–Shparlinski’s conjecture. Journal of Number Theory, 236(??):334–348, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002791.

Li:2022:EOJ

[636] Yan-Bin Li, Nils-Peter Skoruppa, and Haigang Zhou. Eichler orders and Jacobi forms of squarefree level. Journal of Number Theory, 236(??):349–387, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002808.

Pritsker:2022:HAI

[637] Igor E. Pritsker. House of algebraic integers symmetric about the unit circle. Journal of Number Theory, 236(??):388–403, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002821.

Essouabri:2022:MVM

[638] Driss Essouabri, Christoper Salinas Zavala, and László Tóth. Mean values of multivariable multiplicative functions and applications to the average number of cyclic subgroups and multivariable averages associated with the LCM function. Journal of Number Theory, 236(??): 404–442, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100281X.

Das:2022:PPS

[639] Pranabesh Das, Pallab Kanti Dey, Angelos Koutsianas, and Nikos Tzanakis. Perfect powers in sum of three fifth powers. Journal of Number Theory, 236(??):443–462, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002857.

Mitankin:2022:GOF

[640] Vladimir Mitankin and Cecília Salgado. Genus one fibrations and vertical Brauer elements on del Pezzo surfaces of degree 4. Journal of Number Theory, 236(??):463–478, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002833.

Gaal:2022:CNG

[641] István Gaál and Michael E. Pohst. On calculating the number N(D) of global cubic fields F of given discriminant D. Journal of Number Theory, 236(??):479–491, July 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003346.

Anonymous:2022:EBh

[642] Anonymous. Editorial Board. Journal of Number Theory, 237(??): ifc, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000610.

Pellarin:2022:NDT

[643] Federico Pellarin. New developments in the theory of modular forms over function fields. Journal of Number Theory, 237(??):1–2, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000361.

Balakrishnan:2022:VLC

[644] Jennifer S. Balakrishnan, William Craig, and Ken Ono. Variations of Lehmer’s Conjecture for Ramanujan’s tau-function. Journal of Number Theory, 237(??):3–14, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301189.

Cojocaru:2022:GDE

[645] Alina Carmen Cojocaru and Mihran Papikian. The growth of the discriminant of the endomorphism ring of the reduction of a rank 2 generic Drinfeld module. Journal of Number Theory, 237(??):15– 39, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21001438.

Demangos:2022:MIR

[646] L. Demangos and T. M. Gendron. Modular invariant of rank 1 Drinfeld modules and class field generation. Journal of Number Theory, 237(??):40–66, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303279.

Armana:2022:STB

[647] Cécile Armana and Fu-Tsun Wei. Sturm-type bounds for modular forms over function fields. Journal of Number Theory, 237(??):67– 98, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20302171.

Chen:2022:SAG

[648] Chien-Hua Chen. Surjectivity of the adelic Galois representation associated to a Drinfeld module of rank 3. Journal of Number Theory, 237(??):99–123, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20301918.

Bandini:2022:DCF

[649] Andrea Bandini and Maria Valentino. Drinfeld cusp forms: oldforms and newforms. Journal of Number Theory, 237(??):124–144, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301220.

Garai:2022:CER

[650] Sumita Garai and Mihran Papikian. Computing endomorphism rings and Frobenius matrices of Drinfeld modules. Journal of Number Theory, 237(??):145–164, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19304184.

Pal:2022:ACA

[651] Ambrus Pál. p-adic cohomology and arithmetic applications: Conference in Banff, 2017. Journal of Number Theory, 237(??):165–166, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000592.

Agrawal:2022:DOI

[652] Shishir Agrawal. Deformations of overconvergent isocrystals on the projective line. Journal of Number Theory, 237(??):167–241, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19304135.

Ertl:2022:NPV

[653] Veronika Ertl. A new proof of a vanishing result due to Berthelot, Esnault, and Rülling. Journal of Number Theory, 237(??):242–256, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301271.

Gregory:2022:CRD

[654] Oliver Gregory. Crystals of relative displays and Calabi–Yau threefolds. Journal of Number Theory, 237(??):257–284, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000051.

Gros:2022:TDP

[655] Michel Gros, Bernard Le Stum, and Adolfo Quirós. Twisted divided powers and applications. Journal of Number Theory, 237(??):285– 331, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930085X.

Huyghe:2022:ADO

[656] Christine Huyghe, Tobias Schmidt, and Matthias Strauch. Arithmetic differential operators with congruence level structures: First results and examples. Journal of Number Theory, 237(??):332–352, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302847.

Kedlaya:2022:NI

[657] Kiran S. Kedlaya. Notes on isocrystals. Journal of Number Theory, 237(??):353–394, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200004X.

Lazda:2022:NED

[658] Christopher Lazda. A note on effective descent for overconvergent isocrystals. Journal of Number Theory, 237(??):395–410, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303464.

Upton:2022:DSA

[659] James Upton. Discriminant-stability in p-adic Lie towers of number fields. Journal of Number Theory, 237(??):411–428, August 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302440.

Anonymous:2022:PS

[660] Anonymous. Pages 1–1106 (September 2022). Journal of Number Theory, 238(??):1–1106, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBi

[661] Anonymous. Editorial Board. Journal of Number Theory, 238(??):ifc, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000701.

Micheli:2022:LGP

[662] Giacomo Micheli, Severin Schraven, and Violetta Weger. A local to global principle for expected values. Journal of Number Theory, 238(??):1–16, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002869.

Ganguly:2022:KTA

[663] Arijit Ganguly. Khintchine’s theorem for affine hyperplanes in positive characteristic. Journal of Number Theory, 238(??):17–36, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21002924.

Kim:2022:MUC

[664] Kyoungmin Kim, Jeongwon Lee, and Byeong-Kweon Oh. Minimal universality criterion sets on the representations of binary quadratic forms. Journal of Number Theory, 238(??):37–59, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002948.

Nomoto:2022:RCE

[665] Keiichiro Nomoto. The rank of a CM elliptic curve and a recurrence formula. Journal of Number Theory, 238(??):60–81, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21002936.

Ray:2022:RAV

[666] Jishnu Ray. Rigid analytic vectors of crystalline representations arising in p-adic Langlands. Journal of Number Theory, 238(??):82–105, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100295X.

Ei:2022:ELC

[667] Hiromi Ei, Hitoshi Nakada, and Rie Natsui. On the existence of the Legendre constants for some complex continued fraction expansions over imaginary quadratic fields. Journal of Number Theory, 238 (??):106–132, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002973.

Nathanson:2022:BCA

[668] Melvyn B. Nathanson. The Bose–Chowla argument for Sidon sets. Journal of Number Theory, 238(??):133–146, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21002985.

Jha:2022:MSG

[669] Somnath Jha, Tathagata Mandal, and Sudhanshu Shekhar. Multiplicities in Selmer groups and root numbers of Artin twists. Journal of Number Theory, 238(??):147–182, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003024.

Chakraborty:2022:MNN

[670] Kalyan Chakraborty and Krishnarjun Krishnamoorthy. On moments of non-normal number fields. Journal of Number Theory, 238(??):183–196, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003000.

Tomita:2022:AEP

[671] Takuki Tomita. The absolute Euler product representation of the absolute zeta function for a torsion free Noetherian F_{1}-scheme. Journal of Number Theory, 238(??):197–220, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003012.

Franco:2022:PHC

[672] Jorge Duque Franco. Periods of Hodge cycles and special values of the Gauss’ hypergeometric function. Journal of Number Theory, 238 (??):221–252, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003048.

Zhao:2022:POS

[673] Kevin Zhao and Qinghai Zhong. On product-one sequences with congruence conditions over non-abelian groups. Journal of Number Theory, 238(??):253–268, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003036.

Grishkov:2022:FCM

[674] A. Grishkov, D. Logachev, and A. Zobnin. L-functions of Carlitz modules, resultantal varieties and rooted binary trees — I. Journal of Number Theory, 238(??):269–312, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003073.

Aiylam:2022:SBS

[675] Dhroova Aiylam and Tanya Khovanova. Stern–Brocot sequences from weighted mediants. Journal of Number Theory, 238(??):313–330, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003061.

Momihara:2022:PGS

[676] Koji Momihara. Powers of Gauss sums in quadratic fields. Journal of Number Theory, 238(??):331–352, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21002961.

Elliott:2022:AEP

[677] Jesse Elliott. Asymptotic expansions of the prime counting function. Journal of Number Theory, 238(??):353–395, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003085.

Jeon:2022:CAC

[678] Daeyeol Jeon, Byoung Du Kim, and Chang Heon Kim. Construction of anti-cyclotomic Euler systems of modular abelian varieties, and the ranks of their Mordell–Weil groups. Journal of Number Theory, 238 (??):396–443, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003097.

Geisser:2022:KBM

[679] Thomas H. Geisser and Baptiste Morin. On the kernel of the Brauer–Manin pairing. Journal of Number Theory, 238(??):444–463, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003206.

Hoque:2022:CI

[680] Azizul Hoque. On a conjecture of Iizuka. Journal of Number Theory, 238(??):464–473, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21002870.

Lee:2022:FNC

[681] Junguk Lee, Wan Lee, Hayan Nam, and Myungjun Yu. Families of non-congruent numbers with odd prime factors of the form 8k + 3. Journal of Number Theory, 238(??):474–501, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2100322X.

Becher:2022:SRP

[682] Karim Johannes Becher and Parul Gupta. Square-reflexive polynomials. Journal of Number Theory, 238(??):502–534, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003231.

Lee:2022:MDS

[683] Eun Hye Lee and Ramin Takloo-Bighash. On a multiple Dirichlet series associated to binary cubic forms. Journal of Number Theory, 238 (??):535–556, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2100319X.

Yukie:2022:DUT

[684] Akihiko Yukie. On the density of unnormalized Tamagawa numbers of orthogonal groups III. Journal of Number Theory, 238(??):557–610, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003255.

Pollack:2022:MFI

[685] Aaron Pollack and Gordan Savin. Modular forms on indefinite orthogonal groups of rank three. Journal of Number Theory, 238(??):611–675, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003280.

Gejima:2022:SRC

[686] Kohta Gejima. Some remarks on the critical values of Rankin–Selberg L-series. Journal of Number Theory, 238(??):676–709, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003292.

Hirose:2022:OTR

[687] Minoru Hirose, Hideki Murahara, and Masataka Ono. Ohno-type relation for interpolated multiple zeta values. Journal of Number Theory, 238(??):710–730, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003309.

Stucky:2022:FSS

[688] Joshua Stucky. The fractional sum of small arithmetic functions. Journal of Number Theory, 238(??):731–739, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003334.

Brent:2022:MSE

[689] Richard P. Brent, David J. Platt, and Timothy S. Trudgian. The mean square of the error term in the prime number theorem. Journal of Number Theory, 238(??):740–762, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003322.

Yu:2022:CEL

[690] Wang-Xing Yu and Yong-Gao Chen. On a conjecture of Erdős and Lewin. Journal of Number Theory, 238(??):763–778, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003358.

Knyszewski:2022:IPR

[691] Franciszek Knyszewski. Integral parts of reciprocals of the Hurwitz zeta function and the polynomial families that arise in the study thereof. Journal of Number Theory, 238(??):779–807, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003450.

Fang:2022:BBI

[692] Jin-Hui Fang. On bounded basis of integers. Journal of Number Theory, 238(??):808–822, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003462.

Genao:2022:TBT

[693] Tyler Genao. Typically bounding torsion on elliptic curves with rational j-invariant. Journal of Number Theory, 238(??):823–841, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003619.

Ahmad:2022:ABQ

[694] Faseeh Ahmad, Robert L. Benedetto, Jennifer Cain, Gregory Carroll, and Lily Fang. The arithmetic basilica: a quadratic PCF arboreal Galois group. Journal of Number Theory, 238(??):842–868, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003620.

Cao:2022:FES

[695] Wei Cao. L-functions of exponential sums and Hasse polynomials. Journal of Number Theory, 238(??):869–882, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003632.

Keilthy:2022:MMZ

[696] Adam Keilthy. Motivic multiple zeta values and the block filtration. Journal of Number Theory, 238(??):883–919, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003656.

Garcia:2022:EEA

[697] Stephan Ramon Garcia and Ethan Simpson Lee. Explicit estimates for Artin L-functions: Duke’s short-sum theorem and Dedekind zeta residues. Journal of Number Theory, 238(??):920–943, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003644.

Wang:2022:BPH

[698] Haowu Wang and Brandon Williams. Borcherds products of halfintegral weight. Journal of Number Theory, 238(??):944–950, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003668.

Bell:2022:DPP

[699] Zoë Bell, Jasmine Camero, Karina Cho, Trevor Hyde, Chieh-Mi Lu, Bianca Thompson, and Eric Zhu. Density of periodic points for Lattès maps over finite field. Journal of Number Theory, 238(??):951–966, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100367X.

Wong:2022:SOC

[700] Peng-Jie Wong. Selberg’s orthogonality conjecture and symmetric power L-functions. Journal of Number Theory, 238(??):967–977, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003693.

Bharadwaj:2022:SVD

[701] Abhishek Bharadwaj and Siddhi Pathak. On special values of Dirichlet series with periodic coefficients. Journal of Number Theory, 238 (??):978–1011, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003711.

Benjamin:2022:DAI

[702] Elliot Benjamin and C. Snyder. Determination of all imaginary quadratic fields for which their Hilbert 2-class fields have 2-class groups of rank 2. Journal of Number Theory, 238(??):1012–1043, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2100370X.

Hamieh:2022:DV

[703] Alia Hamieh and Rory McClenagan. The distribution of values of \frac{L′}{L} (1/2 + ϵ, χ_{D}). Journal of Number Theory, 238(??):1044–1062, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003723.

Mihailescu:2022:DEL

[704] Preda Mihăilescu and Michael T. Rassias. Double exponential lower bounds for possible solutions in the Second Case of the Fermat Last Theorem. Journal of Number Theory, 238(??):1063–1080, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003735.

Billingsley:2022:ACI

[705] Tyler Raven Billingsley. An algorithm for checking injectivity of specialization maps from elliptic surfaces. Journal of Number Theory, 238 (??):1081–1106, September 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003772.

Anonymous:2022:PO

[706] Anonymous. Pages 1–500 (October 2022). Journal of Number Theory, 239(??):1–500, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBj

[707] Anonymous. Editorial Board. Journal of Number Theory, 239(??): ifc, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001159.

Khan:2022:WVE

[708] Rizwanur Khan, Micah B. Milinovich, and Unique Subedi. A weighted version of the Erdős–Kac theorem. Journal of Number Theory, 239(??): 1–20, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003681.

Melistas:2022:PAR

[709] Mentzelos Melistas. Purely additive reduction of abelian varieties with torsion. Journal of Number Theory, 239(??):21–39, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003759.

Fujita:2022:NNS

[710] Yasutsugu Fujita and Maohua Le. A note on the number of solutions of the Pillai type equation ∥a^{x} − b^{y}∥ = k. Journal of Number Theory, 239(??):40–56, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003760.

Weingartner:2022:NPF

[711] Andreas Weingartner. The number of prime factors of integers with dense divisors. Journal of Number Theory, 239(??):57–77, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21003784.

Drewitt:2022:LEE

[712] Joshua Drewitt. Laplace-eigenvalue equations for length three modular iterated integrals. Journal of Number Theory, 239(??):78–112, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003838.

Keating:2022:GSG

[713] Kevin Keating and Paul Schwartz. Galois scaffolds and Galois module structure for totally ramified C_{p^{2}}-extensions in characteristic 0. Journal of Number Theory, 239(??):113–136, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2100384X.

Richard:2022:MMP

[714] Rodolphe Richard. Manin–Mumford par le critère de Weyl. Journal of Number Theory, 239(??):137–150, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X21003863.

Matsumoto:2022:BMZ

[715] Kohji Matsumoto, Toshiki Matsusaka, and Ilija Tanackov. On the behavior of multiple zeta-functions with identical arguments on the real line. Journal of Number Theory, 239(??):151–182, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X21003802.

Jeon:2022:VWH

[716] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. On values of weakly holomorphic modular functions at divisors of meromorphic modular forms. Journal of Number Theory, 239(??):183–206, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003851. See [960].

Nathanson:2022:SSL

[717] Melvyn B. Nathanson. Sidon sets for linear forms. Journal of Number Theory, 239(??):207–227, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X21003826.

Kataoka:2022:FSI

[718] Takenori Kataoka. Finite submodules of Iwasawa modules for multivariable extensions. Journal of Number Theory, 239(??):228–250, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003917.

Martirosyan:2022:LAV

[719] Lilit Martirosyan and Pieter Moree. Large absolute values of cyclotomic polynomials at roots of unity. Journal of Number Theory, 239(??):251–272, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003899.

Besser:2022:AHV

[720] Amnon Besser. p-adic heights and Vologodsky integration. Journal of Number Theory, 239(??):273–297, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000038.

Minamide:2022:AGT

[721] Arata Minamide and Shota Tsujimura. Anabelian group-theoretic properties of the absolute Galois groups of discrete valuation fields. Journal of Number Theory, 239(??):298–334, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000014.

Jarviniemi:2022:SIE

[722] Olli Järviniemi. Simultaneous insolvability of exponential congruences. Journal of Number Theory, 239(??):335–351, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000063.

Pollack:2022:DSB

[723] Paul Pollack and Akash Singha Roy. Dirichlet, Sierpiński, and Benford. Journal of Number Theory, 239(??):352–364, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000099.

Ogrodnik:2022:LGC

[724] Brooke Logan Ogrodnik. On the local-global conjecture for commutator traces. Journal of Number Theory, 239(??):365–401, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003796.

tChu:2022:GDM

[725] Hùng Viêt. Chu, Dylan King, Noah Luntzlara, Thomas C. Martinez, Steven J. Miller, Lily Shao, Chenyang Sun, and Victor Xu. Generalizing the distribution of missing sums in sumsets. Journal of Number Theory, 239(??):402–444, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003875.

Flammang:2022:TDA

[726] V. Flammang. The total distance for algebraic integers having all their conjugates in a sector. Journal of Number Theory, 239(??):445– 461, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003887.

Schein:2022:SLT

[727] Michael M. Schein and Amir Shoan. Systolic length of triangular modular curves. Journal of Number Theory, 239(??):462–488, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X21003814.

ElFadil:2022:NMC

[728] Lhoussain El Fadil. On non monogenity of certain number fields defined by trinomials x^{6} + ax^{3} + b. Journal of Number Theory, 239 (??):489–500, October 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000026.

Anonymous:2022:PN

[729] Anonymous. Pages 1–754 (November 2022). Journal of Number Theory, 240(??):1–754, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBk

[730] Anonymous. Editorial Board. Journal of Number Theory, 240(??):ifc, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001470.

Freeman:2022:OTF

[731] Jesse Freeman. Optimal test functions for detecting central point vanishing over families of L-functions. Journal of Number Theory, 240(??):1–30, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X21003905.

Banerjee:2022:CFO

[732] Pradipto Banerjee and Ranjan Bera. Classifying families of orthogonal polynomials having Galois group the alternating group. Journal of Number Theory, 240(??):31–49, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000087.

Panti:2022:ADC

[733] Giovanni Panti. Attractors of dual continued fractions. Journal of Number Theory, 240(??):50–73, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000129.

Han:2022:RMV

[734] Jiyoung Han. Rogers’ mean value theorem for S-arithmetic Siegel transforms and applications to the geometry of numbers. Journal of Number Theory, 240(??):74–106, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000117.

Mao:2022:CMP

[735] Renrong Mao. Congruences for s̄pt_{ω}(n) modulo powers of primes. Journal of Number Theory, 240(??):107–123, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000208.

Gunji:2022:FCS

[736] Keiichi Gunji. On the Fourier coefficients of the Siegel Eisenstein series of odd level and the genus theta series. Journal of Number Theory, 240(??):124–144, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200021X.

Maurischat:2022:AIC

[737] Andreas Maurischat. Algebraic independence of the Carlitz period and its hyperderivatives. Journal of Number Theory, 240(??):145–162, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000233.

Philip:2022:VAC

[738] Séverin Philip. Variétés abéliennes CM et grosse monodromie finie sauvage. (French) [Abelian CM manifolds and big finite wild monodromy]. Journal of Number Theory, 240(??):163–195, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22000245.

Gekeler:2022:GPA

[739] Ernst-Ulrich Gekeler. Goss polynomials, q-adic expansions, and sheats compositions. Journal of Number Theory, 240(??):196–253, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22000269. See corrigendum [889].

Jakhar:2022:EIB

[740] Anuj Jakhar. Explicit integral basis of Q(ap_{1}p_{2}). Journal of Number Theory, 240(??):254–271, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000282.

Lalin:2022:HHT

[741] Matilde Lalín and Olivier Mila. Hyperbolic Heron triangles and elliptic curves. Journal of Number Theory, 240(??):272–295, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22000270.

Acharya:2022:ASF

[742] Ratnadeep Acharya, Prahlad Sharma, and Saurabh Kumar Singh. taspect subconvexity for GL(2) × GL(2) L-function. Journal of Number Theory, 240(??):296–324, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000294.

Mosunov:2022:AGB

[743] Anton Mosunov. On the automorphism group of a binary form associated with algebraic trigonometric quantities. Journal of Number Theory, 240(??):325–356, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000373.

Holm:2022:DAB

[744] Kristian Holm. On the distribution of angles between increasingly many short lattice vectors. Journal of Number Theory, 240(??):357– 403, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000452.

Shahidi:2022:RRM

[745] Freydoon Shahidi and William Sokurski. On the resolution of reductive monoids and multiplicativity of γ-factors. Journal of Number Theory, 240(??):404–438, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000543.

Li:2022:ARI

[746] Zhonghua Li. Algebraic relations of interpolated multiple zeta values. Journal of Number Theory, 240(??):439–470, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000488.

Tanaka:2022:RTM

[747] Tatsushi Tanaka and Noriko Wakabayashi. Rooted tree maps for multiple L-values. Journal of Number Theory, 240(??):471–489, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2200052X.

Mulay:2022:AIV

[748] Shashikant Mulay. Algebraic-integer valued polynomials. Journal of Number Theory, 240(??):490–521, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000506.

Yamamoto:2022:TSR

[749] Yuki Yamamoto. On the types for supercuspidal representations of inner forms of GL_{N}. Journal of Number Theory, 240(??):522–550, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000464.

Zhuang:2022:SCS

[750] Jincheng Zhuang, Qi Cheng, and Jiejing Wen. Solution counts and sums of roots of unity. Journal of Number Theory, 240(??):551–561, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000518.

Fejzullahu:2022:ELF

[751] Endrit Fejzullahu and Kevin Keating. Extensions of local fields given by 3-term Eisenstein polynomials. Journal of Number Theory, 240 (??):562–592, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000476.

Adedji:2022:SDE

[752] Kouèssi Norbert Adédji, Florian Luca, and Alain Togbé. On the solutions of the Diophantine equation F_{n} ± a(10^{m} − 1)9 = k! . Journal of Number Theory, 240(??):593–610, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000075.

Fiori:2022:UDC

[753] Andrew Fiori and Cameron Franc. The unbounded denominators conjecture for the noncongruence subgroups of index 7. Journal of Number Theory, 240(??):611–640, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000130.

Mossinghoff:2022:ELB

[754] Michael J. Mossinghoff, Valeriia V. Starichkova, and Timothy S. Trudgian. Explicit lower bounds on |L(1, χ)|. Journal of Number Theory, 240(??):641–655, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000105.

Tengely:2022:DPR

[755] Szabolcs Tengely and Maciej Ulas. Diophantine problems related to cyclic cubic and quartic fields. Journal of Number Theory, 240(??):656–684, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000221.

Morton:2022:APC

[756] Patrick Morton and Serban Raianu. Arithmetic properties of 3-cycles of quadratic maps over Q. Journal of Number Theory, 240(??):685–729, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000257.

Keller:2022:SSG

[757] Christoph A. Keller and Jason M. Quinones. On the space of slow growing weak Jacobi forms. Journal of Number Theory, 240(??):730–750, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2200049X.

McMeekin:2022:CAP

[758] Christine McMeekin. Corrigendum to: “On the asymptotics of a prime spin relation” [J. Number Theor. 200 (2019) 407–426]. Journal of Number Theory, 240(??):751–753, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000531. See

Anonymous:2022:PD

[759] Anonymous. Pages 1–760 (December 2022). Journal of Number Theory, 241(??):1–760, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2022:EBl

[760] Anonymous. Editorial Board. Journal of Number Theory, 241(??):ifc, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001718.

Zou:2022:SRD

[761] Jiandi Zou. Supercuspidal representations of GL_{n} (F) distinguished by an orthogonal involution. Journal of Number Theory, 241(??):1–56, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000889.

Baier:2022:DAP

[762] Stephan Baier, Esrafil Ali Molla, and Arijit Ganguly. Diophantine approximation with prime restriction in function fields. Journal of Number Theory, 241(??):57–90, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000579.

Espinoza:2022:BHZ

[763] Milton Espinoza. The Barnes–Hurwitz zeta cocycle on PGL_{2}Q. Journal of Number Theory, 241(??):91–119, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000555.

Dawsey:2022:SPS

[764] Madeline Locus Dawsey, Matthew Just, and Robert Schneider. A “supernormal” partition statistic. Journal of Number Theory, 241 (??):120–141, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000804.

Hughes:2022:DMV

[765] Christopher Hughes and Andrew Pearce-Crump. A discrete mean-value theorem for the higher derivatives of the Riemann zeta function. Journal of Number Theory, 241(??):142–164, December 2022. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000798.

Comeau-Lapointe:2022:OLD

[766] Antoine Comeau-Lapointe. One-level density of the family of twists of an elliptic curve over function fields. Journal of Number Theory, 241(??):165–197, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000890.

Lin:2022:FCE

[767] Xin Lin and Chao Chen. L-functions of certain exponential sums over finite fields II. Journal of Number Theory, 241(??):198–220, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000816.

Caginalp:2022:PRA

[768] Gunduz Caginalp and Bogdan Ion. Probabilistic renormalization and analytic continuation. Journal of Number Theory, 241(??):221–246, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000828.

Kim:2022:RNR

[769] Mingyu Kim and Byeong-Kweon Oh. The rank of new regular quadratic forms. Journal of Number Theory, 241(??):247–261, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000853.

Wu:2022:PFC

[770] Chenyan Wu. Periods and (χ, b)-factors of cuspidal automorphic forms of metaplectic groups. Journal of Number Theory, 241(??):262–296, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000841.

Walji:2022:DSP

[771] Nahid Walji. On the decomposition of symmetric powers of automorphic representations for GL(3) and GL(4). Journal of Number Theory, 241(??):297–313, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000907.

Liu:2022:GVL

[772] Kui Liu, Meijie Lu, and Xianchang Meng. Generalized visibility of lattice points in higher dimensions. Journal of Number Theory, 241 (??):314–329, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200083X.

Kim:2022:MVF

[773] Seungki Kim. Mean value formulas on sublattices and flags of the random lattice. Journal of Number Theory, 241(??):330–351, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000920.

Sun:2022:SFS

[774] Cui-Fang Sun and Zhi Cheng. On the structure of finite sets with the same representation functions. Journal of Number Theory, 241 (??):352–363, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000944.

Gonzalez:2022:EES

[775] Oscar E. González. Effective estimates for the smallest parts function. Journal of Number Theory, 241(??):364–386, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000919.

Jindal:2022:FLP

[776] Ankita Jindal and Shanta Laishram. Families of Laguerre polynomials with alternating group as Galois group. Journal of Number Theory, 241(??):387–429, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001068.

Kaneko:2022:DER

[777] Hajime Kaneko and Makoto Kawashima. The digit exchanges in the rotational beta expansions of algebraic numbers. Journal of Number Theory, 241(??):430–449, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22000993.

Jin:2022:NNO

[778] Seokho Jin and Wenjun Ma. A note on non-ordinary primes for some genus-zero arithmetic groups. Journal of Number Theory, 241 (??):450–464, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001020.

Wang:2022:MQM

[779] Weijia Wang and Hao Zhang. Meromorphic quasi-modular forms and their L-functions. Journal of Number Theory, 241(??):465–503, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001081.

Martinez-Ranero:2022:EDA

[780] Carlos Martínez-Ranero, Javier Utreras, and Xavier Vidaux. Existential decidability for addition and divisibility in holomorphy subrings of global fields. Journal of Number Theory, 241(??):504–530, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22000981.

Kuperberg:2022:PPD

[781] Vivian Kuperberg. On pseudo-polynomials divisible only by a sparse set of primes and α-primary pseudo-polynomials. Journal of Number Theory, 241(??):531–541, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000968.

Hsu:2022:PCH

[782] Chi-Yun Hsu. Partial classicality of Hilbert modular forms. Journal of Number Theory, 241(??):542–562, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2200107X.

Quang:2022:EFF

[783] Si Duc Quang. An effective function field version of Schmidt’s subspace theorem for projective varieties, with arbitrary families of homogeneous polynomials. Journal of Number Theory, 241(??):563–580, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2200097X.

Stromberg:2022:RAH

[784] Fredrik Strömberg. A reduction algorithm for Hilbert modular groups. Journal of Number Theory, 241(??):581–602, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000567.

Chen:2022:CRT

[785] Imin Chen, Gleb Glebov, and Ritesh Goenka. Chudnovsky–Ramanujan type formulae for non-compact arithmetic triangle groups. Journal of Number Theory, 241(??):603–654, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000580.

Rojas:2022:RRF

[786] J. Maurice Rojas and Yuyu Zhu. Root repulsion and faster solving for very sparse polynomials over p-adic fields. Journal of Number Theory, 241(??):655–699, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000877.

Kadiri:2022:PCD

[787] Habiba Kadiri and Peng-Jie Wong. Primes in the Chebotarev density theorem for all number fields (with an appendix by Andrew Fiori). Journal of Number Theory, 241(??):700–737, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22000865.

Gao:2022:RZS

[788] Weidong Gao, Siao Hong, Wanzhen Hui, Xue Li, Qiuyu Yin, and Pingping Zhao. Representation of zero-sum invariants by sets of zero-sum sequences over a finite abelian group II. Journal of Number Theory, 241(??):738–760, December 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22000932.

Anonymous:2023:PJa

[789] Anonymous. Pages 1–754 (January 2023). Journal of Number Theory, 242(??):1–754, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBa

[790] Anonymous. Editorial Board. Journal of Number Theory, 242(??): ifc, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001962.

Guo:2023:LPS

[791] Jingwei Guo and Tao Jiang. Lattice points in stretched finite type domains. Journal of Number Theory, 242(??):1–20, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001603.

Beelen:2023:CAG

[792] Peter Beelen, Leonardo Landi, and Maria Montanucci. Classification of all Galois subcovers of the skabelund maximal curves. Journal of Number Theory, 242(??):46–72, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22001652.

Wei:2023:SFH

[793] Xunjing Wei. Smallness of Faltings heights of CM abelian varieties. Journal of Number Theory, 242(??):73–101, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001676.

Bag:2023:BBS

[794] Nilanjan Bag and Igor E. Shparlinski. Bounds on bilinear sums of Kloosterman sums. Journal of Number Theory, 242(??):102–111, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001597.

Andersen:2023:HSP

[795] Nickolas Andersen and Han Wu. Hybrid subconvexity and the partition function. Journal of Number Theory, 242(??):154–180, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001688.

Nathanson:2023:UEF

[796] Melvyn B. Nathanson. Underapproximation by Egyptian fractions. Journal of Number Theory, 242(??):208–234, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001615.

Wu:2023:GOC

[797] Han Wu. On genus one curves violating the local-global principle. Journal of Number Theory, 242(??):235–243, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001858.

Kim:2023:STC

[798] Seoyoung Kim. The Sato–Tate conjecture and Nagao’s conjecture. Journal of Number Theory, 242(??):245–277, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001093.

Yoo:2023:RCD

[799] Hwajong Yoo. The rational cuspidal divisor class group of X_{0}(N). Journal of Number Theory, 242(??):278–401, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001056.

Koperecz:2023:TRF

[800] Julien Koperecz. Triquadratic p-rational fields. Journal of Number Theory, 242(??):402–408, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001007.

Schmidt:2023:AES

[801] Matthew Schmidt. T -adic exponential sums over affinoids. Journal of Number Theory, 242(??):409–435, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000956.

Valloni:2023:FDK

[802] Domenico Valloni. Fields of definition of K3 surfaces with complex multiplication. Journal of Number Theory, 242(??):436–470, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001019.

Shparlinski:2023:GPM

[803] Igor E. Shparlinski. On the Glasner property for matrices with polynomial entries. Journal of Number Theory, 242(??):522–531, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001044.

Assim:2023:GGC

[804] J. Assim and Z. Boughadi. On Greenberg’s generalized conjecture. Journal of Number Theory, 242(??):576–602, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001287.

Tamiozzo:2023:THS

[805] Matteo Tamiozzo. The tame Hilbert symbol via K-theory and central extensions. Journal of Number Theory, 242(??):603–625, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001263.

Hajdu:2023:EPR

[806] L. Hajdu and O. Herendi. Extrema of polynomials with real roots and Diophantine equations. Journal of Number Theory, 242(??):626– 646, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001275.

Pankowski:2023:ZCN

[807] Łukasz Pańkowski and Mattia Righetti. Zeros of L(s) + L(2s) + ⋯ + L(Ns) in the region of absolute convergence. Journal of Number Theory, 242(??):647–659, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001317.

Heap:2023:PPA

[808] Winston Heap, Anurag Sahay, and Trevor D. Wooley. A paucity problem associated with a shifted integer analogue of the divisor function. Journal of Number Theory, 242(??):660–668, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001299.

Minac:2023:APE

[809] Ján Mináč, Andrew Schultz, and John Swallow. Arithmetic properties encoded in the Galois module structure of K^{×}K^{×p^{m}}. Journal of Number Theory, 242(??):669–708, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22001305.

Hu:2023:JCM

[810] Guangwei Hu and Guangshi Lü. Jutila’s circle method and GL2 × GL 2 shiftedconvolutionsums.JournalofNumberTheory, 242(??) : 709 − −724, January2023.CODENJNUT A9.ISSN0022−314X(print), 1096−

1658(electronic).URLhttpwwwsciencedirectcomsciencearticlepiiS0022314X22001330.

Palacios:2023:FEA

[811] Luis Santiago Palacios. Functional equation of the p-adic L-function of Bianchi modular forms. Journal of Number Theory, 242(??):725– 753, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001329.

Barrios:2023:GEC

[812] Alexander J. Barrios. Good elliptic curves with a specified torsion subgroup. Journal of Number Theory, 242(??):21–43, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001640.

Smoot:2023:CFE

[813] Nicolas Allen Smoot. A congruence family for 2-elongated plane partitions: an application of the localization method. Journal of Number Theory, 242(??):112–153, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200169X.

Ju:2023:PUQ

[814] Jangwon Ju, Daejun Kim, Kyoungmin Kim, Mingyu Kim, and Byeong-Kweon Oh. Primitively universal quaternary quadratic forms. Journal of Number Theory, 242(??):181–207, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001627.

Wang:2023:MRC

[815] Liuquan Wang and Yifan Yang. Moments of ranks and cranks, and quotients of Eisenstein series and the Dedekind eta function. Journal of Number Theory, 242(??):471–521, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001032.

Minac:2023:FPQ

[816] Ján Mináč, Tung T. Nguyen, and Nguy˜ ên Duy Tân. Fekete polynomials, quadratic residues, and arithmetic. Journal of Number Theory, 242(??):532–575, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001251.

Hirakawa:2023:CUP

[817] Yoshinosuke Hirakawa and Hideki Matsumura. Corrigendum to “A unique pair of triangles” [J. Number Theory 194 (2019) 297–302]. Journal of Number Theory, 242(??):44–45, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001585.

Aygin:2023:CMP

[818] Zafer Selcuk Aygin and Khoa D. Nguyen. Corrigendum to “Monogenic pure cubics” [J. Number Theory 219 (2021) 356–367]. Journal of Number Theory, 242(??):244, January 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001639.

Anonymous:2023:PF

[819] Anonymous. Pages 1–728 (February 2023). Journal of Number Theory, 243(??):1–728, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBb

[820] Anonymous. Editorial Board. Journal of Number Theory, 243(??):ifc, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002219.

Tang:2023:NGR

[821] Shiang Tang. A note on Galois representations valued in reductive groups with open image. Journal of Number Theory, 243(??):1–12, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001871.

Gun:2023:CIR

[822] Sanoli Gun, Olivier Ramaré, and Jyothsnaa Sivaraman. Counting ideals in ray classes. Journal of Number Theory, 243(??):13–37, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001901.

Frahm:2023:TRZ

[823] Jan Frahm and Polyxeni Spilioti. Twisted Ruelle zeta function at zero for compact hyperbolic surfaces. Journal of Number Theory, 243(??):38–61, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001895.

Luo:2023:BRN

[824] Zhilin Luo, Qinghua Pi, and Han Wu. Bias of root numbers for Hilbert newforms of cubic level. Journal of Number Theory, 243(??):62–116, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001937.

Bernstein:2023:PGI

[825] Joseph Bernstein and Andre Reznikov. Periods and global invariants of automorphic representations. Journal of Number Theory, 243 (??):117–159, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001664.

Stucky:2023:SMA

[826] Joshua Stucky. The sixth moment of automorphic L-functions. Journal of Number Theory, 243(??):160–201, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001925.

Hauke:2023:PCH

[827] Manuel Hauke and Agamemnon Zafeiropoulos. Poissonian correlations of higher orders. Journal of Number Theory, 243(??):202–240, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001913.

Arala:2023:DSP

[828] Nuno Arala. On the density of a set of primes associated to an elliptic curve. Journal of Number Theory, 243(??):241–255, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22001883.

Noguchi:2023:MST

[829] Hiroshi Noguchi. On multiplier systems and theta functions of halfintegral weight for the Hilbert modular group SL_{2}(o). Journal of Number Theory, 243(??):298–327, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22002049.

Adolphson:2023:HSR

[830] Alan Adolphson and Steven Sperber. Hyperkloosterman sums revisited. Journal of Number Theory, 243(??):328–351, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001457.

Su:2023:JFH

[831] Ren-He Su. On Jacobi forms of half-integral weight and matrix level. Journal of Number Theory, 243(??):352–384, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001445.

Battistoni:2023:AEN

[832] Francesco Battistoni and Hassan Oukhaba. Arithmetic equivalence for non-geometric extensions of global function fields. Journal of Number Theory, 243(??):385–411, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22001421.

Morin:2023:SVF

[833] Adrien Morin. Special values of L-functions on regular arithmetic schemes of dimension 1. Journal of Number Theory, 243(??):412–474, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001433.

Leslie:2023:FLH

[834] Spencer Leslie. A fundamental lemma for the Hecke algebra: the Jacquet–Rallis case. Journal of Number Theory, 243(??):475–494, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22001366.

Huang:2023:HPM

[835] Jing Huang, Wenguang Zhai, and Deyu Zhang. Higher power moments of symmetric square L-function. Journal of Number Theory, 243(??):495–517, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001391.

Weissauer:2023:EPL

[836] Rainer Weissauer. Exceptional poles of local L-functions for GSp(4) with respect to split Bessel models. Journal of Number Theory, 243 (??):518–560, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001378.

Lev:2023:SDC

[837] Vsevolod F. Lev. Small doubling in cyclic groups. Journal of Number Theory, 243(??):561–614, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22001408.

Bergelson:2023:FST

[838] Vitaly Bergelson and Andrew Best. The Furstenberg–Sárközy theorem and asymptotic total ergodicity phenomena in modular rings. Journal of Number Theory, 243(??):615–645, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22001342.

Kawashima:2023:PAC

[839] Makoto Kawashima and Anthony Poëls. Padé approximation for a class of hypergeometric functions and parametric geometry of numbers. Journal of Number Theory, 243(??):646–687, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22001354.

Fowler:2023:ETS

[840] Guy Fowler. Equations in three singular moduli: the equal exponent case. Journal of Number Theory, 243(??):256–297, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002116.

Martin:2023:CFN

[841] Daniel E. Martin. Continued fractions over non-Euclidean imaginary quadratic rings. Journal of Number Theory, 243(??):688–714, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2200141X.

Barsakci:2023:SRP

[842] Burcu Barsakçı and Mohammad Sadek. Simultaneous rational periodic points of degree-2 rational maps. Journal of Number Theory, 243(??):715–728, February 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200138X.

Anonymous:2023:PMa

[843] Anonymous. Pages 1–462 (March 2023). Journal of Number Theory, 244 (??):1–462, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBc

[844] Anonymous. Editorial Board. Journal of Number Theory, 244(??): ifc, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002360.

Ghioca:2023:ZDO

[845] Dragos Ghioca and Sina Saleh. Zariski dense orbits for endomorphisms of a power of the additive group scheme defined over finite fields. Journal of Number Theory, 244(??):1–23, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22002074.

Sakai:2023:ACW

[846] Kazuhiro Sakai. Algebraic construction of Weyl invariant E_{8} Jacobi forms. Journal of Number Theory, 244(??):42–62, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002153.

Gun:2023:LID

[847] Sanoli Gun, Neelam Kandhil, and Patrice Philippon. On linear independence of Dirichlet L values. Journal of Number Theory, 244(??): 63–83, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002165.

Simonic:2023:AFM

[848] Aleksander Simonič and Valeriia V. Starichkova. Atkinson’s formula for the mean square of ζ(s) with an explicit error term. Journal of Number Theory, 244(??):111–168, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22002141.

Schleischitz:2023:OTI

[849] Johannes Schleischitz. Optimality of two inequalities for exponents of Diophantine approximation. Journal of Number Theory, 244(??):169– 203, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002037.

Maletto:2023:GRB

[850] Simone Maletto. Galois representations with big image in the general symplectic group GSp_{4}(Z_{p}). Journal of Number Theory, 244(??):204– 223, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002189.

Ernvall-Hytonen:2023:EFS

[851] Anne-Maria Ernvall-Hytönen, Tapani Matala-aho, and Louna Seppälä. Euler’s factorial series, Hardy integral, and continued fractions. Journal of Number Theory, 244(??):224–250, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22002086.

Dokchitser:2023:NHC

[852] Vladimir Dokchitser and Adam Morgan. A note on hyperelliptic curves with ordinary reduction over 2-adic fields. Journal of Number Theory, 244(??):264–278, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22001949.

Murakami:2023:WFG

[853] Kazuaki Murakami. A weak form of Greenberg’s generalized conjecture for imaginary quadratic fields. Journal of Number Theory, 244(??):308–338, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002025.

Chiloyan:2023:IFI

[854] Garen Chiloyan. Infinite families of isogeny-torsion graphs. Journal of Number Theory, 244(??):369–417, March 2023. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22002128.

Meiners:2023:SZF

[855] Alexander Meiners and Boris Vertman. Spectral zeta function on discrete tori and Epstein–Riemann Hypothesis. Journal of Number Theory, 244(??):418–461, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2200186X.

Raska:2023:RMR

[856] Martin Raška. Representing multiples of m in real quadratic fields as sums of squares. Journal of Number Theory, 244(??):24–41, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2200213X.

Hoshi:2023:NOT

[857] Akinari Hoshi, Kazuki Kanai, and Aiichi Yamasaki. Norm one tori and Hasse norm principle, II: Degree 12 case. Journal of Number Theory, 244(??):84–110, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002062.

Kristensen:2023:BTF

[858] Simon Kristensen and Oskar Mathiasen. BBP-type formulas — an elementary approach. Journal of Number Theory, 244(??):251–263, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002013.

Li:2023:BVC

[859] Jiahui (Stella) Li and Steven J. Miller. Bounding vanishing at the central point of cuspidal newforms. Journal of Number Theory, 244(??): 279–307, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002190.

Ding:2023:PSL

[860] Jintai Ding, Seungki Kim, Tsuyoshi Takagi, Yuntao Wang, and Bo yin Yang. A physical study of the LLL algorithm. Journal of Number Theory, 244(??):339–368, March 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002104.

Anonymous:2023:PAa

[861] Anonymous. Pages 1–324 (April 2023). Journal of Number Theory, 245 (??):1–324, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBd

[862] Anonymous. Editorial Board. Journal of Number Theory, 245(??): ifc, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000021.

Oki:2023:CCS

[863] Yasuhiro Oki. On the connected components of Shimura varieties for CM unitary groups in odd variables. Journal of Number Theory, 245(??):1–64, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002177.

Ren:2023:NPE

[864] Rufei Ren. Newton polygon of exponential sums in two variables with triangular base. Journal of Number Theory, 245(??):119–149, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2200227X.

Forst:2023:ZMP

[865] Maxwell Forst and Lenny Fukshansky. On zeros of multilinear polynomials. Journal of Number Theory, 245(??):169–186, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002347.

Cheng:2023:EMS

[866] Chuangxun Cheng and Jiuyi Lu. On the existence of maximal spanning vectors in L^{2}(Q_{2}) and L^{2}(F_{2}((T))). Journal of Number Theory, 245(??):187–202, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002311.

Aoki:2023:CBA

[867] Miho Aoki and Shin ya Koyama. Chebyshev’s bias against splitting and principal primes in global fields. Journal of Number Theory, 245(??):233–262, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002335.

Chung:2023:FCE

[868] Kwun Chung. Factorization of coefficients for exponential and logarithm in function fields. Journal of Number Theory, 245(??):263– 291, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002281.

Ding:2023:CRA

[869] Yuchen Ding. On a conjecture of Ramírez Alfonsín and Skałba. Journal of Number Theory, 245(??):292–302, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22002426.

Li:2023:PPD

[870] Ruofan Li. Primitive prime divisors in backward orbits. Journal of Number Theory, 245(??):303–324, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X22002451.

Gil-Munoz:2023:RIH

[871] Daniel Gil-Muñoz. The ring of integers of Hopf-Galois degree p extensions of p-adic fields with dihedral normal closure. Journal of Number Theory, 245(??):65–118, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002323.

Melistas:2023:DRB

[872] Mentzelos Melistas. A divisibility related to the Birch and Swinnerton-Dyer conjecture. Journal of Number Theory, 245(??):150–168, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2200230X.

Tinkova:2023:ATP

[873] Magdaléna Tinková, Ezra Waxman, and Mikuláš Zindulka. Artin twin primes. Journal of Number Theory, 245(??):203–232, April 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22002293.

Anonymous:2023:PM

[874] Anonymous. Pages 1–332 (May 2023). Journal of Number Theory, 246 (??):1–332, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBe

[875] Anonymous. Editorial Board. Journal of Number Theory, 246(??): ifc, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000185.

Kim:2023:DQQ

[876] Kyoungmin Kim. Diagonal quinary quadratic forms with a strong regularity property. Journal of Number Theory, 246(??):1–14, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22002505.

Wagner:2023:JFC

[877] Ian Wagner. Jacobi forms with CM and applications. Journal of Number Theory, 246(??):15–48, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002542.

Gillespie:2023:DSM

[878] Timothy Gillespie, Praneel Samanta, and Yangbo Ye. Double square moments and bounds for resonance sums of cusp forms. Journal of Number Theory, 246(??):166–188, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002487.

Sun:2023:UIB

[879] Chia-Liang Sun. Unlikely intersections between additive and multiplicative structures. Journal of Number Theory, 246(??):252–264, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002414.

Wong:2023:ALP

[880] Siman Wong. Abelian lifts of polynomials. Journal of Number Theory, 246(??):265–278, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002098.

Stokes:2023:GFT

[881] Matt Stokes. On Gauss factorials and their connection to the cyclotomic λ-invariants of imaginary quadratic fields. Journal of Number Theory, 246(??):279–293, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002530. See corrigendum [959].

Chu:2023:CRN

[882] Rena Chu and Julie Desjardins. Constant root number on integer fibres of elliptic surfaces. Journal of Number Theory, 246(??):294– 325, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002463.

Ash:2023:CCS

[883] Avner Ash and Dan Yasaki. Cohomology of congruence subgroups of SL_{3}(Z), Steinberg modules, and real quadratic fields. Journal of Number Theory, 246(??):49–86, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X22002438.

Deng:2023:TCC

[884] Taiwang Deng. Torsions in cohomology of SL_{2}(Z) and congruence of modular forms. Journal of Number Theory, 246(??):87–156, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22002529.

Wong:2023:RCQ

[885] Siman Wong. On the rank of cubic and quartic twists of elliptic curves by primes. Journal of Number Theory, 246(??):157–165, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002050.

DallAva:2023:ABT

[886] Luca Dall’Ava. Approximations of the balanced triple product padic L-function. Journal of Number Theory, 246(??):189–226, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X22002475.

Lang:2023:DHL

[887] Hannah Lang, Hamilton Wan, and Nancy Xu. Distributions of hook lengths divisible by two or three. Journal of Number Theory, 246(??): 227–251, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X22002517.

Daowsud:2023:CCF

[888] Katthaleeya Daowsud and Thomas A. Schmidt. Corrigendum to: “Continued fractions for rational torsion” [J. Number Theory 189 (2018) 115– 130]. Journal of Number Theory, 246(??):326–327, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002554.

Gekeler:2023:CGP

[889] Ernst-Ulrich Gekeler. Corrigendum to: “Goss polynomials, q-adic expansions, and Sheats compositions” [J. Number Theory 240 (2022) 196– 253]. Journal of Number Theory, 246(??):328–331, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002566.

See [739].

Anonymous:2023:PJb

[890] Anonymous. Pages 1–210 (June 2023). Journal of Number Theory, 247 (??):1–210, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBf

[891] Anonymous. Editorial Board. Journal of Number Theory, 247(??): ifc, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000434.

Gao:2023:MSQ

[892] Peng Gao and Liangyi Zhao. Mean squares of quadratic twists of the Möbius function. Journal of Number Theory, 247(??):1–14, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23000094.

Burcroff:2023:SAM

[893] Amanda Burcroff and Noah Luntzlara. Sets arising as minimal additive complements in the integers. Journal of Number Theory, 247(??): 15–34, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000082.

Li:2023:SIS

[894] Jiamin Li and Jing Ma. On a sum involving small arithmetic function and the integral part function. Journal of Number Theory, 247(??): 35–45, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000148.

Hu:2023:UBR

[895] Yu Hu, Yan Huang, and Derong Kong. Univoque bases of real numbers: Simply normal bases, irregular bases and multiple rationals. Journal of Number Theory, 247(??):46–74, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000136.

Andrews:2023:GCS

[896] George Andrews and William Keith. A general class of Schmidt theorems. Journal of Number Theory, 247(??):75–99, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000100.

Cully-Hugill:2023:PBC

[897] Michaela Cully-Hugill. Primes between consecutive powers. Journal of Number Theory, 247(??):100–117, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000112.

Matschke:2023:QFA

[898] Benjamin Matschke and Abhijit S. Mudigonda. Quadratic fields admitting elliptic curves with rational j-invariant and good reduction everywhere. Journal of Number Theory, 247(??):162–210, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2200244X.

Languasco:2023:USC

[899] Alessandro Languasco. A unified strategy to compute some special functions of number-theoretic interest. Journal of Number Theory, 247(??):118–161, June 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000124.

Anonymous:2023:PJc

[900] Anonymous. Pages 1–446 (July 2023). Journal of Number Theory, 248 (??):1–446, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBg

[901] Anonymous. Editorial Board. Journal of Number Theory, 248(??): ifc, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000653.

Honnor:2023:CTF

[902] Matthew H. Honnor. Comparing two formulas for the Gross–Stark units. Journal of Number Theory, 248(??):368–419, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000410.

Mondal:2023:DIA

[903] Amiya Kumar Mondal and Santosh Nadimpalli. On the duality involution for p-adic general spin groups. Journal of Number Theory, 248(??):1–13, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000306.

Chattopadhyay:2023:RCQ

[904] Jaitra Chattopadhyay, H. Laxmi, and Anupam Saikia. On the prationality of consecutive quadratic fields. Journal of Number Theory, 248(??):14–26, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000240.

Adersh:2023:NSK

[905] V. K. Adersh, M. Manickam, and M. M. Sreejith. On newforms and Saito–Kurokawa lifts. Journal of Number Theory, 248(??):27– 53, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300029X.

Boylan:2023:JES

[906] Hatice Boylan. Jacobi–Eisenstein series over number fields. Journal of Number Theory, 248(??):54–77, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2300032X.

Lei:2023:LIE

[907] Austin Lei, Tianyu Ni, and Hui Xue. Linear independence of even periods of modular forms. Journal of Number Theory, 248(??):120– 139, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000276.

Katsurada:2023:TKM

[908] Hidenori Katsurada and Henry H. Kim. Twisted Koecher–Maass series of the Ikeda type lift for the exceptional group of type E_{7,3}. Journal of Number Theory, 248(??):140–171, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000264.

Hou:2023:TCS

[909] Fei Hou. Triple correlation sums of coefficients of cuspidal forms. Journal of Number Theory, 248(??):172–207, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000288.

Hindes:2023:IPQ

[910] Wade Hindes, Reiyah Jacobs, and Peter Ye. Irreducible polynomials in quadratic semigroups. Journal of Number Theory, 248(??):208– 241, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000331.

Chanana:2023:SFC

[911] Himanshi Chanana and Saurabh Kumar Singh. Sum of the Fourier coefficients of SL(3, Z) Hecke Maass forms over quadratics. Journal of Number Theory, 248(??):242–260, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000355.

Khaochim:2023:EBT

[912] Narissara Khaochim and Riad Masri. Effective bounds for traces of weak Maass forms. Journal of Number Theory, 248(??):261– 293, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000252.

Miezaki:2023:PNH

[913] Tsuyoshi Miezaki and Gabriele Nebe. Pseudo-normalized Hecke eigenform and its application to extremal 2-modular lattices. Journal of Number Theory, 248(??):294–309, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000409.

Ray:2023:DDF

[914] Chiranjit Ray. Distribution and divisibility of the Fourier coefficients of certain Hauptmoduln. Journal of Number Theory, 248(??):343– 367, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000379.

Li:2023:BDF

[915] Yao-Qiang Li. Bifurcations of digit frequencies in unique expansions. Journal of Number Theory, 248(??):420–446, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000367.

Gras:2023:CHF

[916] Georges Gras. The Chevalley–Herbrand formula and the real abelian Main Conjecture (new criterion using capitulation of the class group). Journal of Number Theory, 248(??):78–119, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000343.

Jackson:2023:UBB

[917] Faye Jackson and Misheel Otgonbayar. Unexpected biases between congruence classes for parts in k-indivisible partitions. Journal of Number Theory, 248(??):310–342, July 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000318.

Anonymous:2023:PAb

[918] Anonymous. Pages 1–602 (August 2023). Journal of Number Theory, 249 (??):1–602, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBh

[919] Anonymous. Editorial Board. Journal of Number Theory, 249(??): ifc, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300080X.

Kwon:2023:BMS

[920] Jaesung Kwon. Bianchi modular symbols and p-adic L-functions. Journal of Number Theory, 249(??):274–313, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000549.

Choi:2023:NPH

[921] Dohoon Choi and Youngmin Lee. Number of proportions of Hecke eigenvalues in two newforms. Journal of Number Theory, 249(??):1– 20, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000483.

Peng:2023:UIP

[922] Wayne Peng. Unlikely intersection problems for restricted lifts of a p-th power. Journal of Number Theory, 249(??):21–48, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000392.

Lee:2023:CDA

[923] Jungin Lee. Counting 3-dimensional algebraic tori over Q. Journal of Number Theory, 249(??):49–92, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000525.

Hittmeir:2023:IFS

[924] Markus Hittmeir. Integer factorization as subset-sum problem. Journal of Number Theory, 249(??):93–118, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000574.

Casazza:2023:AFP

[925] Daniele Casazza and Carlos de Vera-Piquero. On p-adic L-functions for GL(2) × GL(3) via pullbacks of Saito–Kurokawa lifts. Journal of Number Theory, 249(??):131–182, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000550.

Ouaggag:2023:ERA

[926] Zouhair Ouaggag. Effective rational approximation on spheres. Journal of Number Theory, 249(??):183–208, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000562.

Blackman:2023:RAL

[927] John Blackman. Reformulating the p-adic Littlewood Conjecture in terms of infinite loops modp^{k}. Journal of Number Theory, 249(??):209– 236, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000598.

Bajpai:2023:GFC

[928] Jitendra Bajpai, Subham Bhakta, and Renan Finder. Growth of Fourier coefficients of vector-valued automorphic forms. Journal of Number Theory, 249(??):237–273, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000513.

Blakestad:2023:UAS

[929] Clifford Blakestad and David Grant. On the universal p-adic sigma and Weierstrass zeta functions. Journal of Number Theory, 249(??):348– 376, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000380.

Dicks:2023:CRC

[930] Robert Dicks. Congruence relations for r-colored partitions. Journal of Number Theory, 249(??):377–399, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22002499.

Jing:2023:ESP

[931] Mengyao Jing and Huaning Liu. Exponential sums with Piatetski– Shapiro sequences and their application to modular hyperbolas. Journal of Number Theory, 249(??):400–421, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000628.

Kim:2023:QFR

[932] Mingyu Kim and Byeong-Kweon Oh. Quadratic forms representing large integers only. Journal of Number Theory, 249(??):422–440, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000501.

Schwagenscheidt:2023:BTF

[933] Markus Schwagenscheidt and Brandon Williams. Binary theta functions and Borcherds products. Journal of Number Theory, 249(??):441– 461, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000495.

Avelar:2023:NBC

[934] D. V. Avelar, F. E. Brochero Martínez, and S. Ribas. A note on Bass’ conjecture. Journal of Number Theory, 249(??):462–469, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000616.

Charlier:2023:SAN

[935] Émilie Charlier, Célia Cisternino, Zuzana Masáková, and Edita Pelantová. Spectrum, algebraicity and normalization in alternate bases. Journal of Number Theory, 249(??):470–499, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2300063X.

Kamaraj:2023:USG

[936] Ramanujam Kamaraj, Ben Kane, and Ryoko Oishi-Tomiyasu. Universal sums of generalized heptagonal numbers. Journal of Number Theory, 249(??):500–536, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000537.

Vermeulen:2023:AEN

[937] Floris Vermeulen. Arithmetically equivalent number fields have approximately the same successive minima. Journal of Number Theory, 249(??):119–130, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000586.

Jia:2023:IBD

[938] Dennis X. Q. Jia. Inequalities for the broken k-diamond partition functions. Journal of Number Theory, 249(??):314–347, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23000604.

Zhou:2023:DCO

[939] Hai-Gang Zhou and Xiao-Jie Zhu. Double coset operators and eta-quotients. Journal of Number Theory, 249(??):537–601, August 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23000707.

Anonymous:2023:PS

[940] Anonymous. Pages 1–206 (September 2023). Journal of Number Theory, 250(??):1–206, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBi

[941] Anonymous. Editorial Board. Journal of Number Theory, 250(??):ifc, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000987.

Wang:2023:DPC

[942] Victor Y. Wang. Dichotomous point counts over finite fields. Journal of Number Theory, 250(??):1–34, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2300077X.

Majure:2023:APV

[943] Mitchell Majure. Algebraic properties of the values of newform Dedekind sums. Journal of Number Theory, 250(??):35–48, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23000732.

Langeveld:2023:GSS

[944] Niels Langeveld, Lucía Rossi, and Jörg M. Thuswaldner. Generalizations of Sturmian sequences associated with N-continued fraction algorithms. Journal of Number Theory, 250(??):49–83, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23000781.

Torcomian:2023:AFS

[945] Lucas Villagra Torcomian. Asymptotic Fermat for signature (4, 2, p) over number fields. Journal of Number Theory, 250(??):124–138, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000720.

Creutz:2023:LGP

[946] Brendan Creutz and Sheng (Victor) Lu. The local–global principle for divisibility in CM elliptic curves. Journal of Number Theory, 250 (??):139–154, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000719.

Bieker:2023:IWR

[947] Patrick Bieker. Invariants for the Weil representation and modular units for orthogonal groups of signature (2, 2). Journal of Number Theory, 250(??):155–182, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000756.

Kim:2023:AMC

[948] Chan-Ho Kim and Matteo Longo. Anticyclotomic main conjecture and the non-triviality of Rankin–Selberg L-values in Hida families. Journal of Number Theory, 250(??):183–205, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000744.

Berend:2023:SVI

[949] Daniel Berend, Rishi Kumar, and Andrew Pollington. Simultaneous visibility in the integer lattice. Journal of Number Theory, 250 (??):84–123, September 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000768.

Anonymous:2023:PO

[950] Anonymous. Pages 1–210 (October 2023). Journal of Number Theory, 251(??):1–210, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBj

[951] Anonymous. Editorial Board. Journal of Number Theory, 251(??): ifc, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001105.

Katsurada:2023:ASE

[952] Hidenori Katsurada and Shoyu Nagaoka. On p-adic Siegel Eisenstein series. Journal of Number Theory, 251(??):3–30, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0022314X23000884. See corrigendum [1088].

Gajdzica:2023:LCR

[953] Krystian Gajdzica. Log-concavity of the restricted partition function p_{A}(n, k) and the new Bessenrodt–Ono type inequality. Journal of Number Theory, 251(??):31–65, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000914.

Kumar:2023:ATN

[954] Balesh Kumar and Biplab Paul. Additive twists of L^{2}-norm of Fourier– Jacobi coefficients of Siegel cusp forms. Journal of Number Theory, 251(??):70–83, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000951.

Kim:2023:HDI

[955] Taehyeong Kim, Seonhee Lim, and Frédéric Paulin. On Hausdorff dimension in inhomogeneous Diophantine approximation over global function fields. Journal of Number Theory, 251(??):102–146, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000896.

Moree:2023:PDE

[956] Pieter Moree and Pietro Sgobba. Prime divisors of ℓ-Genocchi numbers and the ubiquity of Ramanujan-style congruences of level ℓ. Journal of Number Theory, 251(??):147–184, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000938.

Keith:2023:PCC

[957] William J. Keith and Fabrizio Zanello. Parity of the coefficients of certain eta-quotients, II: the case of even-regular partitions. Journal of Number Theory, 251(??):84–101, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000963.

Languasco:2023:NEL

[958] Alessandro Languasco. Numerical estimates on the Landau–Siegel zero and other related quantities. Journal of Number Theory, 251(??):185– 209, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300094X.

Stokes:2023:CGF

[959] Matt Stokes. Corrigendum to “On Gauss factorials and their connection to the cyclotomic λ-invariants of imaginary quadratic fields” [J. Number Theory 246 (2023) 279–293]. Journal of Number Theory, 251(??):1–2, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23000860.

Jeon:2023:CVW

[960] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Corrigendum to “On values of weakly holomorphic modular functions at divisors of meromorphic modular forms” [J. Number Theory 239 (2022) 183–206]. Journal of Number Theory, 251(??):66–69, October 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23000926. See

[716].

Anonymous:2023:PN

[961] Anonymous. Pages 1–404 (November 2023). Journal of Number Theory, 252(??):1–404, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBk

[962] Anonymous. Editorial Board. Journal of Number Theory, 252(??):ifc, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001415.

Diaconu:2023:STA

[963] Adrian Diaconu and Henry Twiss. Secondary terms in the asymptotics of moments of L-functions. Journal of Number Theory, 252 (??):243–297, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001178.

Buyukboduk:2023:MCH

[964] Kâzım Büyükboduk and Tadashi Ochiai. Main conjectures for higher rank nearly ordinary families — I. Journal of Number Theory, 252(??): 1–73, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23000872.

Cojocaru:2023:NCE

[965] Alina Carmen Cojocaru and McKinley Meyer. Non-CM elliptic curves with infinitely many almost prime Frobenius traces. Journal of Number Theory, 252(??):74–117, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23000902.

Ding:2023:SSG

[966] Yuchen Ding and Haiyan Zhou. Subset sums over Galois rings II. Journal of Number Theory, 252(??):118–136, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2300118X.

Roehrig:2023:STS

[967] Christina Roehrig. Siegel theta series for quadratic forms of signature (m − 1, 1). Journal of Number Theory, 252(??):137–156, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001166.

Ding:2023:RCP

[968] Zhiguo Ding and Michael E. Zieve. Roots of certain polynomials over finite fields. Journal of Number Theory, 252(??):157–176, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001233.

Suzuki:2023:LCN

[969] Masatoshi Suzuki. Li coefficients as norms of functions in a model space. Journal of Number Theory, 252(??):177–194, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001221.

Takada:2023:LIK

[970] Yuta Takada. Lattice isometries and K3 surface automorphisms: Salem numbers of degree 20. Journal of Number Theory, 252(??):195–242, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300121X.

Kim:2023:RSP

[971] Jiseong Kim. On the Rankin–Selberg problem in families. Journal of Number Theory, 252(??):298–325, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001154.

Gekeler:2023:DMF

[972] Ernst-Ulrich Gekeler. On Drinfeld modular forms of higher rank VI: the simplicial complex associated with a coefficient form. Journal of Number Theory, 252(??):326–378, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001191.

Espinosa:2023:ZFC

[973] Malors Espinosa. Zeta functions of certain quadratic orders. Journal of Number Theory, 252(??):379–404, November 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001208.

Anonymous:2023:PD

[974] Anonymous. Pages 1–378 (December 2023). Journal of Number Theory, 253(??):1–378, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2023:EBl

[975] Anonymous. Editorial Board. Journal of Number Theory, 253(??):ifc, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001580.

Kaczorowski:2023:TDC

[976] J. Kaczorowski and A. Perelli. Twists by Dirichlet characters and polynomial Euler products of L-functions. Journal of Number Theory, 253(??):1–16, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001257.

Zhang:2023:ACC

[977] Huimin Zhang. Averages of coefficients of a class of degree seven L-functions. Journal of Number Theory, 253(??):137–156, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23001245.

Kim:2023:IMZ

[978] Doyon Kim. Infinitely many zeros of additively twisted L-functions on the critical line. Journal of Number Theory, 253(??):157–187, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001361.

Hu:2023:CGQ

[979] Yong Hu, Ahmed Laghribi, and Peng Sun. Chow groups of quadrics in characteristic two. Journal of Number Theory, 253(??):188–214, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001348.

Kala:2023:CFP

[980] Vítězslav Kala and Piotr Miska. On continued fraction partial quotients of square roots of primes. Journal of Number Theory, 253 (??):215–234, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001385.

Bostan:2023:RSC

[981] Alin Bostan, Armin Straub, and Sergey Yurkevich. On the representability of sequences as constant terms. Journal of Number Theory, 253 (??):235–256, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001373.

Huynh:2023:BPT

[982] Dinh Tuan Huynh, Ruiran Sun, and Song-Yan Xie. Big Picard theorem for jet differentials and non-Archimedean Ax–Lindemann theorem. Journal of Number Theory, 253(??):257–277, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001300.

Konieczny:2023:CMS

[983] Jakub Konieczny. Characterisation of Meyer sets via the Freiman– Ruzsa theorem. Journal of Number Theory, 253(??):278–294, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300135X.

Thakur:2023:AVP

[984] Steve Thakur. Abelian varieties with prescribed embedding and full embedding degrees. Journal of Number Theory, 253(??):295–316, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001294.

Beckwith:2023:TTC

[985] O. Beckwith, A. Caione, Z. Chen, M. Diluia, O. Gonzalez, and J. Su. Theta-type congruences for colored partitions. Journal of Number Theory, 253(??):317–336, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001282.

Friedlander:2023:TWG

[986] Holley Friedlander. Twisted Weyl group multiple Dirichlet series over the rational function field. Journal of Number Theory, 253(??):337–367, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001312.

Gow:2023:GGL

[987] Rod Gow and Gary McGuire. On Galois groups of linearized polynomials related to the general linear group of prime degree. Journal of Number Theory, 253(??):368–377, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001476.

Ye:2023:CCY

[988] Dongxi Ye. On a conjecture of Chen and Yui: Fricke groups. Journal of Number Theory, 253(??):17–68, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001269.

Grechuk:2023:DET

[989] Bogdan Grechuk, Tetiana Grechuk, and Ashleigh Wilcox. Diophantine equations with three monomials. Journal of Number Theory, 253 (??):69–108, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001336.

Kodrnja:2023:HIM

[990] Iva Kodrnja and Goran Muić. Hilbert’s irreducibility, modular forms, and computation of certain Galois groups. Journal of Number Theory, 253(??):114–136, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001270.

Djankovic:2023:CIF

[991] Goran Djanković, Dragan – Dokić, Nikola Lelas, and Ilija Vrećica. Corrigendum to “The integrated fourth moment of Dirichlet L-functions over rational function fields” [J. Number Theory 218 (2021) 334–369]. Journal of Number Theory, 253(??):109–113, December 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001397. See

[287].

Anonymous:2024:PJa

[992] Anonymous. Pages 1–358 (January 2024). Journal of Number Theory, 254(??):1–358, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBa

[993] Anonymous. Editorial Board. Journal of Number Theory, 254(??): ifc, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001749.

Mazur:2024:EDD

[994] Barry Mazur, Karl Rubin, and Alexandra Shlapentokh. Existential definability and diophantine stability. Journal of Number Theory, 254(??): 1–64, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001324.

Pasten:2024:SCA

[995] Hector Pasten. Shimura curves and the abc conjecture. Journal of Number Theory, 254(??):214–335, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2300149X.

Lalin:2024:EAM

[996] Matilde N. Lalin and Subham Roy. Evaluations of the areal Mahler measure of multivariable polynomials. Journal of Number Theory, 254(??):103–145, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001531.

Komornik:2024:QEA

[997] Vilmos Komornik, Paola Loreti, and Marco Pedicini. A quasi-ergodic approach to non-integer base expansions. Journal of Number Theory, 254(??):146–168, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001555.

Charlier:2024:PAB

[998] Émilie Charlier, Célia Cisternino, and Savinien Kreczman. On periodic alternate base expansions. Journal of Number Theory, 254(??):184– 198, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001506.

Yang:2024:LVZ

[999] Qiyu Yang. Large values of ζ(s) for 1/2 < ℜ(s) < 1. Journal of Number Theory, 254(??):199–213, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001518.

Dao:2024:BMO

[1000] Quang-Duc Dao. Brauer–Manin obstruction for integral points on Markoff-type cubic surfaces. Journal of Number Theory, 254(??):65– 102, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001488.

Fan:2024:IRS

[1001] Kai Fan and Carl Pomerance. An inequality related to the sieve of Eratosthenes. Journal of Number Theory, 254(??):169–183, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001543.

Kumchev:2024:EBL

[1002] Angel Kumchev, Wade McCormick, Nathan McNew, Ariana Park, Russell Scherr, and Willow Ziehr. Explicit bounds for large gaps between squarefree integers. Journal of Number Theory, 254(??):336– 357, January 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300152X.

Anonymous:2024:PF

[1003] Anonymous. Pages 1–230 (February 2024). Journal of Number Theory, 255(??):1–230, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBb

[1004] Anonymous. Editorial Board. Journal of Number Theory, 255(??): ii, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001828.

Chmiel:2024:CMM

[1005] Tymoteusz Chmiel. Coefficients of the monodromy matrices of oneparameter families of double octic Calabi–Yau threefolds at a halfconifold point. Journal of Number Theory, 255(??):8–36, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23001658.

Kaur:2024:SDM

[1006] Amrinder Kaur and Ayyadurai Sankaranarayanan. The Selberg–Delange method and mean value of arithmetic functions over short intervals. Journal of Number Theory, 255(??):37–61, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001634.

Maiti:2024:SNV

[1007] Gopal Maiti and Kummari Mallesham. Simultaneous non-vanishing of central values of GL(2) × GL(3) and GL(3)L-functions. Journal of Number Theory, 255(??):62–84, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001646.

Wang:2024:AIV

[1008] Wei Wang. Algebraic independence of values of quasi-modular forms. Journal of Number Theory, 255(??):85–97, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001713.

Friedberg:2024:WRG

[1009] Solomon Friedberg and David Ginzburg. On the Whittaker range of the generalized metaplectic theta lift. Journal of Number Theory, 255(??):98–147, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001683.

Wang:2024:DNP

[1010] Wei Wang and Chuangxun Cheng. Distinguishing newforms by the prime divisors of their Fourier coefficients. Journal of Number Theory, 255(??):148–165, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2300166X.

Xue:2024:OSS

[1011] Jiangwei Xue and Chia-Fu Yu. Optimal spinor selectivity for quaternion orders. Journal of Number Theory, 255(??):166–187, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23001671.

Yiasemides:2024:VRS

[1012] Michael Yiasemides. The variance of a restricted sum-of-squares function over short intervals in F_{q}[T]. Journal of Number Theory, 255 (??):188–229, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001695.

Walsh:2024:SGR

[1013] P. G. Walsh. Specializations of a generic rank-2 curve of Shioda. Journal of Number Theory, 255(??):1–7, February 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23001725.

Anonymous:2024:PMa

[1014] Anonymous. Pages 1–396 (March 2024). Journal of Number Theory, 256 (??):1–396, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBc

[1015] Anonymous. Editorial Board. Journal of Number Theory, 256(??): ii, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002226.

Chattopadhyay:2024:NPB

[1016] Jaitra Chattopadhyay and Anupam Saikia. Non-Pólya bi-quadratic fields with an Euclidean ideal class. Journal of Number Theory, 256(??): 1–7, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001877.

Corrigan:2024:MSV

[1017] Chandler C. Corrigan. Mean square values of Dirichlet L-functions associated to fixed order characters. Journal of Number Theory, 256(??): 8–22, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002007.

Juul:2024:BOC

[1018] Jamie Juul. Backward orbits of critical points. Journal of Number Theory, 256(??):23–36, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001932.

Fink:2024:RDN

[1019] Thomas M. A. Fink. Recursively divisible numbers. Journal of Number Theory, 256(??):37–54, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001701.

Li:2024:RLF

[1020] Rao Li and Fan Lü. Run-length function of the beta-expansion of a fixed real number. Journal of Number Theory, 256(??):55–78, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001890.

Banerjee:2024:FCR

[1021] Pradipto Banerjee and Amit Kundu. Factorization of composition of reciprocal polynomials with monomials. Journal of Number Theory, 256(??):79–96, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001907.

Kedzierski:2024:EDM

[1022] Dawid Edmund Kȩdzierski and Piotr Krasoń. On ext^{1} for Drinfeld modules. Journal of Number Theory, 256(??):97–135, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001944.

Xie:2024:SVS

[1023] Bing Xie, Yigeng Zhao, and Yongqiang Zhao. Special values of spectral zeta functions of graphs and Dirichlet L-functions. Journal of Number Theory, 256(??):136–159, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001993.

Stokes:2024:CEC

[1024] Matt Stokes. On CM elliptic curves and the cyclotomic λ-invariants of imaginary quadratic fields. Journal of Number Theory, 256(??):160– 169, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001968.

Liu:2024:CPP

[1025] Kui Liu and Wei Niu. Counting pairs of polynomials whose product is a cube. Journal of Number Theory, 256(??):170–194, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2300197X.

Banerjee:2024:CSA

[1026] Soumyarup Banerjee. On a conjecture of Sun about sums of restricted squares. Journal of Number Theory, 256(??):253–289, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23001919.

Clark:2024:DIS

[1027] Pete L. Clark, Paul Pollack, Jeremy Rouse, and Katherine Thompson. Densities of integer sets represented by quadratic forms. Journal of Number Theory, 256(??):290–328, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23001956.

Baker:2024:BNM

[1028] Roger C. Baker, Changhao Chen, and Igor E. Shparlinski. Bounds on the norms of maximal operators on Weyl sums. Journal of Number Theory, 256(??):329–353, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23001889.

Cesana:2024:AEP

[1029] Giulia Cesana, William Craig, and Joshua Males. Asymptotic equidistribution for partition statistics and topological invariants. Journal of Number Theory, 256(??):373–396, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002147.

Hiary:2024:IEE

[1030] Ghaith A. Hiary, Dhir Patel, and Andrew Yang. An improved explicit estimate for ζ(1/2 + it). Journal of Number Theory, 256(??):195– 217, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23001920.

Boudreau:2024:FEC

[1031] Félix Baril Boudreau. L-functions of elliptic curves modulo integers. Journal of Number Theory, 256(??):218–252, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23001981.

Hathi:2024:SPS

[1032] Shehzad Hathi and Daniel R. Johnston. On the sum of a prime and a square-free number with divisibility conditions. Journal of Number Theory, 256(??):354–372, March 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002044.

Anonymous:2024:PAa

[1033] Anonymous. Pages 1–354 (April 2024). Journal of Number Theory, 257 (??):1–354, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBd

[1034] Anonymous. Editorial Board. Journal of Number Theory, 257(??): ii, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000052.

Ren:2024:RTT

[1035] Xiumin Ren, Qingqing Zhang, and Rui Zhang. Roth-type theorem for quadratic system in Piatetski–Shapiro primes. Journal of Number Theory, 257(??):1–23, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23002093.

Mahajan:2024:HMP

[1036] Jewel Mahajan and Kaneenika Sinha. Higher moments of the pair correlation function for Sato–Tate sequences. Journal of Number Theory, 257(??):24–97, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X2300207X.

Pauli:2024:BSC

[1037] S. Pauli and F. Saidak. A bound for Stieltjes constants. Journal of Number Theory, 257(??):112–123, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002032.

Nadimpalli:2024:ILA

[1038] Santosh Nadimpalli and Mihir Sheth. On the integrality of locally algebraic representations of GL_{2}(D). Journal of Number Theory, 257(??):

124–145, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002111.

Lu:2024:OCR

[1039] Dawei Lu and Ruoyi Wang. Optimizing the coefficients of the Ramanujan expansion. Journal of Number Theory, 257(??):146–162, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002081.

Zhang:2024:VPN

[1040] Wei Zhang. On a variant of the prime number theorem. Journal of Number Theory, 257(??):163–185, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002123.

Long:2024:RUC

[1041] Hanqing Long and Dasheng Wei. The 3-rd unramified cohomology for norm one torus. Journal of Number Theory, 257(??):186– 201, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002160.

Benmerieme:2024:AAC

[1042] Y. Benmerieme and A. Movahhedi. Ankeny–Artin-Chowla and Mordell conjectures in terms of p-rationality. Journal of Number Theory, 257(??): 202–214, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002056.

Lorenzini:2024:CPG

[1043] Dino Lorenzini. The critical polynomial of a graph. Journal of Number Theory, 257(??):215–248, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002135.

Martin:2024:MFE

[1044] Kimball Martin and Satoshi Wakatsuki. Mass formulas and Eisenstein congruences in higher rank. Journal of Number Theory, 257(??):249– 272, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2300210X.

Garcia:2024:SSP

[1045] Jonathan García, Carlos A. Gómez, and Florian Luca. Solving Skolem’s problem for the k-generalized Fibonacci sequence with negative indices. Journal of Number Theory, 257(??):273–299, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23002184.

Ma:2024:MDC

[1046] Jing Ma and Ronghui Wu. Möbius disjointness conjecture on the product of a circle and the Heisenberg nilmanifold. Journal of Number Theory, 257(??):300–319, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23002068.

Ghioca:2024:COO

[1047] Dragos Ghioca. Collision of orbits for a one-parameter family of Drinfeld modules. Journal of Number Theory, 257(??):320–340, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23002342.

Shallit:2024:RTM

[1048] Jeffrey Shallit. Rarefied Thue–Morse sums via automata theory and logic. Journal of Number Theory, 257(??):98–111, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23002159.

Chow:2024:VFP

[1049] Sam Chow and Owen Jones. On the variance of the Fibonacci partition function. Journal of Number Theory, 257(??):341–353, April 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X23002329.

Anonymous:2024:PM

[1050] Anonymous. Pages 1–450 (May 2024). Journal of Number Theory, 258 (??):1–450, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBe

[1051] Anonymous. Editorial Board. Journal of Number Theory, 258(??): ii, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000209.

Im:2024:GTG

[1052] Bo-Hae Im and Hansol Kim. Growth of torsion groups of elliptic curves over number fields without rationally defined CM. Journal of Number Theory, 258(??):1–21, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23002391.

Kimura:2024:CMD

[1053] Shotaro Kimura. Construction of modular differential equations by using Rankin–Cohen brackets. Journal of Number Theory, 258(??):22– 39, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002275.

Quadrelli:2024:MPG

[1054] Claudio Quadrelli. Massey products in Galois cohomology and the elementary type conjecture. Journal of Number Theory, 258(??):40– 65, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002263.

Chen:2024:IF

[1055] Yong-Gao Chen andJi Zhen Xu. On integers of the form p + 2k_{1}^{r_{1}} + ⋯ + 2k_{t}^{r_{t}}. Journal of Number Theory, 258(??):66–93, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23002408.

Hung:2024:CNM

[1056] Fan-Yun Hung and Chia-Fu Yu. Class numbers of multinorm-one tori. Journal of Number Theory, 258(??):94–121, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23002378.

Dalal:2024:BSW

[1057] Tarun Dalal. A basis for the space of weakly holomorphic Drinfeld modular forms of level T. Journal of Number Theory, 258(??):122– 145, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X23002330.

Xu:2024:EVD

[1058] Max Wenqiang Xu and Daodao Yang. Extreme values of Dirichlet polynomials with multiplicative coefficients. Journal of Number Theory, 258(??):173–180, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23002317.

Au:2024:MIC

[1059] Kam Cheong Au. Moments of ideal class counting functions. Journal of Number Theory, 258(??):181–211, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X23002354.

Uttenthal:2024:DSR

[1060] Peter Vang Uttenthal. Density of Selmer ranks in families of even Galois representations, Wiles’ formula, and global reciprocity. Journal of Number Theory, 258(??):212–268, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002305.

Ikeda:2024:RZH

[1061] Karin Ikeda. On real zeros of the Hurwitz zeta function. Journal of Number Theory, 258(??):269–280, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2300238X.

Das:2024:MPD

[1062] Mithun Kumar Das and Sudhir Pujahari. Mean of the product of derivatives of Hardy’s Z-function with Dirichlet polynomial. Journal of Number Theory, 258(??):334–367, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X23002366.

Lin:2024:OLD

[1063] Hua Lin. One-level density of zeros of Dirichlet L-functions over function fields. Journal of Number Theory, 258(??):368–413, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X24000015.

Venkatasubbareddy:2024:CKF

[1064] Kampamolla Venkatasubbareddy and Ayyadurai Sankaranarayanan. On certain kernel functions and shifted convolution sums of the Fourier coefficients. Journal of Number Theory, 258(??):414–450, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X24000118.

Capunay:2024:ARC

[1065] Alex Capuñay, Milton Espinoza, and Eduardo Friedman. Attractor– repeller construction of Shintani domains for totally complex quartic fields. Journal of Number Theory, 258(??):146–172, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X23002299.

Huber:2024:RTC

[1066] Timothy Huber, Nathaniel Mayes, Jeffery Opoku, and Dongxi Ye. Ramanujan type congruences for quotients of Klein forms. Journal of Number Theory, 258(??):281–333, May 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X23002287.

Anonymous:2024:PJb

[1067] Anonymous. Pages 1–438 (June 2024). Journal of Number Theory, 259 (??):1–438, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBf

[1068] Anonymous. Editorial Board. Journal of Number Theory, 259(??): ii, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000477.

Yang:2024:UST

[1069] Zichen Yang. Universal sums of triangular numbers and squares. Journal of Number Theory, 259(??):1–15, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X24000155.

Lunia:2024:QDF

[1070] Rashi Lunia. On quotients of derivatives of L-functions inside the critical strip. Journal of Number Theory, 259(??):16–37, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X24000088.

Oshita:2024:FEG

[1071] Kazuki Oshita and Masao Tsuzuki. Functional equations and gamma factors of local zeta functions for the metaplectic cover of SL_{2}. Journal of Number Theory, 259(??):57–81, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2400012X.

Akande:2024:NPI

[1072] Agbolade Akande. The number of preimages of iterates of φ and σ. Journal of Number Theory, 259(??):82–92, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X2400009X.

Bordelles:2024:SSS

[1073] Olivier Bordellès, Randell Heyman, and Dion Nikolic. Sparse sets that satisfy the prime number theorem. Journal of Number Theory, 259(??):93–111, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000106.

Lagarias:2024:FFG

[1074] Jeffrey C. Lagarias and Wijit Yangjit. The factorial function and generalizations, extended. Journal of Number Theory, 259(??):131– 170, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000143.

Bravo:2024:QBT

[1075] Claudio Bravo. Quotients of the Bruhat–Tits tree by function field analogs of the Hecke congruence subgroups. Journal of Number Theory, 259(??):171–218, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000234.

Chidambaram:2024:MGR

[1076] Shiva Chidambaram. Mod-p Galois representations not arising from abelian varieties. Journal of Number Theory, 259(??):219–237, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000374.

Muller:2024:NGG

[1077] Peter Müller. A note on Galois groups of linearized polynomials. Journal of Number Theory, 259(??):238–241, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X24000416.

Zhou:2024:ELC

[1078] Nian Hong Zhou. Eventual log-concavity of k-rank statistics for integer partitions. Journal of Number Theory, 259(??):242–272, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000258.

Vargas-Montoya:2024:AMG

[1079] Daniel Vargas-Montoya. Algebraicity modulo p of generalized hypergeometric series _{n}F_{n−1}. Journal of Number Theory, 259(??):273– 321, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000271.

Heath-Brown:2024:BQW

[1080] D. R. Heath-Brown. Bounds for the quartic Weyl sum. Journal of Number Theory, 259(??):330–337, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2400043X.

Lee:2024:FTR

[1081] Yoonjin Lee and Donghyeok Lim. The finitude of tamely ramified pro-p extensions of number fields with cyclic p-class groups. Journal of Number Theory, 259(??):338–356, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000283.

Kabluchko:2024:DPP

[1082] Zakhar Kabluchko and Alexander Marynych. Divisibility properties of polynomial expressions of random integers. Journal of Number Theory, 259(??):357–377, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000398.

Buium:2024:LIF

[1083] Alexandru Buium. Lie invariant Frobenius lifts. Journal of Number Theory, 259(??):378–418, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000404.

Schmiedt:2024:PCF

[1084] Anja Schmiedt and Christian Weiß. The pair correlation function of multi-dimensional low-discrepancy sequences with small stochastic error terms. Journal of Number Theory, 259(??):422–437, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X24000295.

Scherr:2024:QIP

[1085] Zachary Scherr and Katherine Thompson. Quartic integral polynomial Pell equations. Journal of Number Theory, 259(??):38–56, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000167.

Advocaat:2024:COR

[1086] Bryan Advocaat. Computations on overconvergence rates related to the Eisenstein family. Journal of Number Theory, 259(??):112– 130, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000131.

Campbell:2024:RTS

[1087] John M. Campbell. On a Ramanujan-type series associated with the Heegner number 163. Journal of Number Theory, 259(??):322– 329, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000325.

Katsurada:2024:CAS

[1088] H. Katsurada and S. Nagaoka. Corrigendum to “On p-adic Siegel Eisenstein series” [J. Number Theory 251 (2023) 3–30]. Journal of Number Theory, 259(??):419–421, June 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000428. See [952].

Anonymous:2024:PJc

[1089] Anonymous. Pages 1–266 (July 2024). Journal of Number Theory, 260 (??):1–266, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBg

[1090] Anonymous. Editorial Board. Journal of Number Theory, 260(??): ii, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000726.

Gazda:2024:RSF

[1091] Quentin Gazda and Andreas Maurischat. Residue of special functions of Anderson A-modules at the characteristic graph. Journal of Number Theory, 260(??):1–28, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000386.

Geisser:2024:ICF

[1092] Thomas H. Geisser and Baptiste Morin. On integral class field theory for varieties over p-adic fields. Journal of Number Theory, 260(??): 41–70, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2400026X.

Dong:2024:HOT

[1093] Janet J. W. Dong and Kathy Q. Ji. Higher order Turán inequalities for the distinct partition function. Journal of Number Theory, 260(??): 71–102, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000301.

Bell:2024:RSM

[1094] Jason Bell, Dragos Ghioca, and Zinovy Reichstein. Rational self-maps with a regular iterate on a semiabelian variety. Journal of Number Theory, 260(??):103–119, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000350.

Fedosova:2024:SCS

[1095] Ksenia Fedosova and Kim Klinger-Logan. Shifted convolution sums motivated by string theory. Journal of Number Theory, 260(??):151– 172, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000362.

Dembele:2024:FFG

[1096] Lassina Dembélé and René Schoof. Finite flat group schemes over Z killed by 19. Journal of Number Theory, 260(??):191–195, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X24000568.

Jiang:2024:ZSP

[1097] Xiao Jiang and Wenkai Yang. A zero-sum problem related to the max gap of the unit group of the residue class ring. Journal of Number Theory, 260(??):198–211, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000520.

Aslanyan:2024:NUI

[1098] Vahagn Aslanyan and Christopher Daw. A note on unlikely intersections in Shimura varieties. Journal of Number Theory, 260(??):212– 222, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000611.

Fang:2024:CNM

[1099] Jin-Hui Fang. On Cilleruelo–Nathanson’s method in Sidon sets. Journal of Number Theory, 260(??):223–231, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X24000507.

Wang:2024:UDA

[1100] Bo Wang, Bing Li, and Ruofan Li. Uniform Diophantine approximation with restricted denominators. Journal of Number Theory, 260(??): 232–265, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000581.

Bertok:2024:RTE

[1101] Csanád Bertók and Lajos Hajdu. The resolution of three exponential Diophantine equations in several variables. Journal of Number Theory, 260(??):29–40, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000313.

Fan:2024:NEE

[1102] Kai Fan. Numerically explicit estimates for the distribution of rough numbers. Journal of Number Theory, 260(??):120–150, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000349.

Blaser:2024:STT

[1103] Avalon Blaser, Molly Bradley, Daniel A. N. Vargas, and Kathy Xing. Sato–Tate type distributions for matrix points on elliptic curves and some K3 surfaces. Journal of Number Theory, 260(??):173–190, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000337.

Sun:2024:CDF

[1104] Haiwei Sun and Yangbo Ye. Corrigendum to “Double first moment for L(\frac{1}{2}, Sym^{2}f × g by applying Petersson’s formula twice” [J. Number Theory 202 (2019) 141–159]. Journal of Number Theory, 260(??): 196–197, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000519. See [7].

Anonymous:2024:PAb

[1105] Anonymous. Pages 1–382 (August 2024). Journal of Number Theory, 261 (??):1–382, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2024:EBh

[1106] Anonymous. Editorial Board. Journal of Number Theory, 261(??): ii, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X2400101X.

Lou:2024:SCS

[1107] Miao Lou. Shifted convolution sums of divisor functions with Fourier coefficients. Journal of Number Theory, 261(??):1–21, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000659.

Peng:2024:HDC

[1108] Liuqing Peng, Jun Wu, and Jian Xu. Hausdorff dimension of certain sets related to random αβ-orbits which are not dense. Journal of Number Theory, 261(??):22–35, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000532.

Fischler:2024:RBV

[1109] S. Fischler and T. Rivoal. Relations between values of arithmetic Gevrey series, and applications to values of the gamma function. Journal of Number Theory, 261(??):36–54, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2400057X.

Harun:2024:SCS

[1110] Mohd Harun and Saurabh Kumar Singh. Shifted convolution sum with weighted average: GL(3) × GL(3) setup. Journal of Number Theory, 261(??):55–94, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000623.

Cohen:2024:KVD

[1111] Jonathan Cohen. Klingen vectors for depth zero supercuspidals of GSp(4). Journal of Number Theory, 261(??):95–124, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X2400060X.

Gao:2024:FMC

[1112] Peng Gao and Liangyi Zhao. First moment of central values of some primitive Dirichlet L-functions with fixed order characters. Journal of Number Theory, 261(??):125–142, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X24000593.

Chan:2024:SAM

[1113] Tsz Ho Chan. Spectrum of all multiplicative functions with application to powerful numbers. Journal of Number Theory, 261(??):143– 160, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000660.

Bahnson:2024:TPL

[1114] Erik Bahnson, Mark McConnell, and Kyrie McIntosh. Tempered perfect lattices in the binary case. Journal of Number Theory, 261(??):161– 189, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000672.

Browning:2024:SFV

[1115] Tim D. Browning and Igor E. Shparlinski. Square-free values of random polynomials. Journal of Number Theory, 261(??):220–240, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000635.

Clozel:2024:ULB

[1116] Laurent Clozel and Peter Sarnak. A universal lower bound for certain quadratic integrals of automorphic L-functions. Journal of Number Theory, 261(??):252–298, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000684.

Jackson:2024:PIP

[1117] Faye Jackson and Misheel Otgonbayar. Parts in k-indivisible partitions always display biases between residue classes. Journal of Number Theory, 261(??):299–311, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X24000556.

Cho:2024:DMV

[1118] Hansle Cho. Discrete mean values of the product of derivatives of shifted Dirichlet L-functions. Journal of Number Theory, 261(??):312– 326, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000787.

An:2024:LFZ

[1119] Chen An. Log-free zero density estimates for automorphic L -functions. Journal of Number Theory, 261(??):327–382, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X2400088X.

Grantham:2024:FPP

[1120] Jon Grantham and Andrew Granville. Fibonacci primes, primes of the form 2^{n} − k and beyond. Journal of Number Theory, 261(??):190– 219, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000544.

Carrillo:2024:FGS

[1121] Benjamin Carrillo. Finding Galois splitting models to compute local invariants. Journal of Number Theory, 261(??):241–251, August 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X24000647.